A Complete Bibliography of The Mathematical Gazette (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

20 May 2024 Version 1.10

Title word cross-reference

$ [Tol21b]. (1/4)(a + b + c + d) [Lor23l]. (2n − 1) [Cag21]. (3k − 4) [Cag21]. (n − 1) [Cag21]. \frac{(ϕ(m^{r}))^{a}}{(ϕ(n^{s}))^{b}} [Vu22]. 1 + 1 = 11 [PS22a]. 1/sinθ [Jam23b]. $15.95 [Aar22]. 2 + 2 = 5 [PS22a, Mac23c]. $25 [Gri20, Hal21]. $26.99 [Ste22b]. 2 × 2 [Has21b, Spo24]. $32 [Haw20]. $35 [Hun24]. 3 × 3 [Dow20]. $44.95 [Lev24]. $51 [Hun22e]. $85 [Hun22b]. $89.00 [Hun22d]. b > e [Cha24]. b^{2} = \frac{1}{2} (3 + \sqrt{5}) ac [McL20]. b^{e} < e^{b} [Cha24]. c^{2} = a^{2} + bd [Lau23a]. cos 2x + sin 2x = 1 [Ale22]. cos β − cos α [AU22]. e^{x/y} [Gho22a]. k [Cag21, Cag23]. lim_{n→∞} ^{pn} \sqrt{p_{1}p_{2} ⋯ p_{n}} = e [Far21]. log x [Jam23d]. N! [Sul22, Has24]. n [Hun22i, RV24, ST23, San23, Sin22, Ste23b]. N^{p/q} [Gho22b]. ϕ [GA23]. π [LN22b, Sao20, Sin20]. sin 3x = 3 sin x − 4 sin^{3} x [ST22]. sin α + sin β [AU22]. \sqrt{2} [MGC20]. ∑_{n=0}^{∞} \frac{1}{(2n+1)^{2}} = \frac{π^{2}}{8} [Mar22]. ∑_{n=2}^{∞} 1/(nH_{n−1}) [AP21]. ∑_{n=2}^{∞} 1/(nH_{n−1}^{1+ϵ}) [AP21]. ∑ k^{2} [Lor22h]. ∑ k^{3} [Lor22h]. tan(5π/12) = 2 + \sqrt{3} [Luk24a]. tan(π/12) = 2 − \sqrt{3} [Luk24a]. θ/sinθ [Jam23b]. x = a^{x} [Bea22b]. x^{3} + y^{3} = (x + y)(x^{2} − xy + y^{2}) [Cha21]. x^{3} − y^{3} = (x − y)(x^{2} + xy + y^{2}) [Cha21]. x^{y} = y^{x} [BG20]. xy = cos(x + y) [Jew24]. y = E(x) [BG20]. y = mx − 2m − m^{3} [AT21]. ζ(3) [Sin20].

-dimensional [Hun22i]. -gonal [Cag21, Cag23]. -gons [RV24].

1 [Tol24]. 1900 [Mac24]. 1950 [Hun23b].

2 [Lev22d]. 2017 [Ryb20, Lev20a]. 2018 [Lev20b]. 2019 [Lev22e]. 2020 [Ano20c, Ano20e]. 2021 [Ano20d, Ano21d, Gib23b, Lev23d]. 2022 [Lor23i, YC23]. 2023 [Ano22a, Ano23c, Fos23]. 2024 [Ano23d, Ano24a]. 20758-2 [Lor23d]. 279-9 [Lor22c]. 2nd [Dav22, Hun22g, Lor20b, Mac20].

3000 [Gib23b]. 3rd [Tol22a, Tol22d].

4th [Tol22e].

58221-0 [Gib23b].

978-0-00-832458-2 [Hun21e]. 978-0-12-820788-8 [Tol23e]. 978-0-12-823417-4 [Tol22a]. 978-0-19-875535-7 [Cri22b]. 978-0-19-882971-3 [Slo24]. 978-0-19-883344-4 [Tol21d]. 978-0-19-883567-7 [Tol21c]. 978-0-19-884638-3 [Lev22f]. 978-0-19-884643-7 [Lev22b]. 978-0-19-886902-3 [Tol23h]. 978-0-19882-122-9 [Hun20c]. 978-0-19883-160-0 [Bay20b]. 978-0-2280-0373-1 [Cri22d]. 978-0-24423-100-2 [Tol22c]. 978-0-262-04664-0 [Lev24]. 978-0-262-54300-2 [Cri24]. 978-0-26253-902-9 [Aar22]. 978-0-300-25539-3 [Tol22f]. 978-0-30024- [Lor22c]. 978-0-367-19557-1 [Tol22b]. 978-0-367-36272-0 [Cri22c]. 978-0-367-56303-5 [Lor23e]. 978-0-367-62837-6 [Shi24]. 978-0-36721-936-9 [Shi22a]. 978-0-46509-481-3 [Haw21a]. 978-0-46509-760-9 [Haw20]. 978-0-521-72839-3 [Hun22a]. 978-0-578-61682-7 [Ste22b]. 978-0-69-118278-0 [Hun21d]. 978-0-691- [Lor23d]. 978-0-691-15883-9 [Lor21c]. 978-0-691-17691-8 [Bay21a]. 978-0-691-17863-9 [Lev20a]. 978-0-691-18264-3 [Cri22a]. 978-0-691-18276-6 [Lev20b]. 978-0-691-18277-3 [Hoa20]. 978-0-691-18367-1 [Tol22g]. 978-0-691-19232-1 [Pra21]. 978-0-691-19835-4 [Lev22e]. 978-0-691-20219-8 [Hew23]. 978-0-691-20613-4 [Hew22]. 978-0-691-22348-3 [Mac23b]. 978-0-691-22570-8 [Lev23d]. 978-0-691-23524-0 [Hun23b]. 978-0-69117-941-4 [Rua23b]. 978-0-69119-922-1 [Tol23f]. 978-0-69121-435-1 [Tol22h]. 978-0-69121-876-2 [Hun23c]. 978-0-8153-7097-0 [Dav22]. 978-0-90658-889-5 [Haw21b]. 978-019884759 [Tol22e]. 978-0198847618 [Tol22d]. 978-1-009-00919-5 [Hal23]. 978-1-009-09629-4 [Col24]. 978-1-00900-162-5 [Yeo23]. 978-1-107-13057-9 [Tol21e]. 978-1-107-15613-5 [dV21a]. 978-1-107-17790-1 [Bay22]. 978-1-10717-314-9 [Yeo20]. 978-1-10718-233-2 [Hun21b]. 978-1-108-41089-2 [Hun21c]. 978-1-108-41090-8 [Hal22c]. 978-1-108-43224-5 [Hun21a]. 978-1-108-43679-3 [Lor20b]. 978-1-108-44102-5 [Hun20a]. 978-1-108-45644-9 [Hun22c]. 978-1-108-70300-0 [Hop22a]. 978-1-108-70937-8 [Luk23b]. 978-1-108-70945-3 [Luk23b]. 978-1-108-72262-9 [Hun22g]. 978-1-108-73838-5 [Lor22e]. 978-1-108-78998-1 [Tol22i]. 978-1-108-79920-1 [Shi23a]. 978-1-108-83496-4 [Tol23d]. 978-1-108-83665-4 [Tol23b]. 978-1-108-92740-6 [Mac23a]. 978-1-108-95823-3 [Tol23c]. 978-1-10833-745-8 [Gib22a]. 978-1-10842-579-7 [Hun20b]. 978-1-10843-953-4 [Gib22a]. 978-1-10847-059-9 [Sch21a]. 978-1-10895-972-8 [Lor22d]. 978-1-316-51027-8 [Mac23b]. 978-1-42142-407-1 [Jac20]. 978-1-42143-308-0 [Hun22h]. 978-1-4704-4871-4 [Hun22e]. 978-1-4704-5084-7 [Hun22d]. 978-1-4704-5342-8 [Hun22b]. 978-1-4704-5425-8 [Hun22f]. 978-1-4704-6307-6 [Hun23a]. 978-1-4704-6858-3 [Hun24]. 978-1-4717-6147-8 [Lor24c]. 978-1-4718-5307-4 [Lev22d]. 978-1-4718-8648-5 [Lev22c]. 978-1-5104-1451-8 [Lev23b, Lev23c]. 978-1-5104-3337-3 [Lev23a]. 978-1-5294-1125-6 [Haw24]. 978-1-5416-0036-2 [Bay23a]. 978-1-5416-9949-6 [Bay23b]. 978-1-61614-931-4 [Bay21b]. 978-1-63388-373-4 [Tol21b]. 978-1-63388-387-1 [Gri20]. 978-1-63388-520-2 [Hal21]. 978-1-73314-663-0 [Gib22b]. 978-1-78125-433-2 [Jac21]. 978-1-78205-290-6 [Cri21b]. 978-1-78747-542-7 [Shi22b]. 978-1-78816-228-9 [Tol21a]. 978-1-80061-121-4 [Tol24]. 978-1-90600-133-9 [Hal20a]. 978-1-91161-600-9 [Hal20b]. 978-1-911616-08-5 [Hal22b]. 978-1-912827-03-9 [Hal22a]. 978-107-12413-4 [Cri21a]. 978-3-030- [Gib23b]. 978-3-030-04036-9 [Mal22]. 978-3-030-63812-2 [Rua23a]. 978-3-030-70574-9 [Mac24]. 978-3-030-79430-9 [Luk24b]. 978-3-030-87173-4 [Shi23c]. 978-3-03052-810-2 [Tol23g]. 978-3-0348-0621-3 [Cri21c]. 978-3-319-70631-3 [Rou20b]. 978-3-319-74747-7 [Mac20]. 978-3-319-77636-1 [Sch20]. 978-3-319-77835-8 [Rou20a]. 978-3-319-90914-1 [Bay20a]. 978-3-662-62688-7 [Gib24]. 978-3-662-62689-4 [Gib24]. 978-981-15-9218-8 [Shi23b]. 978-981-16-6364-2 [Tol23a]. 99 [Lor21c].

ABC [Gri24]. ABC-triangles [Gri24]. Ablowitz [Lor22d]. above [Bea21a]. Abstract [Tol23b]. Academic [Tol22a, Tol23e]. Acheson [Lev22f]. Acknowledgements [Ano23a, Lev22a]. across [Ban23]. activities [SWT21]. Adam [Hal22a]. Adams [Hun24]. addition [Spo23]. additive [Sur20].

Address [Pri22, Fos23]. Advanced [Ste22b]. Advocate [Lor24e]. Aesop [All21]. Africa [Mal22]. AG [Gib23b]. again [Lor22g, Ste22a]. Alfred [Bay21b, Gri20, Hal21]. algebra [Gib22b, Hun22d, Hun22e, Hun22f, Mac20, Tol23b, Bay22, Tol23g, Lor23e, Tol23c]. algebraic [Tos23, Rou20a]. Algorithms [Aar22]. Allan [Hun22b]. Almost [Fal21]. along [Kob23]. Alsamraee [Ste22b]. Also [Bay21b, Gib22a, Gib23b, Haw21a, Hun21b, Mac20, Sch20, Tol22b, Tol23a]. Alternating [Ber20]. alternative [MY22a, Ort20]. altitudes [Luk20a, Tho22e]. Aluffi [Tol23c]. AM [Haj20, Mes23]. Amendment [YC23]. American [Hun22b, Hun22d, Hun22e, Hun22f, Hun23a, Hun23b, Hun24]. Amir [Tol22i]. among [Haj21, Spo24]. AMS [Hun22d]. Anachronisms [Tol23d]. analogue [Sur20]. analyses [Lor21b]. analysis [Gup24, Gib23b, Hun21b, Hun22a, Lor20b, Lor23d, Mac20, Shi23b, Yeo20]. analytic [Fre22, HS20]. anarchy [Wap20]. Andrei [Luk24b, Shi23b]. Andrew [Cri24, Hal20a, Hun20c, Tol23h, Tol24]. Andrews [Lev24]. Angell [Shi24]. angle [ASAMHH21, Ban23, Dal21, Haj22, Lam21, Lor21d]. angled [Nie21]. angles [ASH20a]. answer [Bea22c, LH22, Mac21c]. Anthony [Tol22h]. antiderivatives [EH23]. Anton [Cri22a]. anything [Hun21e]. appearance [Lor24b]. application [Tho22d]. Applications [VVK21, GS22, Gîr23, Lor22d]. Applied [Mac20]. appreciation [LR21]. appreciations [Mah21a]. approach [All20, HP23, Lor23c, Rou20b]. approached [HS20]. approximate [Hun22c]. Approximating [Bel22, Jam23a]. approximation [Jam20, Mah22b, Vil23]. approximations [Lor21d]. apps [Wap20]. April [Ryb20]. arbitrariness [Haj22]. archery [Sko20]. Archimedean [DT23]. Archimedes [Dub22, Wil23]. architecture [Cri24]. area [Mey20, SWT21, Vol22, Hal22b]. areas [Fox20, Haj21, Bea21a]. Ariel [Tol22i]. arising [Ale22]. arithmetic [ASH20b, Cer23, Pla22a, SE20a, SO23]. arithmetical [Les20]. art [Bay23a, Lev24]. Arts [Cri22a, Cri22a]. Association [Hal20b, Hal22b, Haw21b, Ryb20]. Athanassios [Lor22d]. Atrium [Cri21b]. attached [Mac22]. Aubel [PdV22]. Ault [Jac20]. available [Bay21b, Gib22a, Gib23b, Haw21a, Hun21b, Mac20, Sch20, Tol22b, Tol23a]. average [SS20]. averages [Jam23c, Lau23b]. axis [HE24].

B [Hun21d, Tol21c, Yeo20]. Back [Ano20f, Ano20h, Ano21e, Ano21g, Ano22c, Ano22e, Ano22g, Ano23e, Ano23g, Ano23i, Ano24b, Hun21e, Mal22, Ste22c]. Baldwin [Lev22c]. Bankoff [Jia22, Luk23d]. Barrow [Luk23d]. barycentric [Vol22]. base [FL22]. based [Pla22a, Tos23]. Basic [Bay23a, Bay23b, Haw20, Jac21, Tol23e]. beautiful [Gib24]. Beetham [Nar20]. being [Gho22a]. below [Bea21a]. Ben [Lev22c]. Bernoulli [Kac20]. Bernstein [Jam20]. Bertrand [Sil24]. best [Hun23c, Lev20a, Lev20b, Lev22e, Lev23d]. better [Bay23a]. betting [AMMW22]. between [Abe20, Cag21, Sar23, Shi23a, Ste23a]. Beyond [Haw21a]. Bicentric [Ste23a, Jos22]. Bickle [Hun22b]. bifocal [RZ21]. big [Hun20c]. Bijan [Tol23a]. bijections [Bea23b]. billions [Hoa20]. Binet [SE20a]. binomial [Abe20, KM21, Lor24e]. Birkhäuser [Cri21c, Gib23b, Bay20a]. birthday [Bev22]. bisect [dV21b]. bisect-diagonal [dV21b]. bisectors [ASAMHH21, Haj22]. Bjorn [Hun20b]. Blog [Ano22b]. boarding [GS21, KNS23]. Bohemia [Kit21]. Boissonnat [Hun21c]. Bolton [Hun20a]. Bonnet [Rid21a]. Book [Aar22, Bay20b, Bay20a, Bay21a, Bay21b, Bay22, Bay23a, Bay23b, Col24, Cri21b, Cri21c, Cri21a, Cri22a, Cri22b, Cri22d, Cri22c, Cri24, Dav22, Gib22a, Gib22b, Gib23b, Gib24, Gri20, Hal20b, Hal20a, Hal21, Hal22a, Hal22b, Hal22c, Hal23, Haw20, Haw21a, Haw21b, Haw24, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22b, Hun22c, Hun22a, Hun22d, Hun22e, Hun22f, Hun22h, Hun22g, Hun23a, Hun23b, Hun24, Hun20c, Hun21e, Hun21d, Hun23c, Jac20, Jac21, Lev20a, Lev20b, Lev22e, Lev22b, Lev22d, Lev22c, Lev22f, Lev23d, Lev23b, Lev23c, Lev23a, Lev24, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23e, Lor23d, Lor24c, Luk23b, Luk24b, Mac20, Mac23a, Mac23b, Mac24, Mal22, Pra21, Rou20a, Rou20b, Rua23b, Rua23a, Sch20, Sch21a, Shi22b, Shi22a, Shi23a, Shi23b]. Book [Shi23c, Shi24, Slo24, Ste22b, Tol21a, Tol21b, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22f, Tol22c, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23e, Tol23f, Tol23a, Tol23g, Tol23h, Tol24, Yeo20, Yeo23, dV21a, Bay21b]. Books [Ano20a, Ano21a, Bay21b, Bay23a, Bay23b, Gri20, Hal21, Hal22a, Haw20, Haw21a, Haw24, Jac21, Shi22b, Tol21a, Tol21b]. Boole [Cri21b]. Bottema [Pel22]. bound [Has24]. bounded [She23, Shi23a]. bounds [Ayd24, Zam20]. Bourchtein [Luk24b, Shi23b]. box [Jah22]. Brahmagupta [Mey20]. Brian [Haw24, Hoa20]. brief [Mah21b]. Bring [Kul21]. Brocard [ASH20a, LS21]. Broughan [Shi23a]. Brown [Tol21c]. Brummelen [Rua23b]. Bueno [Tol22h]. Burger [Hun21d]. Butterworth [Haw24].

C [Hun22c, Hun22a, Hun20c, Slo24, Tol23e]. calculate [Hun21e, Pat22]. calculated [Lam21]. Calculating [Jac21]. calculus [Cri21c, Ste22b, Lor22c, Sch20]. calendar [Gad20]. Cambridge [Bay22, Col24, Cri21a, Gib22a, Gib22b, Hal22c, Hal23, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22c, Hun22a, Hun22g, Lor20b, Lor22d, Lor22e, Luk23b, Mac23a, Mac23b, Sch21a, Shi23a, Tol21e, Tol22i, Tol23c, Tol23d, Tol23b, Yeo20, Yeo23]. Can [Haw24]. card [Pra21]. Cardano [LSS23, Ohy20, Ohy22]. cardioids [Sil21b]. cards [De22b]. care [SA20]. Carlo [Sie20]. case [Lor20a]. Cassini [Bea22a]. Catalan [Ste20c]. Cath [Lev22d]. Cauchy [Dun23, Far23, Fre22]. cautionary [Lor24d, Uhl23]. Cayley [Lor23f]. Ceccherini [Hun21b]. Ceccherini-Silberstein [Hun21b]. Cédric [Lev22b]. central [Cag20]. centre [Lor23l]. centred [Cag23]. centres [Luk21a]. Centroid [Ste20b]. centroids [Fri22]. century [Lev24]. certain [Bou21]. Ceva [Haj23a]. Chain [Sie20]. Challal [Tol22b]. chances [Tol23h]. Change [Mac24]. changed [Cri22a]. changers [Tol21b]. Chapman [Cri22c, Hop22a, Shi24]. Characterisation [Nak22, Jos20, RV24]. characterisations [HK23a, HK23b, Jos22]. characteristic [Con23, Wap21]. charms [Mac23a]. Chazal [Hun21c]. chemist [Lor22g, Ste22a]. Cheng [Haw21a]. Cheryl [Cri22b]. Chiossi [Shi23c]. Chris [Hal20b, Hal22b]. Christian [Gri20, Hal21]. Chvátal [Mac23a]. Cinderella [Lev21a]. circle [Bea23b, Gor23b, Lev21a, Lor23a, Luk23a, PdV22, Sil23a]. circle-preserving [Bea23b]. Circles [Spo22b, Bea24, DT23, Ste23a]. circular [De22a, Nie21]. circum [Hum20]. circum-medial [Hum20]. circumcircle [Luk23a]. circumcircles [Ste23a]. circummidarc [Luk20b]. circumradii [Rea23]. Claire [Lev22c]. class [Hun21a, LN22a, Rea22]. classic [Fri22]. classical [Jah22]. classroom [SWT21]. clearly [Tol22h]. Clever [Hun21e, Hun23c]. clock [Che23]. close [Jam20]. coalescing [Sam23]. coefficients [Jam23b, Zam20]. cofactor [Dow20]. Cohen [Cri21b]. coincidence [Gad20]. Colin [Hun24]. Collatz [Sam23]. Collins [Hun21e]. colonial [Mal22]. come [Lor22e]. Comic [Lor24c]. commensurability [RV24]. companion [Tol22f]. complementary [Nic20b]. complex [Hau23, Kir20, Lor22d, Lor23d, Sil23b, Lor20b, Shi23b]. composite [Con23, Gri21]. composites [LM24]. comprehensive [Tol23b]. computations [Uhl23]. computer [Dub22, Sul22]. Computing [NT23]. concerning [Fri22]. conclusive [SR20]. Concurrent [Kac20]. Conditional [Has21b]. conditions [Sam23]. cone [Sin21, Sin24]. cones [De22a, Rid21a]. configuration [PdV22, Vig20]. congruence [Kon20, SE20b, Seb22]. conical [De24, Sta20]. conics [Bea21b, Bea23c, RZ21]. conjecture [Sam23, Tho22c]. conjunctions [Che23]. connections [Cal20a, Sar23]. consecutive [Ste20a]. conservation [SWT21]. considerations [SA23d]. constant [Mer23, Ste20c]. constants [Sch21a]. Constructing [Fox20]. constructions [KF24]. contact [Luk23a]. containing [CJS20]. continued [KS21]. continuous [Jam23a]. convergence [Mer23]. convergents [LN22b]. converges [AP21]. converse [KNS23]. converses [Sil23a]. convex [Dal22, Jam22b]. convexity [BG20, Jam21d, Ott24]. Conway [Ryb20, Yeo20]. coordinates [KF21b, Lor24d, Vol22]. copy [Tol22i]. Core [Lev22c]. Cork [Cri21b]. Corner [Ano24d, Dol20c, L.21a, L.21b, L.22c, L.22a, L.22b, L.23, L.24, Lor20c, Lor20d, Lor23j, Lor23k]. Correct [Mac21c, Bea22c, LH22]. Correspondence [Mah22a, Por22, Ric24, San24]. cosine [Gor23a, VVK21]. cosmos [Jac21]. Could [Lam21]. count [Haw24]. counterexample [Sam23]. Counterexamples [Yeo23]. course [Hun20a, Hun20b, Hun21a, Lor23d, Tol23a, Yeo20]. cousins [CF21]. Cover [Ano20g, Ano20f, Ano20i, Ano20h, Ano21f, Ano21e, Ano21h, Ano21g, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23f, Ano23e, Ano23h, Ano23g, Ano23j, Ano23i, Ano24c, Ano24b]. Cramer [Hu22]. cranks [Lor22c]. CRC [Cri22c, Dav22, Lor23e, Shi22a, Shi24, Tol22b]. CRC/Taylor [Lor23e]. criterion [Nat21]. Croarken [Cri22c]. Crossword [Sta21b]. Cryptography [Tos23]. cubes [Ber22, Jam22a, Kat23, Rid21a]. cubic [Fri22, LSS23, MY22a, Nic20b, Ohy20]. cubics [De21a]. culture [Cri24]. Cumulative [Ano21b, Ric21a, Ric21b]. curios [Lor23b]. Curiosities [Bay21b]. Curious [Ste22b, Nic20a, Hew22]. curve [Bea21a]. curves [AT21, Fri22, Hew22]. cut [Kat23]. cycles [Nor20]. cyclic [Dal21, Fri22, Mey20].

Daniel [Col24, Gib22a, Hun21c]. data [Tol21e, Tol22h]. David [Hun22h, Lev22f, Lor20b, Rua23a, Shi24, Tol22d, Tol22e]. Davis [Hal22a]. Davvaz [Tol23a]. death [Shi22b]. DeBevoise [Lev22b]. debt [SA23d]. December [MC21, Ryb20]. decimal [Pas21]. defined [BG20]. defining [Ott24]. definite [Abe23, HP23]. definition [Bau21]. deflection [Lam21]. delights [Hew23]. derivation [Bau22, Kul21]. derivative [Ham20, SR20]. Desargues [Gîr23, Sil20]. Descartes [Kit21]. Desmond [Cri21b]. determinant [Dow20, Has21b]. determinants [MGC20]. Developing [Mah21c, ST23]. development [Luk23b]. Devil [Lor24e]. diagonal [dV21b]. dice [Tol21a]. difference [Cag21, Cag23, Luc24, NT23]. differential [All20, Geo20a, Hun20a, Hun20b, Mac24, Tol22a]. differentiating [KF23]. diffusion [Hop22a]. digamma [KS23]. Digit [PPP23]. Digital [Gup24]. Digne [Hun22g]. dimensional [Hun22i]. dimensions [Pra21]. diophantine [MS23]. direct [Dal21]. direction [Whi22]. Dirk [Mal22]. disc [LS21]. discontinuous [Ham20]. discover [PS22b]. discovered [Hun23c]. Discrete [Lev23a, Cri20, Mac23a, She23, Ste21a, Hun21b]. discs [LS22]. Dissecting [Mac22]. dissections [Fre21]. distance [Tho22a]. distances [Gib23a]. distinguishing [Tol21e]. distribution [Abe20]. Divergence [Fel20]. diverges [AP21]. divisibility [Seb22]. divisible [Bro22]. divisors [Sin22]. do [Lor22e, PS22a, Tol21a, Tol23f]. doctrine [Rua23b]. dodecahedra [Kat23]. dodgy [Bea22c, LH22, Mac21c]. does [Ban23, Lor23l, OS23]. Donoghue [Mac23b]. Dropping [Hop22b]. dualities [Dal22]. duality [Dal21, KF21a, KF24]. Dundas [Hun20b]. dynamic [Bea21b].

e-book [Gib22a, Gib23b, Haw21a, Mac20, Sch20, Tol22b, Tol23a, Bay21b]. e-copy [Tol22i]. e-version [Hal22c]. ears [Haj21]. Eastaway [Hun21e]. easy [Jew24, Mey20]. eBook [Gib24, Hun21b]. ed [Lev20a, Lev20b, Lev22e, Lev23d]. edited [Cri22c, Haw21b, Tol23d]. edition [Gib22a, Hal23, Lor23e, Luk23b, Tol22d, Tol22e]. Editor [Ano23b]. edn [Dav22, Hun22g, Lor20b, Mac20, Tol22a]. Education [Lev22d, Lev22c, Lev23b, Lev23c, Lev23a]. Edward [Hun21d, Tol22f]. effective [Sib23]. efficiency [Sib23]. Eilers [dV21a]. Elap [Sta21b]. elementary [Dow20, Kro24, Mac21a]. elements [Has21b]. elephant [Hal20b]. Eli [Hew23]. Elisabeth [Kit21]. Elizabeth [Bay22]. Ellina [Bay20a]. Elliott [Hun20c, Tol23h]. Elsevier [Tol23e]. Elsevir [Tol22a]. emergence [FL22]. Emmy [Rua23a]. encountered [SS20]. end [SA20]. enigma [Cus21]. Enomoto [Oku23]. entertainments [Bay21b]. envelope [Hun21e]. Equal [Jam22a, Abe23]. equalisation [Nak22]. Equality [Haj21]. equation [KF23, MS23, MY22a, MY22b]. Equations [Bur20, All20, Geo20a, Jew24, Luc24, Mac24, NT23, TT24, Tol22a]. equilateral [Ber20]. equivalent [Lor23a]. era [Hun23b]. erase [Tol23f]. Erban [Hop22a]. Erdos [Mac23a]. Eric [Hal22c, Tol21c]. error [Abe20, Bay20b, Lor21b]. essential [Sch21b, Shi23c]. estimates [Jam21c]. Ethan [Tol22h]. Euclid [Haj23b]. Euclidean [Tol22b]. Eugenia [Haw21a]. Euler [BB20, CC24, GA23, HLM24, Mer23, Nim22]. Eureka [Dol21b, Dol24]. evaluation [Lor22h, Mah20]. even [Jam21b]. Every [Cag20]. everyday [Gri20]. everyone [Gib22b]. Evolution [Nim22]. evolving [De24]. example [Uhl23]. examples [AMMW22]. Exarc [Luk22a]. Excel [Hau23]. excentral [Luk20a]. excess [Cri22b]. exercises [Tol22d]. exist [OS23]. expansion [Dow20, Lor24e]. expansions [Lor21a, Pas21]. Expected [CK21]. experimental [dV21a]. exploration [DL21]. explored [Ste22b]. exponentials [FP24]. expression [Con23]. expressions [Gri21]. Extending [Ohy20]. extension [Lau23a, PdV22]. Extensions [BB20, DT24, Now20, Tra24]. extraordinaire [Rua23a].

Fabio [Hun21b, Tol22g]. face [Fox20]. fact [Lor23f, Tol21e]. Factorial [Sil21a]. factorials [Gad20]. factors [Sil21a]. fall [Bay21a]. Families [AT21]. family [Cri20, Ste21a]. Farewell [Ano24d]. Farey [Sch21b]. Feedback [Ano20b, Ano21c, Bur22, Lor23i, YC23]. Fermat [Kon20, Sco22b, Sin21, TT24]. Fermat-like [TT24]. Fibonacci [Bea22a, Gri21, SE20a, SE20b, Spo22b, Spo22a, Ste23e, Hal22a]. fiction [Tol21e]. field [Hun21a, Mac23b]. fifth [Bau22]. figurate [Cag22]. figures [Bur20, Fre21]. Filip [Sch20]. Filippo [Hun21b]. final [OI23]. financial [SA23d]. Finch [Sch21a]. Finding [Hum20]. findings [Mal22]. finite [Hun22g, Now20, Sie20]. Finsler [Lor24a, Luk20b, Luk22a]. first [Hun20a, Mac23b, ST23, San23, Sco22b, Sin22, Ste23b, Tol23a, Yeo20]. fish [Haw24]. Fleisch [Col24]. Flood [Cri22c]. fluency [Lor20a]. flying [Cri22d]. Focus [Pri22]. Fokas [Lor22d]. forgotten [Ste22c]. form [ASH20b, OS23, Sco22a, SO23, Vu22]. formula [Cer23, Mar22, Mey20, MY22a, Nim22, SE20a, Tol22g]. formulae [Lau22, Lor21b, NO22]. Formulations [Cri24]. founded [Sci24]. Four [Gib23a]. Fourier [Jam23b]. Fowler [Tol22h]. Fraction [Spo22b]. fractional [TT24]. fractions [KS21, LN22b]. Francis [Lor23e]. Francois [Hun22g]. Frank [Tol23b, Cri22b]. Franziska [Yeo23]. Frédéric [Hun21c]. Frederick [Hun22e, Hun22f]. free [Bro22, Jam21b]. French [Tol24]. frequency [Sta21a]. Friday [Sta21a]. Front [Ano20g, Ano20i, Ano21f, Ano21h, Ano22d, Ano22f, Ano22h, Ano23f, Ano23h, Ano23j, Ano24c]. FRS [Ryb20]. Fruit [MS23]. frustum [Bel22]. full [Lor22f]. function [BG20, Con23, EH23, GA23, Ham20, KS23, Mun20, Pla22a, Roi23, Ste22c]. function-based [Pla22a]. functional [Hun22a, KF23]. functions [Bev22, CJS20, Jam21a, Jam22b, Jam23a, Nie21]. Fundamentals [Hun22b]. Further [Lev22c, Lev23b, Lev23c, Lev23a].

G [Shi23c]. game [CK21, FL22, Wap23, Pra21, Tol21b]. games [AMMW22]. gap [Sib23]. gaps [Shi23a]. Garfunkel [Jia22, Luk23d]. Garfunkel-Bankoff [Jia22]. Garrity [Hal23]. Gary [Pra21]. Gauss [Dol21b, Dol24, Rid21a, Shi20]. Gazette [Ano21b]. Geere [Lev23a]. general [KF23, LSS23]. generalisation [HE24, Jah22]. generalisations [Tra23, Tra24]. generalised [Nat22, PdV22, KM21, Lau23b]. generate [Jam21d]. generation [Lor20a]. generator [DT23]. gentle [Hun21a]. Geoffrey [Tol22d, Tol22e, Gar23]. geometric [Dol20a, Dow20, Hu22, Lig23, Mac21b, Muk23, Mun20, OS23, Pla22a, Hun21c]. Geometrical [SWT21]. geometry [Bea21b, Dal21, Dav22, Fre22, HS20, Kro24, Lev21a, Lev24, Lor23e, Oku23, Tra24, Hun20a, Lev22f, Hun23a]. George [Cri21b]. Georgiev [Cri21c]. Gerretsen [Sco22a]. Giessen [Buh23]. Gilbert [Gib22b]. give [Lor23l]. given [Fox20, Has21b, Hum20, Whi22]. giving [Vig20]. Glen [Rua23b]. GM [Haj20, Mes23]. God [Tol21a]. Godfried [Dav22]. Goldbach [Fal21]. Golden [Sci22, Cus21, Lor24b, Spo21b, Yos22, Sci24]. Goldie [Lev22d]. gonal [Cag21, Cag23]. gons [RV24]. good [Dav22, GT23, Lev21b, Luk22b, Ste23d]. Gordon [Pra21]. GPS [Wap20]. graduate [Hal23]. Graham [LR21]. graph [GA23, Hun22b]. gravity [Bay21a]. Gray [Mac24]. Greenleaf [Hun22e, Hun22f]. Greenwich [Cri22c]. grid [SS24]. Griffiths [Slo24, Cer23]. Grigorieva [Bay20a]. Grimmett [Tol22d, Tol22e]. group [All20, Spo21a, Spo21b, Tol23a]. Groups [Bea23c, Hun22g, Mac21d, Hun22c]. Guicciardini [Tol23d]. guide [Col24]. Guillot [Hun21a].

H [Hun22c]. habit [Ask20]. Hadwiger [Lor24a, Luk20b, Luk22a]. haiku [Hil21]. Hall [Cri22c, Shi24]. Hall/CRC [Cri22c, Shi24]. Halving [Whi22]. Hamza [Ste22b]. handbook [Cri21c]. Hannah [Pra21]. Hanrahan [Lev22d]. hard [Bay20b, Bay20a, Bay21a, Bay22, Bay23b, Cri22a, Cri22b, Gib22b, Gib23b, Gri20, Haw20, Haw24, Hoa20, Hun20b, Hun21b, Hun22d, Hun20c, Hun21e, Hun21d, Jac20, Jac21, Lor21c, Lor23d, Mac20, Mac23b, Mac24, Rou20a, Rou20b, Rua23b, Sch21a, Shi23c, Shi24, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22g, Tol23d, Tol23b, Tol23f, Tol23a, Tol23g, Tol23h, dV21a]. hardback [Gib22a, Mal22]. Hardy [Jam23c]. harmonic [Fel20, Hun21b, Pat22, Pla22a, Sco20, Vil23]. Harper [Hun21e]. Hart [Hal22a]. Hart-Davis [Hal22a]. Havil [Hew22]. Hawk [Lor22c]. Hayk [Rou20a, Rou20b]. Heard [Cri21a]. Heinz [Gib24]. Henry [Tol22a]. Hermite [Nat21, Nat22]. Heronian [Dol20b]. Hex [Ber22]. highway [SS20]. history [Cri22c, Gîr23, Mac24, Tol23d]. Hoare [LR21]. Hodder [Lev22d, Lev22c, Lev23b, Lev23c, Lev23a]. Hole [Fre21]. Honorary [Ryb20]. Hopkins [Hun22h, Jac20]. Horton [Ryb20]. Howard [Tol21e]. Howes [Cri22a]. Howson [Gar23]. Hunger [Hun23b]. Huygens [Lor22a]. Huylebrouck [Mal22]. hyperbolas [Spo22a]. hyperbolic [Bea20, Nie21]. hypergeometric [Abe20].

Ian [Bay23b, Hun20b, Jac21, Lor20b, Tol21a]. icosahedra [Kat23]. identities [Kir20, Ste23c]. identity [Bea22a, BB20, Kol22, Spo24, Ste24]. II [Hun22f, Sch21a]. illuminated [RZ21]. Illustrating [Hau23]. illustration [Lig23]. illustrative [Bau22]. implicit [Jew24]. Impossibility [Ohy22]. impress [Lor22g, Ste22a]. improvement [Jia22]. Improving [Fos23]. inarc [Luk21a, Luk23a]. incentre [Luk21a, LS22]. incircles [Ste23a]. inclusive [SR20]. incorrect [Pau21]. increasing [Zam20]. indefinite [NT23]. Indeterminate [FP24]. Index [Ano21b, Ric21a, Ric21b, AMMW22]. inductive [Mes23]. inequalities [BK21, HP23, Jam23d, KS23, Lor22a, Luk23d, Pla22a, Sco22a, Tho22e, Rou20a]. inequality [Far23, Fre22, Haj20, HL23, Jam23c, Jia22, Lor22a, Lor24a, Luk20b, Luk20a, Luk22a, Luk23d, Mes23, OS22, SS22, Tho22d, Tra23]. inference [Hun21c]. infinite [Lig23, Luk24b, Pas21]. Infinitely [LM24, Ale22]. infinitesimal [Bau21]. infinity [Haw21a]. Ingmar [Bay21b]. inradius [Spo21a, Tho22e]. inscribed [Pla20b]. inspired [Ste20c]. Institute [Cri24]. instructive [Tol21f]. Integer [Rea23, Rea22, Ste23b]. integers [Cha22, ST23, San23]. Integral [BK21, Sin24, Mah20, NO22, Ort20, Sin20, Ste20c, Ste23e]. Integrals [CJS20, Abe23, Jam21c, LN22a, NT23]. Integrating [Gor23a]. integration [HP23, Yeo23]. interesting [DB22, DT23, LN22a, Lor23a, Tho22d]. interevent [She23]. interpretation [Hu22, Uhl23]. intriguing [De22b]. introduction [Hun22a, Lor23e, Mac23b, Tol22a, Jac20, Tol23b, Hun22c, Lor22d, Tol22b, Tol23g, dV21a]. Introductory [Tol21c]. invariant [Lor22f]. inverse [EH23, HL23, Ste22c]. inverses [Cal20b]. inverted [Sin21]. investigation [Cal20a]. involution [Gîr23]. involving [Cag23, De22a, Tho22e]. irrational [Gho22a]. irrationality [Gho22b, Tho22b, MGC20, Shi24]. ISBN [Aar22, Bay20b, Bay20a, Bay21a, Bay21b, Bay22, Bay23a, Bay23b, Col24, Cri21b, Cri21c, Cri21a, Cri22a, Cri22b, Cri22d, Cri22c, Cri24, Dav22, Gib22a, Gib22b, Gib23b, Gib24, Gri20, Hal20b, Hal20a, Hal21, Hal22a, Hal22b, Hal22c, Hal23, Haw20, Haw21a, Haw21b, Haw24, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22b, Hun22c, Hun22a, Hun22d, Hun22e, Hun22f, Hun22h, Hun22g, Hun23a, Hun23b, Hun24, Hun20c, Hun21e, Hun21d, Jac20, Jac21, Lev20a, Lev20b, Lev22e, Lev22b, Lev22d, Lev22c, Lev22f, Lev23d, Lev23b, Lev23c, Lev23a, Lev24, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23e, Lor23d, Lor24c, Luk23b, Luk24b, Mac20, Mac23a, Mac23b, Mac24, Mal22, Pra21, Rou20a, Rou20b, Rua23b, Rua23a, Sch20, Sch21a, Shi22b, Shi22a, Shi23a, Shi23b, Shi23c]. ISBN [Shi24, Slo24, Ste22b, Tol21a, Tol21b, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22f, Tol22c, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23e, Tol23f, Tol23a, Tol23g, Tol23h, Tol24, Yeo20, Yeo23, dV21a]. Ishango [Mal22]. isomorphism [Mac21d]. issue [Ano20g, Ano20f, Ano20i, Ano20h, Ano21f, Ano21e, Ano21h, Ano21g, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23f, Ano23e, Ano23h, Ano23g, Ano23j, Ano23i, Ano24c, Ano24b].

J [Bay20b, Bay21a, Gib22a, Hun20a, Hun23c, Lor22d, Tol22a]. Jackson [Jam23a]. Jakob [GS22]. James [Hun22a, Hun23a]. Jean [Hun21c, Hun22g, Lev22d, Lev23b, Lev23c]. Jean-Daniel [Hun21c]. Jean-Paul [Lev22d, Lev23b, Lev23c]. Jeremy [Mac24]. Jessica [Tol23f]. João [Cri21c]. Jobbings [Hal20a]. Johannes [Coo21]. Johansen [dV21a]. John [Cri21a, Mac23b, Tol23b, Yeo20, Ryb20]. Johns [Hun22h, Jac20]. Johnston [Tol23g]. Jonathan [Hop22a]. journey [Cri21a]. journeys [Hal22b]. Joy [Pra21, Tol21d]. Judea [Haw20]. Julian [Hew22]. July [Ano20c, Ano21d, Ano23c, Ano24a, Bur22, Lor23i]. justification [Dow20].

Kahn [Tol21c]. Karen [Hun23b]. Katz [Lor22c]. keeps [Vig20]. Kepler [Coo21, Sci24]. Kernighan [Hoa20]. Kevin [Shi23a]. King [Hun23a]. Kirkman [Sil20]. Kit [Shi22b]. Klaus [Gib24]. know [Hal23]. Körner [Lor22e]. Kühn [Yeo23].

L [Hun20a, Yeo23]. Langley [BM23, Sil22]. Laplace [Col24, KM21]. large [Hun23c]. last [Sin21]. lattice [Lor23b]. law [Lau22]. Lawrence [Shi22a, Tol23b]. Lawson [Lor23e]. LCM [Sin22]. Learning [De21b, Sko20, Tol21e]. least [Hun23c]. left [Pas21]. left-sided [Pas21]. Legendre [Shi20]. Lehmann [Bay21b]. Lehmus [Haj23a, Haj23b]. Leibniz [Awa24]. Lemoine [HE24, Sco23]. length [CK21]. lengths [OS23, SO23]. Less [Fos23]. letter [Tol22c]. level [Lev22d, Lev22c, Lev23b, Lev23c, Lev23a]. Lie [Hun22g, All20]. life [Shi22b, SA20, Bay23a, Gri20]. light [Cri21b]. like [TT24, Tol21e, Wap23]. limit [Bau21, CC24]. limit-revisited [CC24]. Limits [Geo20a, Pla20a]. Lindsay [Hun22h]. line [Ban23, Lor24a, Sco23]. linear [Mac20, Tol23g, Bay22, Gib22b, Hun22d, Hun22e, Hun22f]. Linearly [KS21]. lines [AT21, Lor22f]. link [Nie21]. Lipschitz [Jam23a]. Listener [Sta21b]. Liz [Pra21]. local [Hun21a]. Location [Luk23a]. Locus [Fri22]. Loftus [Tol23e]. logarithm [Ort20]. Logic [Lor24c, Tol22c, Slo24]. Long [ASAMHH21, Lev24, Ste22c]. long-forgotten [Ste22c]. look [Ste22c, Sul22]. Lorenzo [Mac23b]. lost [GS21, KNS23, Hun24]. Louridas [Aar22]. lover [Tol22f]. lower [Has24]. Lucas [SE20b, Spo22a]. Ludmila [Luk24b, Shi23b]. Lukarevski [Tho22c].

M [Hun20a]. MA [Hal22b, Ano22b]. MacHale [Cri21b, Lor24c, Tol22c]. Maclaurin [Gor23a, Lor21a, Nim22]. Madhava [Awa24]. MAG [Ano20g, Ano20f, Ano20i, Ano20h, Ano21f, Ano21e, Ano21h, Ano21g, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23f, Ano23e, Ano23h, Ano23g, Ano23j, Ano23i, Ano24c, Ano24b]. majorisation [Jam22b]. make [Hun23c]. makers [Hal21]. makes [Dav22, GT23, Lev21b, Luk22b, Ste23d]. making [Cal20a, Coo21, Hun21d]. Malcolm [Lev22b]. manifesto [Slo24]. Mann [Cri22c]. many [Ale22, Hun23c, LM24, Pra21]. Maor [Hew23]. mappings [Hau23]. Maps [PPP23]. March [Ano20d, Ano22a, Ano23d]. Marcus [Bay23a]. Mariette [Hun21c]. Mark [Bay22, Lor22d, Lor23e].

Markov [Sie20]. Marques [Hun22e, Hun22f]. Mary [Cri22c]. mass [Lor23l]. Massachusetts [Cri24]. Math [Hun24, Ste22b, Bay23a, Hal21]. Mathematical [Hal20b, Hal22b, Haw21b, Hun22b, Hun22d, Hun22e, Hun22f, Hun23a, Hun24, Ryb20, Ask20, Cal20a, Kob23, Lor20a, Mac23a, Mah21a, Mah22b, Pra21, SA23d, Bay21b, Sch21a]. mathematically [Hew22, Pau21]. mathematician [Rua23a, Hun22d, Shi22a]. mathematicians [Cri22d, Hun23c, Lor20a, Mac23b]. Mathematics [Hal20a, Lev22d, Lev22c, Lev23b, Lev23c, Lev23a, Row22, Cri22c, Cri24, Gri20, Hal22b, Hal23, Hun23b, Lev20a, Lev20b, Lev22e, Lev23d, Mal22, Shi23c, Tol21c, Tol22f, dlR22, Cri21a, Lor23e, Luk23b, Rou20b, Tol23d, dV21a, Cri22c, Gib24, Lev22b]. Mathematikum [Buh23]. maths [GT23, Shi22b, Hun21e]. Matrices [Cal20b, Bea24, Has21b]. Matrix [Kob20, Hil21, Mac20, Spo24, Tol23g]. matter [Ano20g, Ano20f, Ano20i, Ano20h, Ano21f, Ano21e, Ano21h, Ano21g, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23f, Ano23e, Ano23h, Ano23g, Ano23j, Ano23i, Ano24c, Ano24b]. Matthew [Hun22c]. May [Gar23]. McGill [Cri22d]. McGill-Queen [Cri22d]. McMahon [Pra21]. mean [Pla22a]. measure [Yeo23]. Mechanics [Lev23b, De22b]. Meckes [Bay22]. medial [Hum20]. median [Vu21, Vu21]. median-orthologic [Vu21]. Median-parallelogic [Vu21]. medians [ASAMHH21]. Member [Ryb20]. members [Hal22b, Haw21b]. memory [Mac21b]. mensuration [Lor21b]. meridian [Cri22c]. Mesquita [Tol22h]. method [Bea22c, LH22, Mac21c, MY22b, Pat22]. Methods [Bay20a]. metric [Gor23b]. Michael [MC21]. Michel [Hun22g, Mac23b]. midpoint [Jam21c]. midsquare [Jos20]. Millions [Hoa20]. mind [Ask20, Hun21d]. minds [Cri22a]. Minimum [Pla20b, Sam23]. Mircea [Lev20a, Lev20b, Lev22e, Lev23d]. Misak [Cri22b]. missed [Hal23]. mistake [De21b]. Möbius [Joh21]. model [Che23]. Modelling [Sko20, Hop22a]. Modern [Hal22a, Dun23, GS22, Tol22a]. modified [Fel20]. modular [SE20a]. Moivre [Luc24]. Mollweide [Lau22]. monist [Slo24]. monotone [EH23]. Monotonic [LN22b, Jam21d, Jam21a]. Monotonicity [Jam21c]. Monte [Sie20]. Moore [Lev22d]. Morais [Cri21c]. moral [All21]. Morley [Sil21b]. Muldoon [Tol21c]. Multiplication [Spo23]. multiplicative [Sur20]. Muscat [Lev22d, Lev23b, Lev23c]. Museum [Hun24]. musical [Dav22]. mystic [Ste23c].

Nahin [Bay21a, Hun23c]. Nairi [Rou20a, Rou20b]. Nathaniel [Tol23g]. nation [Cri22a]. natural [Sin22]. Nature [Gib23b]. navigation [Wap20]. near [LN22b]. needed [Hal23]. nested [Lor22b]. Nets [Lor22c]. Neumann [MC21]. news [Row22]. Newton [Lam21, Lau22]. next [Lor20a]. Niccolò [Tol23d]. Nick [Lev23a]. nim [FL22]. NJL [Ano24d]. no [Mel21, Pas21]. Noam [Lev24]. Noether [Rua23a]. non [Mac21d]. non-isomorphism [Mac21d]. note [Lor21d]. Notes [ASH20a, Tol23c, Sta21b]. novel [MY22b]. November [Gar23, Ano20e, YC23]. number [Bay20a, Cag20, Hun20c, Mac21a, Pat22, Row22, Sar23, Shi24, Sin22, Ste24, Tol22c]. numbers [Bea22a, Ber22, Bro22, Cag21, Cag22, Cag23, Gup24, Jam21b, Kir20, Les20, Lor22e, Now20, SE20a, Sil23b, Sin22, Ste23e, Vu22, Hun22h]. numerical [Uhl23]. Nurturing [Lor20a].

OBE [MC21]. observations [Lor22a, Has22]. Obtaining [KF23]. OCR [Lev22d, Lev22c, Lev23b, Lev23c, Lev23a]. octagonal [Cag20]. odd [Cag20, Jam21b, Kol22, Lor23f, ST23, San23]. odd-powers [Kol22]. odyssey [Ste20c]. off [DB22]. old [Lor21b, Nic20a]. one [Lor24a, Sch21b, Slo24, Tol22d]. one-line [Lor24a]. open [Jah22]. open-top [Jah22]. optimisation [De22a, Lor22i, Lor23c]. optimization [Tol22b]. orbits [Sam23]. order [Geo20a]. ordinary [Geo20a]. Ording [Lor21c]. orthogonal [AT21]. orthologic [Vu21]. Osterlind [Bay20b]. other [Cag22, GS21, Jew24, Pat22]. Owen [Slo24]. own [Hun21d]. Oxford [Bay20b, Cri22b, Hun20c, Lev22b, Lev22f, Slo24, Tol21c, Tol21d, Tol22d, Tol22e, Tol23h].

P [Hun22e, Hun22f, Cri21c]. Pack [Hal22b]. pair [Jam23d]. Pairing [KF21a]. pairwise [Bea24]. pairwise-tangent [Bea24]. palindromic [OI23]. pancake [Sao20]. Panos [Aar22]. Paolo [Tol23c]. paper [Aar22, Bay21b, Col24, Cri21b, Cri21a, Cri22d, Cri22c, Cri24, Gib22a, Gib24, Hal20b, Hal20a, Hal22a, Hal22c, Hal23, Haw21a, Haw21b, Hew22, Hew23, Hun20a, Hun21c, Hun21a, Hun22c, Hun22a, Hun22e, Hun22f, Hun22g, Hun23a, Hun23b, Hun23c, Lor20b, Lor22c, Lor22d, Lor22e, Lor23e, Lor24c, Luk24b, Mac23a, Mac23b, Pra21, Rua23a, Sch20, Shi22b, Shi22a, Shi23a, Shi23b, Ste22b, Tol21a, Tol21b, Tol22f, Tol22c, Tol22d, Tol22e, Tol22i, Tol22h, Tol23c, Tol23e, Tol24, Yeo23]. paperback [Dav22, Hop22a, Hun24, Luk23b]. parabolas [KF21a]. Parabolic [KF21b]. paradox [Bev22]. Parallel [Geo20b, Kob23]. parallelogic [Vu21]. parallelograms [HK23a, HK23b]. parameters [CJS20]. parental [CF21]. Parshall [Hun23b]. parts [TT24]. Pascal [Sil20]. Paseau [Slo24]. pass [GS21, KNS23]. Passionate [Kit21]. paths [Kob23]. pattern [Cal20a]. Patterns [Spo24]. Paul [Bay21a, Hun23c, Lev22d, Lev23b, Lev23c, Mac23a]. Paying [SA20]. Pearl [Haw20]. Pedro [Cri21c]. Pell [SE20b]. pendulum [De24]. penguin [Hal22b]. Pentagonal [Cag22]. perfect [Les20, Sha24]. perimeter [Pla20b]. periodic [KS21, Joh21]. periods [SE20a]. perpendicularities [Tos23]. perspectives [Jam23d]. pertinent [Ayd24]. Peter [MC21]. Philip [Lor21c]. physical [Mah21a, Mah22b]. Pic [Sch21b]. picture [ST23]. Pie [Haw21b]. Pierre [Hun21a]. Pipeline [Ano20c, Ano20d, Ano20e, Ano21d, Ano22a, Ano23c, Ano23d, Ano24a]. Pitici [Lev20a, Lev20b, Lev22e, Lev23d]. Pitot [Bea21b]. planar [Ber22, Fre21]. plane [Bea20, Bea23b, Bur20, HS20]. planetary [Che23]. plates [Hop22b]. play [Tol21a]. playing [Dub22]. plot [Jew24]. plus [Lor24c]. poetry [Lev22b]. point [GS22, Luk23a, PdV22, Sco22b]. points [ASH20a, Gib23a]. polar [Lor24d]. pole [Lor24d]. polygonal [Sar23]. polygons [Lau23b, Lor22b, Lor23b, Ste23a, Kir20]. polyhedra [Ott24]. polyhedrists [Lev24]. polynomial [HLM24, Nic21]. polynomials [Jam20, Jam23a, Kob20, Nat21, Nat22, Zam20, dlR22, Mel21]. Posamentier [Bay21b, Gri20, Hal21]. positive [Cha22]. possible [Hun23c]. postulate [Sil24]. Potter [Mah21b]. Power [GS22, Wap21]. powers [Bau22, Cer23, Cha22, Kol22, Sha24, Cri22b]. pp [Aar22, Bay20b, Bay20a, Bay21a, Bay21b, Bay22, Bay23a, Bay23b, Col24, Cri21b, Cri21a, Cri22a, Cri22b, Cri22d, Cri22c, Cri24, Dav22, Gib22a, Gib22b, Gib23b, Gib24, Gri20, Hal20b, Hal20a, Hal21, Hal22a, Hal22b, Hal22c, Hal23, Haw20, Haw21a, Haw21b, Haw24, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22b, Hun22c, Hun22a, Hun22d, Hun22e, Hun22f, Hun22h, Hun22g, Hun23a, Hun23b, Hun24, Hun20c, Hun21e, Hun21d, Hun23c, Jac20, Jac21, Lev20a, Lev20b, Lev22e, Lev22b, Lev22d, Lev22c, Lev22f, Lev23d, Lev23b, Lev23c, Lev23a, Lev24, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23e, Lor23d, Lor24c, Luk23b, Luk24b, Mac20, Mac23a, Mac23b, Mac24, Mal22, Pra21, Rou20a, Rou20b, Rua23b, Rua23a, Sch20, Sch21a, Shi22b, Shi22a, Shi23a, Shi23b, Shi23c]. pp [Shi24, Slo24, Ste22b, Tol21a, Tol21b, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22f, Tol22c, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23e, Tol23f, Tol23a, Tol23g, Tol23h, Tol24, Yeo23, dV21a]. practical [GS21, Jac20]. predictive [Wap21]. prelude [Mac23b]. present [Shi22a]. preserving [Bea23b]. Presidential [Fos23, Ano22b, Pri22]. Press [Aar22, Bay20b, Bay21a, Bay22, Col24, Cri21b, Cri21a, Cri22a, Cri22b, Cri22d, Cri24, Dav22, Gib22a, Gib22b, Hal22c, Hal23, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22c, Hun22a, Hun22h, Hun22g, Hun23b, Hun20c, Hun21d, Hun23c, Jac20, Lev20a, Lev20b, Lev22e, Lev22b, Lev22f, Lev23d, Lev24, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23d, Lor24c, Luk23b, Mac23a, Mac23b, Pra21, Rua23b, Sch21a, Shi22a, Shi23a, Slo24, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22f, Tol22c, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23e, Tol23f, Tol23h, Yeo20, Yeo23]. Press/Elsevier [Tol23e]. Press/Elsevir [Tol22a]. pretty [Jam22c, Ste21b]. price [Wap20]. prime [Con23, Has21b, HLM24]. prime-composite [Con23]. prime-producing [HLM24]. primes [Ayd24, Shi23a]. Princeton [Bay21a, Cri22a, Hew22, Hew23, Hoa20, Hun23b, Hun21d, Hun23c, Lev20a, Lev20b, Lev22e, Lev23d, Lor21c, Lor23d, Mac23b, Pra21, Rua23b, Tol22g, Tol22h, Tol23f]. principle [Jam22b]. Pritchard [Hal20b, Hal22b]. probabilistic [PS22b]. probabilistically [Tol22i]. probabilities [Geo20b]. Probability [Wap22, Bou21, BK21, CK21, Tol22d, Tol22e]. Problem [Ano24d, BM23, Dol20c, L.21a, L.21b, L.22c, L.22a, L.22b, L.23, L.24, Lor20c, Lor20d, Lor23j, Lor23k, Bos21, De22a, Jah22, Kro24, KNS23, Lor22b, Oku23, Sil22, Sur20, Tho22a, Vol22]. Problems [Dol21c, Rob20, Sa23c, Sa23a, Sa23b, Sa24, Woo21, WS22, Woo22a, Woo22b, Fri22, GS21, HS20, Lor22i, Lor23c, Nak22, Bay20a, Hal20a]. processes [Hop22a, Tol22e]. producing [HLM24]. products [BL20, Luk23b]. Profile [Haw21a, Tol21a]. progression [ASH20b, Cer23, OS23, SO23]. promenade [Kob23]. Prometheus [Bay21b, Gri20, Hal21, Tol21b]. Proof [Ber22, Luk22b, Cha24, Dal21, Das22, Dol21a, Gho22a, Gho22b, Has22, Hun22i, Lau23a, Lor24a, Mes23, Ort20, Pel22, Pla22a, Sil24, Tho22c, Cag20, Fre22, Has21a, HL23, Has24, Lev21b, Luk24a, Pla20b, Pla22b, Ste23d, ST22, Lor21c]. Proofs [Ber20, Cha21, Cha22, Haj20, KS23, Mar22, ST23]. properties [KF24, Sch21b, dV21b]. property [RZ21, Ste20a, Wap23]. proportion [Bro22]. prove [Gib22a]. Proving [HP23]. Proximity [Luk21a]. Ptolemy [RZ21, Tho22d]. Publications [Ste22b]. puzzle [Nic20a, Tra24]. puzzles [Haw21b, Tol22c]. PWW [AU22, Ban23, OS22]. Pythagorean [Hun22i, Kro24, Lau23a, Spo21a].

quadrangle [Pla20b]. quadratic [Vil23]. quadratics [Bou21]. quadrature [Mah21c]. quadrilateral [Dal21, Fri22, Lor23l, Mey20, Sci22, VVK21, dV21b]. quadrilaterals [Dal22, Jos20, Jos22]. Quantitative [Hal22c]. quantum [Mac23b]. quartic [MY22b]. quasi [OI23]. quasi-palindromic [OI23]. quaternionic [Cri21c]. Queen [Cri22d, Cri21a]. Quercus [Haw24, Shi22b]. question [Les20, Tol21f]. questionable [Wap22]. quick [Lor22h, KS23]. quintic [Ohy22].

R [Hun23a, Sch21a, Tol22d, Tol22e, Tol23e]. rabbits [Hal22a]. Radek [Hop22a]. radical [Kul21]. radii [Luk22a]. rainbow [PS22b]. Ramanujan [Gad20, Sil24]. Ramsey [Cri22b]. random [Bou21, NO22, Nak22, Tol22e]. range [Sco20]. Ranjan [Luk23b]. Ransome [Haw21b]. ratio [Cus21, Lor24b, Yos22]. rational [Now20, Tho22a, Vu22]. Ratios [Dol20b, Jam21a, Jam21d]. Raymond [Cri22c]. reaction [Hop22a]. Real [Nat22, Bea22b, Bou21, Mel21, NO22, Nat21, Cri21c]. Real-rooted [Nat22, Nat21]. real-valued [NO22]. Reasoning [Ask20, Hal22c]. Rebecca [Pra21]. Received [Ano20a, Ano21a]. Reconciling [Lor21a]. Rectangles [Rid21b]. recurrence [Bea23c, Dol20a]. refinement [Cer23]. regular [Ott24, RV24]. related [GA23, KS23, Les20, Lor22a, Lor23b, Pat22]. Relating [KF24, Sin20, Uhl23]. relation [Gor23b, Luk23d]. relations [Bea23c, Dol20a, SE20b]. relationships [Cag22]. relative [Abe20]. remainders [Gor23a, Lor21a]. remark [Far21, Far23]. remarkable [Tra24]. Remarks [Sha24]. Removing [Fos23]. René [Yeo23, Kit21]. renewal [She23]. repayment [SA23d]. Repeated [BL20]. representation [Ste23e, Vu22]. Representations [Hun22g]. represented [Yos22]. Republic [Hun22h]. result [KF23]. results [KNS23, Mac21a]. Review [Aar22, Bay20b, Bay20a, Bay21a, Bay21b, Bay22, Bay23a, Bay23b, Col24, Cri21b, Cri21c, Cri21a, Cri22a, Cri22b, Cri22d, Cri22c, Cri24, Dav22, Gib22a, Gib22b, Gib23b, Gib24, Gri20, Hal20b, Hal20a, Hal21, Hal22a, Hal22b, Hal22c, Hal23, Haw20, Haw21a, Haw21b, Haw24, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22b, Hun22c, Hun22a, Hun22d, Hun22e, Hun22f, Hun22h, Hun22g, Hun23a, Hun23b, Hun24, Hun20c, Hun21e, Hun21d, Hun23c, Jac20, Jac21, Lev20a, Lev20b, Lev22e, Lev22b, Lev22d, Lev22c, Lev22f, Lev23d, Lev23b, Lev23c, Lev23a, Lev24, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23e, Lor23d, Lor24c, Luk23b, Luk24b, Mac20, Mac23a, Mac24, Mal22, Pra21, Rou20a, Rou20b, Rua23b, Rua23a, Sch20, Sch21a, Shi22b, Shi22a, Shi23a, Shi23b, Shi23c, Shi24]. Review [Slo24, Ste22b, Tol21a, Tol21b, Tol21c, Tol21d, Tol21e, Tol22b, Tol22a, Tol22f, Tol22c, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23e, Tol23f, Tol23a, Tol23g, Tol23h, Tol24, Yeo20, Yeo23, dV21a]. reviewed [Hal22c, Tol22i]. Reviews [Mac23b]. revisited [BK21, CC24, HLM24, Jam22c, Lor22b, Muk23, Sil21a, Ste21b, Tho22a]. Revisiting [LSS23, Jam21b]. revolution [Bel22]. Rewriting [dlR22]. rhythm [Dav22]. Ricardo [Tol22a]. Riemann [Mun20, Pla22b]. right [De22a, Sci22, Nie21]. Right-angled [Nie21]. Rindler [Sch20]. rivals [Dun23]. Rob [Hun21e]. Roberts [Hun22h]. Robinson [Hun22a]. Rods [Mal22]. role [Ott24]. Rolle [Das22]. room [Hal20b]. root [Gup24]. rooted [Nat21, Nat22]. roots [Bou21, Bra23, Nic21, Spo24, Tho22b, Zam20]. rose [Ste23c]. roughly [Hun21e]. Routh [Haj21]. Rowe [Rua23a]. Roy [Luk23b]. Royal [Cri22a]. Royle [Cri22d]. Rudolf [Tol21b]. Rudolphine [Coo21]. rule [Bel22, Hu22]. rules [VVK21]. Rune [dV21a].

S [Bay21b, Bay22, Gri20, Hal21, Hop22a, Lor22d, Mac20]. Saeed [Lor23d]. Samia [Tol22b]. Sangaku [Bos21]. Sautoy [Bay23a]. Scarabotti [Hun21b]. Scheinerman [Tol22f]. Schilling [Yeo23]. school [Hal23]. Schwarz [Far23, Fre22]. science [Wap22, Lev22b, Tol24]. sciences [Mah21a, Mah22b]. Scientific [Tol24]. scientist [Tol21e]. second [Geo20a, Hal23, Lor23e, Luk23b, SR20]. secondary [dlR22]. secret [Tol22g]. sections [Lor24c]. sector [Ste20b]. Sedrakyan [Rou20a, Rou20b]. seemingly [Pra21]. Selvin [Tol21d]. semi [Haj22]. semi-angle-bisectors [Haj22]. Senior [Hal20a]. sense [Row22]. sequences [Jam21d, Joh21, Kac20, Lig23, Luk24b]. series [Ale22, Ber20, Fel20, Jam22c, Lig23, LN22b, Mun20, Sao20, Sch21b, Ste21b, Luk24b, Luk23b]. servant [Cri21a]. SET [Pra21]. Shaun [Jac20]. sheer [Cri22b]. Shores [Mac20]. short [Hun20b, Mar22]. shortcut [Bay23a]. side [Dal21, OS23, SO23]. side-angle [Dal21]. side-lengths [SO23]. sided [Pas21]. sides [ASH20b, Cag21]. significance [McL20]. Silberstein [Hun21b]. similar [Cal20b]. Similarities [Bea23b]. Simon [Shi23c]. simple [Dol21a, Jam23d, Pra21, Ste23e, Ayd24]. Simpler [Kul21]. simplification [LSS23]. simulation [Tol24]. simultaneous [Lor22h]. sine [Gor23a, VVK21]. Singular [Bea24]. six [Gib23a, PdV22]. six-point [PdV22]. sixes [McB21]. sixteenth [Lev24]. slides [CF21]. slower [Bay21a]. slowly [De24]. small [Hun23c, Lor21d]. Smith [Gup24]. Snezana [Shi22a]. Society [Cri22a, Hun22b, Hun22d, Hun22e, Hun22f, Hun23a, Hun24]. solid [Bel22]. Solution [Ohy22, LSS23, Ohy20, OI23]. solutions [Bea22b, Bos21]. solve [MY22b]. solving [All20, Bay20a, Ohy22]. Some [Mah21a, Sil23b, Tra23, Tra24, dV21b, Bur20, Gri21, Lor21b, Lor23c, SA23d, Lor22a, Mac21a, Now20, Sch21b]. Sonar [Gib23b]. Sophie [Hun22e, Hun22f, Lev22d]. Sorbo [Mac23b]. Søren [dV21a]. souls [Kit21]. space [Tol22b]. spaces [Sie20]. Sparks [Lev22c]. special [Pla20a]. specific [Mah20]. speed [Mer23]. spin [DB22]. spin-off [DB22]. spiral [Cal20a, Ste21a]. spirals [Cri20, Rid21b]. sport [Tol21c]. sports [AMMW22]. Spreitzer [Gri20]. Spreizer [Hal21]. Springer [Gib23b, Gib24, Luk24b, Mac20, Mac24, Mal22, Rou20a, Rou20b, Rua23a, Sch20, Shi23b, Shi23c, Tol23a, Tol23g]. Sprößig [Cri21c]. square [Bro22, Cag20, Jam21b, Pla20b, Shi20, Spo24, Tho22b]. square-free [Bro22, Jam21b]. squares [Bea23a, Dol21a, Jam22a, Mac22, ST23, Ste20a]. squaring [Lor23a]. stair [Rou20b]. stair-step [Rou20b]. staircase [Roi23]. standard [Lor23c]. starlight [Lam21]. state [Sie20]. Statistics [Lev23c, Tol21c, Tol23e, Tol21d]. Steiner [GS22, Haj23a, Haj23b]. Steiner-Lehmus [Haj23b]. step [Rou20b]. Stephen [Tol23e]. Steve [Tol21d]. Steven [Bay20b, Hun22d, Sch21a]. Stewart [Bay23b, Jac21, Lor20b, Tol21a]. Stirzaker [Tol22d, Tol22e]. Stochastic [Hop22a]. story [Lor22f, Nic20a]. Strang [Gib22b]. strategies [FL22]. Strick [Gib24]. structure [Nic21, Spo21b]. Student [Dol20c, Dol21c, Rob20, Sa23c, Sa23a, Sa23b, Sa24, Woo21, WS22, Woo22a, Woo22b, Col24]. success [All21]. sufficiency [Dal21]. suggestive [Pau21]. sum [Ayd24, Bau22, Mun20, Nim22, Pla22b]. Sums [Bra23, San23, Bea23a, Cer23, Cha22, Jam22a, Pat22, ST23, Tho22b, Bea23a, Ber22, Ste23b]. surface [Sta20]. surprising [Wap23]. surprisingly [Abe23, Jew24]. Surveying [Bea20]. Susan [Lev22d]. Svetlin [Cri21c]. Switzerland [Gib23b]. Sylvester [Nat21, Nat22]. synthetic [Pel22].

T [Dav22, Lor22e]. tables [Lor23f, Coo21]. tail [NO22]. Talagrand [Mac23b]. tale [De21a, Lor24d, McB21]. Tall [Lor20b]. tangent [Bea24, HL23, Yos22]. tangents [Lau22]. Taschner [Tol21b]. Taylor [Lor23e]. teacher [GT23]. teaching [dlR22]. tears [FP24]. technique [KM21]. Technology [Cri24]. telescope [Muk23]. tennis [CK21, Wap23]. tennis-like [Wap23]. Tessellation [Sta20]. test [SR20]. tetrahedra [Fox20]. Thébault [Vig20]. their [Cag22, Cal20b, LN22b, Ste23a]. themes [Mah21c]. theorem [Awa24, Bea21b, Cag23, Dal21, Dol21b, Dol21a, Dol24, Gîr23, Haj23b, Hun22i, Jam23a, KM21, Lau23a, Lev21a, Luc24, Nat22, PdV22, Pel22, Rid21a, RZ21, Seb22, She23, Shi20, Sil21b, Sin21, Das22, DT24, Haj23a]. theorems [Fal21, KF21a, Kon20, Sil23a, Sil23b, SWT21]. theory [Bay20a, BK21, Mac23b, Mac21a, Row22, Tol22b, Hun21a, Hun22b, Mac23b, Shi24, Tol23a, Luk24b]. there [Pas21]. things [Hun23c]. think [Tol21e]. Thinking [Bay23a, Tol22i, Tol22h]. third [Gib22a]. thirteenth [Sta21a]. Thomas [Gib23b, Hal23, Mac20]. thousand [Tol22d]. three [Has21b, Shi20, LS22]. three-square [Shi20]. Threshold [Bev22]. times [She23]. Tointon [Hun22c]. Tolli [Hun21b]. tool [dlR22]. top [Jah22]. topological [Hun21c]. topology [Jac20, Hil21, Hun20b]. Toscano [Tol22g]. tournament [Wap21]. Toussaint [Dav22]. tractrix [De22b]. transcendence [Shi24]. transform [KM21]. Transformed [Hun23a]. transforms [Col24]. translated [Lev22b]. Trapezia [SO23]. trapezium [Jam21c, OS23]. treasure [Bay21b]. trials [Kac20]. triangle [ASH20a, Ber20, Gor23b, Haj21, Haj22, Hum20, Luk20b, Luk20a, Nar20, Rea22, Sci24, Sco20, Sco22b, Sco23, SS22, Sil23b, Ste24, Tho22e, Whi22]. Triangles [ASH20b, Dol20b, Gri24, Hum20, Nie21, Rea23, Sci22, Sci24, Vu21, Rua23b]. triangular [DL21]. Tricia [Tol21c]. Trigonometric [OS22, Ste23c, Kir20, Ste22c, Hew23]. trigonometrical [Sco22a]. triples [Kro24, Spo21a, Spo21b]. trisect [Ban23]. trisecting [Ban23]. trisectrix [Sil22]. Tropical [Nor20]. trove [Bay21b]. true [Slo24]. Trust [Hal20a]. truth [Bay20b, Tol21e]. truthiness [Tol21e]. Tullio [Hun21b]. ture [Sch21b]. Two [HS20, Jam21d, Lor22i, Mar22, Ber20, De21a, Dol21a, Fri22, KS23, Lor23c, McB21, Abe23, Bos21, Lor23b]. two-squares [Dol21a]. two-variable [Lor23c]. type [Pla20a, Hun22g]. types [Sar23].

Ubiquitous [CF21]. UK [Hal20a, SA20]. uncertain [Wap22]. undergraduates [Shi23c]. underground [Tol23c]. Understanding [Jac20]. unexpected [Bay21b, Wap21]. unfolding [Dol20a]. ungula [Wil23]. Unifying [Kon20]. University [Bay20b, Bay21a, Bay22, Col24, Cri21b, Cri21a, Cri22a, Cri22b, Cri22d, Gib22a, Gib22b, Hal22c, Hal23, Hew22, Hew23, Hoa20, Hop22a, Hun20a, Hun20b, Hun21b, Hun21c, Hun21a, Hun22c, Hun22a, Hun22h, Hun22g, Hun23b, Hun20c, Hun21d, Hun23c, Jac20, Lev20a, Lev20b, Lev22e, Lev22b, Lev22f, Lev23d, Lor20b, Lor21c, Lor22c, Lor22d, Lor22e, Lor23d, Luk23b, Mac23a, Mac23b, Pra21, Rua23b, Sch21a, Shi23a, Slo24, Tol21c, Tol21d, Tol21e, Tol22f, Tol22d, Tol22e, Tol22g, Tol22i, Tol22h, Tol23c, Tol23d, Tol23b, Tol23f, Tol23h, Yeo20, Yeo23, Lor23e]. unsolved [Sur20]. unusual [Kol22, Lor22i, SE20b]. use [Bay23b, Kro24, SWT21]. using [Fre22, Ohy22, SE20a, Spo21a, Vol22].

V [Jac20, Lor23e]. Val [Lev22d]. valued [NO22]. variable [Lor23c]. variables [Lor22d, NO22]. variations [Lor21c, Mac24, Sch20]. Various [Jam23d]. Vasek [Mac23a]. vector [BL20]. Velleman [Gib22a]. velocity [SS20]. Verifying [Mac21d]. Verlag [Gib24, Luk24b, Mac20, Mac24, Mal22, Rou20a, Rou20b, Rua23a, Sch20, Shi23b, Shi23c, Tol23a, Tol23g]. versine [Ste22c]. version [Hal22c]. very [Dol21a, Mah20, Nic20a]. via [HP23, KM21, Luc24, Nak22, RV24]. Victorian [Cri21a]. Viète [Ste20b]. Villani [Lev22b]. Visual [Pri22, Cha24, HP23, Cha21, Cha22]. Visualising [Sar23]. Visually [Pau21]. Vitta [DT24]. Vol [Luk23b]. volume [Ano20g, Ano20f, Ano20i, Ano20h, Ano21f, Ano21e, Ano21h, Ano21g, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23f, Ano23e, Ano23h, Ano23g, Ano23j, Ano23i, Ano24c, Ano24b, Bel22, Tol24].

W [Bay22, Hoa20, Lor22e, Tol23b]. Wainer [Tol21e]. Walk [SS24]. walks [Nak22]. War [Cri22d]. Wasan [Oku23]. way [Mey20, PS22b]. ways [Hun21e, Hun23c, Jam21d]. Weintraub [Hun22d]. Weitzenböck’s [Tra23]. Where [Lor22e]. while [AP21]. Whitehouse [Lev22d]. whose [ASH20b, OS23, SO23]. Wil [Haw21b]. Wilson [Kon20]. Winning [FL22, CK21]. Without [Ber20, HL23, Ber22, Cag20, FP24, Fre22, Has21a, Has22, Has24, Kir20, Lev21b, Luk22b, Luk24a, Pla20b, Pla22b, Ste23d, ST22]. Witt [Cri24]. Wolfgang [Cri21c]. Wolstenholme [Luk23d]. wonder [Lev22f]. Woodward [Hun20a]. Word [Luk22b]. Words [Ber20, Ber22, Cag20, HL23, Pla20b, Fre22, Has21a, Has22, Has24, Luk24a, Pla22b, ST22, Lev21b, Ste23d]. World [Cri22d, Tol24]. Worst [SS20]. writes [Ano23b]. writing [Lev20a, Lev20b, Lev22e, Lev23d]. Wynne [Tol23f].

xx [Gib23b].

Yale [Lor22c, Tol22f]. Yates [Shi22b]. Year [Lev22d, Lev22c, Shi22a]. years [BM23, Gib23b]. young [Hun22d]. youth [Sko20]. Yvinec [Hun21c]. Yvonne [Cri21b].

Zakeri [Lor23d]. Zaslow [Hal22c]. zero [Pla20a]. zero-over-zero [Pla20a]. zeros [Kob20, Mel21]. zeta [Mun20]. zillions [Hoa20]. Zorzitto [Tol23b].

References

Aaronson:2022:BRA

[Aar22] Hugo Aaronson. Book review: Algorithms by Panos Louridas, pp. 312, $15.95 (paper), ISBN 978-0-26253-902-9, MIT Press (2020). The Mathematical Gazette, 106(566):380–381, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

            6328 (electronic). URL https://www.cambridge.org/core/
            journals/mathematical-gazette/article/algorithms-by-
            panos-louridas-pp-312-1595-paper-isbn-9780262539029-
            mit-press-2020/484E5C1645E8FD1F4CFAE799B3E76F5F.

Abel:2020:REB

[Abe20] Ulrich Abel. On the relative error between the binomial and the hypergeometric distribution. The Mathematical Gazette, 104(559):136–142, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

            www.cambridge.org/core/journals/mathematical-gazette/

article/on-the-relative-error-between-the-binomial-and-
the-hypergeometric-distribution/FF9C8D7F86D2CB2E3A570FD3237B470A.

Abel:2023:TDI

[Abe23] Ulrich Abel. 107.35 Two definite integrals that are (not surprisingly) equal. The Mathematical Gazette, 107(570):503–508, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10735-two-
definite-integrals-that-are-not-surprisingly-equal/0C0B33BB30A759C32825C06C1AE55492.

Alegri:2022:IMS

[Ale22] Mateus Alegri. 106.41 Infinitely many series arising from cos 2x + sin 2x = 1. The Mathematical Gazette, 106(567):517– 520, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10641-
infinitely-many-series-arising-from-cos2x-sin2x-1/AE00ACEB8DF80763E272EDAB218ADE61.

Allen:2020:LGA

[All20] Rory Allen. The Lie group approach to solving differential equations. The Mathematical Gazette, 104(559):82–106, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/lie-group-
approach-to-solving-differential-equations/C826E2D74CA15AEC939FC75313DE288B.

Allen:2021:AMS

[All21] Edward J. Allen. Aesop’s moral on success. The Mathematical Gazette, 105(564):481–489, November 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/aesops-moral-on-success/
E82C3934BE749FEF8EF79F916AD85EFD.

Abdin:2022:IBE

[AMMW22] Talaat Abdin, Hosam Mahmoud, Arian Modarres, and Kai Wang. An index for betting with examples from games and sports. The Mathematical Gazette, 106(565):32–40, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/an-index-
for-betting-with-examples-from-games-and-sports/B7C60BC67DA5C2C5E13CEDD0BA041C69.

Anonymous:2020:BR

[Ano20a] Anonymous. Books received. The Mathematical Gazette, 104 (560):382–384, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/books-
received/A238F0651CEC2493307A4044E1E0586B.

Anonymous:2020:F

[Ano20b] Anonymous. Feedback. The Mathematical Gazette, 104(560): 346–352, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/feedback/
C78CFE260B34084344A2C99CD471DB1E.

Anonymous:2020:PJ

[Ano20c] Anonymous. In the pipeline for July 2020. The Mathematical Gazette, 104(559):106, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-july-2020/B3325DE9700BF3D521A256244DCFE3B2.

Anonymous:2020:PM

[Ano20d] Anonymous. In the pipeline for March 2021. The Mathematical Gazette, 104(561):402, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-march-2021/2F526F2FEF7C2B056C96C60A44DEFA68.

Anonymous:2020:PN

[Ano20e] Anonymous. In the pipeline for November 2020. The Mathematical Gazette, 104(560):280, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-november-2020/157BC7FC9EBB37AFD795E2D839A49AE9.

Anonymous:2020:MVIb

[Ano20f] Anonymous. MAG volume 104 issue 559 cover and back matter. The Mathematical Gazette, 104(559):b1–b2, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-104-
issue-559-cover-and-back-matter/10249CCCE52B1B3E7A4C648B39C3F16A.

Anonymous:2020:MVIa

[Ano20g] Anonymous. MAG volume 104 issue 559 cover and front matter. The Mathematical Gazette, 104(559):f1–f2, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-104-
issue-559-cover-and-front-matter/B5DC709C849BEF54DBD42A558A8DEB4B.

Anonymous:2020:MVId

[Ano20h] Anonymous. MAG volume 104 issue 560 cover and back matter. The Mathematical Gazette, 104(560):b1–b2, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-104-
issue-560-cover-and-back-matter/78E8BF0C5E6F7703B6497A2E71CC3E32.

Anonymous:2020:MVIc

[Ano20i] Anonymous. MAG volume 104 issue 560 cover and front matter. The Mathematical Gazette, 104(560):f1–f2, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-104-
issue-560-cover-and-front-matter/164AE04B5650B0B97821F81ACCD78377.

Anonymous:2021:BR

[Ano21a] Anonymous. Books received. The Mathematical Gazette, 105 (562):190–191, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/books-
received/D40408674B23F9D8A33BE5E1BE235EE1.

Anonymous:2021:CGI

[Ano21b] Anonymous. Cumulative Gazette index 1894–2009. The Mathematical Gazette, 105(563):252, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/cumulative-gazette-index-18942009/4FCA0AB8D70372F4F4FEAA7AD38AB9D0.

Anonymous:2021:F

[Ano21c] Anonymous. Feedback. The Mathematical Gazette, 105 (562):163–168, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/feedback/39DBAA3E91F0B432BCA005FF70A93F7B.

Anonymous:2021:PJ

[Ano21d] Anonymous. In the pipeline for July 2021. The Mathematical Gazette, 105(562):97, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-july-2021/C0D05BABD0E6DE27BB6E77C94D0898B3.

Anonymous:2021:MVIb

[Ano21e] Anonymous. MAG volume 105 issue 562 cover and back matter. The Mathematical Gazette, 105(562):b1–b2, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-105-
issue-562-cover-and-back-matter/B8CB5BCC1C3EC5AEB353E2056945DBAA.

Anonymous:2021:MVIa

[Ano21f] Anonymous. MAG volume 105 issue 562 cover and front matter. The Mathematical Gazette, 105(562):f1–f2, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-105-
issue-562-cover-and-front-matter/BE8B22AEB054EAA5C12D92E0F2579F3D.

Anonymous:2021:MVId

[Ano21g] Anonymous. MAG volume 105 issue 563 cover and back matter. The Mathematical Gazette, 105(563):b1–b3, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-105-
issue-563-cover-and-back-matter/F26D0EB2AA84790C8BEE7AE66C12B4A5.

Anonymous:2021:MVIc

[Ano21h] Anonymous. MAG volume 105 issue 563 cover and front matter. The Mathematical Gazette, 105(563):f1–f2, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-105-
issue-563-cover-and-front-matter/3E58915BACE8811A4D4EAF9F4132AEA5.

Anonymous:2022:PM

[Ano22a] Anonymous. In the pipeline for March 2023. The Mathematical Gazette, 106(567):466, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-march-2023/2AAB7D3032A22F4555D6712D2E4919BD.

Anonymous:2022:MPB

[Ano22b] Anonymous. MA Presidential blog. The Mathematical Gazette, 106(565):8, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/ma-presidential-blog/D3BD3492541B419BA66B6BB3673BD8BF.

Anonymous:2022:MVIb

[Ano22c] Anonymous. MAG volume 106 issue 565 cover and back matter. The Mathematical Gazette, 106(565):b1–b3, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-565-cover-and-back-matter/B6F02015407F18330A4D4932ADD7E12C.

Anonymous:2022:MVIa

[Ano22d] Anonymous. MAG volume 106 issue 565 cover and front matter. The Mathematical Gazette, 106(565):f1–f2, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-565-cover-and-front-matter/6B88327B81BF12EED75EB7E94EAC5A69.

Anonymous:2022:MVId

[Ano22e] Anonymous. MAG volume 106 issue 566 cover and back matter. The Mathematical Gazette, 106(566):b1–b3, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-566-cover-and-back-matter/24812E143CBA550CBCA6971DED8AF82D.

Anonymous:2022:MVIc

[Ano22f] Anonymous. MAG volume 106 issue 566 cover and front matter. The Mathematical Gazette, 106(566):f1–f2, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-566-cover-and-front-matter/AB7FC28144016AD615F7EBB61865CCF2.

Anonymous:2022:MVIf

[Ano22g] Anonymous. MAG volume 106 issue 567 cover and back matter. The Mathematical Gazette, 106(567):b1–b3, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-567-cover-and-back-matter/2C8A89329B1B0C644ED5FA8EAC1494B9.

Anonymous:2022:MVIe

[Ano22h] Anonymous. MAG volume 106 issue 567 cover and front matter. The Mathematical Gazette, 106(567):f1–f2, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-106-
issue-567-cover-and-front-matter/CFBD0D2A8D482BC1E8F8F1486D4EF0EA.

Anonymous:2023:A

[Ano23a] Anonymous. Acknowledgements. The Mathematical Gazette, 107(570):575, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/acknowledgements/1059750B99F9C216F43A89BD46783F3E.

Anonymous:2023:EW

[Ano23b] Anonymous. The Editor writes. The Mathematical Gazette, 107(570):545, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/editor-writes/7029875E60D9AF57520C8C18A97B1E20.

Anonymous:2023:PJ

[Ano23c] Anonymous. In the pipeline for July 2023. The Mathematical Gazette, 107(568):159, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-july-2023/7A420B2873436AB3CBCEB6C5387D23EA.

Anonymous:2023:PM

[Ano23d] Anonymous. In the pipeline for March 2024. The Mathematical Gazette, 107(570):544, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-march-2024/FC04EFD046D845725F1AF9963989122E.

Anonymous:2023:MVIb

[Ano23e] Anonymous. MAG volume 107 issue 568 cover and back matter. The Mathematical Gazette, 107(568):b1–b3, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-568-cover-and-back-matter/AE8C47C65946090E25451856CED51C2D.

Anonymous:2023:MVIa

[Ano23f] Anonymous. MAG volume 107 issue 568 cover and front matter. The Mathematical Gazette, 107(568):f1–f2, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-568-cover-and-front-matter/A5801BAD33D0BF4C71064B2B1B09C6C7.

Anonymous:2023:MVId

[Ano23g] Anonymous. MAG volume 107 issue 569 cover and back matter. The Mathematical Gazette, 107(569):b1–b3, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-569-cover-and-back-matter/27C1147CA02428CEB5A9E06C35EFF986.

Anonymous:2023:MVIc

[Ano23h] Anonymous. MAG volume 107 issue 569 cover and front matter. The Mathematical Gazette, 107(569):f1–f2, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-569-cover-and-front-matter/5BB953F07DEED47FA1392B4BE2AB7015.

Anonymous:2023:MVIf

[Ano23i] Anonymous. MAG volume 107 issue 570 cover and back matter. The Mathematical Gazette, 107(570):b1–b3, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-570-cover-and-back-matter/04DD67A944F1EEA3B1664A328BE4B700.

Anonymous:2023:MVIe

[Ano23j] Anonymous. MAG volume 107 issue 570 cover and front matter. The Mathematical Gazette, 107(570):f1–f2, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-107-
issue-570-cover-and-front-matter/E13EC9FE53C3F4D0CC043F1A7F1CA40F.

Anonymous:2024:PJ

[Ano24a] Anonymous. In the pipeline for July 2024. The Mathematical Gazette, 108(571):42, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-july-2024/B26B3595318CD476555D0A7CE806F326.

Anonymous:2024:MVIb

[Ano24b] Anonymous. MAG volume 108 issue 571 cover and back matter. The Mathematical Gazette, 108(571):b1–b3, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-108-
issue-571-cover-and-back-matter/E47C31C7C4A2AD78041C9977DD52F3CB.

Anonymous:2024:MVIa

[Ano24c] Anonymous. MAG volume 108 issue 571 cover and front matter. The Mathematical Gazette, 108(571):f1–f2, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mag-volume-108-
issue-571-cover-and-front-matter/56A64D71AB773F60A75F968D466879F7.

Anonymous:2024:PCF

[Ano24d] Anonymous. Problem corner: Farewell to NJL. The Mathematical Gazette, 108(571):77, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner-farewell-to-njl/9A08A1D28CBC60DF0052DF2F18411771.

On the retirement of Problem Corner editor Nick Lord.

Ash:2021:DWC

[AP21] J. Marshall Ash and Ángel Plaza. ∑_{n=2}^{∞} 1/(nH_{n−1}) diverges while ∑_{n=2}^{∞} 1/(nH_{n−1}^{1+ϵ}) converges. The Mathematical Gazette, 105(562):161–162, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/diverges-while-converges/1DD54F152893D5BA25F52901F0A34988.

Abu-Saymeh:2021:LML

[ASAMHH21] Sadi Abu-Saymeh, Yaqeen Al-Momani, Mowaffaq Hajja, and Mostafa Hayajneh. Long medians and long angle bisectors. The Mathematical Gazette, 105(564):397–409, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/long-
medians-and-long-angle-bisectors/822A56E91FE35463096FFAF746E34394.

Abu-Saymeh:2020:NBP

[ASH20a] Sadi Abu-Saymeh and Mow Affaq Hajja. Notes on the Brocard points and angles of a triangle. The Mathematical Gazette, 104(559):49–62, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/notes-on-the-brocard-points-and-angles-of-a-
triangle/7C971B3A6908834843EC08A4822CB07F.

Abu-Saymeh:2020:TWS

[ASH20b] Sadi Abu-Saymeh and Mowaffaq Hajja. Triangles whose sides form an arithmetic progression. The Mathematical Gazette, 104 (561):469–481, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/triangles-whose-sides-form-an-arithmetic-progression/
0B040986D8F60A451B92513600877D7B.

Askew:2020:RMH

[Ask20] Mike Askew. Reasoning as a mathematical habit of mind. The Mathematical Gazette, 104(559):1–11, March 2020. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/reasoning-as-a-mathematical-habit-of-
mind/A2AD50CFA49800ECF3DD7637D42250EE.

Ahmed:2021:FCO

[AT21] Zafar Ahmed and Pallavi S. Telkar. 105.25 Families of curves orthogonal to the lines y = mx − 2m − m^{3}. The Mathematical Gazette, 105(563):306–309, July 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10525-families-of-curves-orthogonal-
to-the-lines-y-mx-2m-m3/48F51927AE4048E7E493F67B4C84E55B.

Arikan:2022:PSS

[AU22] Elif Esra Arikan and Hasan Unal. 106.38 PWW: sin α + sin β and cos β − cos α. The Mathematical Gazette, 106(567):514, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10638-pww-
sin-sin-and-cos-cos/345B16672F0258AF41BCFC8CC22D07E1.

Awasthi:2024:MLT

[Awa24] Amrit Awasthi. 108.15 The Madhava–Leibniz theorem. The Mathematical Gazette, 108(571):149–151, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10815-the-madhavaleibniz-theorem/725469C50BA3E787FB863D2F05B58241.

Aydin:2024:SBS

[Ayd24] Hazar Aydin. 108.06 Simple bounds on a sum pertinent to primes. The Mathematical Gazette, 108(571):135– 136, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10806-
simple-bounds-on-a-sum-pertinent-to-primes/9A2FCD7DEA53E9DE6FE9A6D6823F847B.

Banks:2023:PTL

[Ban23] Jessica E. Banks. 107.37 PWW: trisecting a line across an angle does not trisect the angle. The Mathematical Gazette, 107(570):512, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10737-pww-trisecting-a-line-across-an-angle-
does-not-trisect-the-angle/40BA7A381D54738D4558F720DD86780F.

Baumslag:2021:IDL

[Bau21] Benjamin Baumslag. 105.08 An infinitesimal definition of limit. The Mathematical Gazette, 105(562):126–129, March 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10508-an-infinitesimal-
definition-of-limit/6717E831A3423282BEC61551AA4941DF.

Baumann:2022:IDS

[Bau22] Michael Heinrich Baumann. An illustrative derivation of the sum of fifth powers. The Mathematical Gazette, 106(565):68–77, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/an-illustrative-
derivation-of-the-sum-of-fifth-powers/27C76DA68709566213C21D688FFE5A16.

Baylis:2020:BRM

[Bay20a] John Baylis. Book review: Methods of solving number theory problems by Ellina Grigorieva, pp. 391, £37.99 (hard), ISBN 978-3-319-90914-1, Birkhäuser (2018). The Mathematical Gazette, 104(560):378–380, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/methods-of-solving-number-theory-problems-by-
ellina-grigorieva-pp-391-3799-hard-isbn-9783319909141-
birkhauser-2018/A73A14B2C9A892111A3A8E581BD053EC.

Baylis:2020:BRE

[Bay20b] John Baylis. Book review: The error of truth by Steven J. Osterlind, pp. 352, £25 (hard), ISBN 978-0-19883-160-0, Oxford University Press (2019). The Mathematical Gazette, 104(560):364–366, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/error-of-truth-by-steven-j-osterlind-pp-352-
25-hard-isbn-9780198831600-oxford-university-press-2019/
26313470E4EDFBCEB79DBE54BB927D2D.

Baylis:2021:BRH

[Bay21a] John Baylis. Book review: How to fall slower than gravity by Paul J. Nahin, pp. 279, £22.00 (hard), ISBN 978-0-691-17691-8, Princeton University Press (2018). The Mathematical Gazette, 105(563):377–378, July 2021. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/how-to-fall-slower-than-gravity-by-paul-j-nahin-
pp-279-2200-hard-isbn-9780691176918-princeton-university-
press-2018/1020561D2854EA75D6CEAEA6B38A2F58.

Baylis:2021:BRM

[Bay21b] John Baylis. Book review: Mathematical Curiosities: a treasure trove of unexpected entertainments by Alfred S. Posamentier and Ingmar Lehmann, pp 382, £14.95 (paper), ISBN 978-1-61614-931-4, Prometheus Books (2014). Also available as e-book. The Mathematical Gazette, 105 (563):382–383, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-curiosities-a-treasure-trove-of-
unexpected-entertainments-by-alfred-s-posamentier-and-
ingmar-lehmann-pp-382-1495-paper-isbn-9781616149314-
prometheus-books-2014-also-available-as-ebook/E84CDD636AA729DD42CED3D205BBC20A.

Baylis:2022:BRL

[Bay22] John Baylis. Book review: Linear algebra by Elizabeth S. Meckes and Mark W. Meckes, pp. 427, £49.99 (hard), ISBN 978-1-107-17790-1, Cambridge University Press (2018). The Mathematical Gazette, 106(565):174–175, March 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/linear-algebra-by-elizabeth-s-meckes-
and-mark-w-meckes-pp-427-4999-hard-isbn-9781107177901-
cambridge-university-press-2018/869D64950E9F49E7D178EA58D4A48860.

Baylis:2023:BRT

[Bay23a] John Baylis. Book review: Thinking better: the art of the shortcut in math and life by Marcus du Sautoy, pp 336, £20, ISBN 978-1-5416-0036-2, Basic Books (2021). The Mathematical Gazette, 107(570):571–573, November 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/thinking-better-the-art-
of-the-shortcut-in-math-and-life-by-marcus-du-sautoy-
pp-336-20-isbn-9781541600362-basic-books-2021/51694B25DD3F170E0A8840239BCE5D1B.

Baylis:2023:BRW

[Bay23b] John Baylis. Book review: What’s the use? by Ian Stewart, pp 336, £20 (hard), ISBN 978-1-5416-9949-6, Basic Books (2021). The Mathematical Gazette, 107(570):571–573, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/whats-the-
use-by-ian-stewart-pp-336-20-hard-isbn-9781541699496-
basic-books-2021/37BEE74EA880713F7BFF30D5EA817271.

Brown:2020:EIE

[BB20] Geoffrey Brown and Narayanaswamy Balakrishnan. Extensions of an identity of Euler. The Mathematical Gazette, 104(560):241–246, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/extensions-of-an-identity-of-euler/30B79A84247855CD0A6088167180C04E.

Beardon:2020:SHP

[Bea20] A. F. Beardon. 104.24 Surveying in the hyperbolic plane. The Mathematical Gazette, 104(560):341–343, July 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10424-surveying-in-the-
hyperbolic-plane/7373B904A0C80E528BA081EBDCE7AFDB.

Beardon:2021:AAB

[Bea21a] A. F. Beardon. 105.26 Areas above and below a curve. The Mathematical Gazette, 105(563):309–311, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10526-areas-above-and-below-a-curve/
85D746B4260E9CAB4EC0B61B43BFE324.

Beardon:2021:PTD

[Bea21b] A. F. Beardon. Pitot’s theorem, dynamic geometry and conics. The Mathematical Gazette, 105(562):52–60, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/pitots-theorem-dynamic-
geometry-and-conics/3E40A70B857EEEAAC5A6FDBAB4BF419B.

Beardon:2022:FNC

[Bea22a] A. F. Beardon. 106.33 Fibonacci numbers and Cassini’s identity. The Mathematical Gazette, 106(567):498–501, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10633-fibonacci-numbers-
and-cassinis-identity/5E7A4CD24EDA06A769F51A4989BCAEA8.

Beardon:2022:RSX

[Bea22b] A. F. Beardon. The real solutions of x = a^{x}. The Mathematical Gazette, 106(566):206–211, July 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/real-solutions-of-x-ax/
77C83CF28975D2E11087876356C4C1CF.

Beardon:2022:CAD

[Bea22c] Alan Beardon. On 105.28 Correct answer — dodgy method. The Mathematical Gazette, 106(566):351, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/on-10528/43A88D60865884743E551C82F752B3AD.

Beardon:2022:X

[Bea22d] Alan Beardon. On 106.17. The Mathematical Gazette, 106 (567):550–551, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10617/71C944D047ADAEB4FB76482DB0CC241F.

Beardon:2023:SSS

[Bea23a] A. F. Beardon. 107.28 Sums, and sums of squares. The Mathematical Gazette, 107(570):488–490, November 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10728-sums-and-sums-of-
squares/FB8494020D2723DCB8BC724D3CD99464.

Beardon:2023:SCP

[Bea23b] A. F. Beardon. 107.36 Similarities and circle-preserving bijections of the plane. The Mathematical Gazette, 107(570):508– 511, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10736-
similarities-and-circlepreserving-bijections-of-the-
plane/458BFCB7B44D6E5ED7534CE552F9B52A.

Beardon:2023:GCR

[Bea23c] A. F. Beardon. Groups, conics and recurrence relations. The Mathematical Gazette, 107(569):193–203, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/groups-conics-and-recurrence-relations/
FFD697ECD6A2A172876A45887EE33107.

Beardon:2024:SMP

[Bea24] A. F. Beardon. Singular matrices and pairwise-tangent circles. The Mathematical Gazette, 108(571):12–19, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/singular-matrices-and-pairwisetangent-
circles/8875EA6401BA48765C3CBD3F4FFA788C.

Belcher:2022:AVS

[Bel22] Paul Belcher. Approximating the volume of a solid of revolution — the frustum rule. The Mathematical Gazette, 106 (567):408–413, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/approximating-the-volume-of-a-solid-of-revolution-
the-frustum-rule/1F3F9DBE6850CEB1B87E8970FCF66D29.

Berendonk:2020:ASE

[Ber20] Stephan Berendonk. 104.10 Alternating series in an equilateral triangle — two proofs without words. The Mathematical Gazette, 104(559):170–171, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10410-alternating-series-in-an-equilateral-triangle-
two-proofs-without-words/A70D4A6E7621AFC9BC884F92D806C49C.

Berendonk:2022:SHN

[Ber22] Stephan Berendonk. 106.16 Sums of hex numbers are cubes — a planar proof without words. The Mathematical Gazette, 106 (565):147–148, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10616-
sums-of-hex-numbers-are-cubes-a-planar-proof-without-
words/F1CD6B2BAA68B746FD8BB3FAB3A6F9A5.

Bevan:2022:TFB

[Bev22] David Bevan. 106.30 Threshold functions and the birthday paradox. The Mathematical Gazette, 106(566):344– 348, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10630-
threshold-functions-and-the-birthday-paradox/3EB9A50E261D857CD590763F3F49C8B0.

Beardon:2020:CFD

[BG20] Alan F. Beardon and Russell A. Gordon. The convexity of the function y = E(x) defined by x^{y} = y^{x}. The Mathematical Gazette, 104(559):36–43, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/convexity-of-the-function-y-ex-defined-by-xy-
yx/65AE9DCC0ED961A8E5738AE62E4DD4F9.

Bougoffa:2021:IIP

[BK21] Lazhar Bougoffa and Panagiotis T. Krasopoulos. Integral inequalities in probability theory revisited. The Mathematical Gazette, 105(563):263–270, July 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/integral-inequalities-in-probability-
theory-revisited/47C44654F595CB749E142E883E357061.

Beardon:2020:RVP

[BL20] A. F. Beardon and N. Lord. Repeated vector products. The Mathematical Gazette, 104(561):460–468, November 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/repeated-vector-products/
36476DE71FA170862ADD6976130B7006.

Bardelang:2023:LPY

[BM23] Reinfried Bardelang and Bernard Murphy. 107.41 Langley’s problem: 100 years on. The Mathematical Gazette, 107(570): 523–526, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10741-
langleys-problem-100-years-on/2C224730F9AFD6BD83A573AE08BA3782.

Bosch:2021:TSS

[Bos21] Robert Bosch. 105.12 Two solutions to a Sangaku problem. The Mathematical Gazette, 105(562):139–142, March 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10512-two-solutions-to-
a-sangaku-problem/BBC17C86A45D00A6A3F26F4599001F56.

Boucher:2021:PCR

[Bou21] Chris Boucher. The probability certain random quadratics have real roots. The Mathematical Gazette, 105(564): 410–415, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/probability-certain-random-quadratics-have-real-
roots/4304E5C2379129D6C0033D43F06A17F2.

Braza:2023:SR

[Bra23] Peter Braza. Sums of roots. The Mathematical Gazette, 107(570): 399–404, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/sums-
of-roots/434C4B38E020A392C3D387489FB4D6FD.

Brown:2022:WPS

[Bro22] Ron Brown. 106.31 What proportion of square-free numbers are divisible by 2? or by 30 but not by 7? The Mathematical Gazette, 106(567):494–497, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10631-what-proportion-of-squarefree-numbers-
are-divisible-by-2-or-by-30-but-not-by-7/C297E82493024EFA7ADD0F10A9DE135F.

Buhmann:2023:MG

[Buh23] Martin Buhmann. The Mathematikum in Giessen. The Mathematical Gazette, 107(568):1–9, March 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mathematikum-in-giessen/
A5EFD99CCFC27B6523D1EBC1EEFE2D39.

Buritica:2020:ESP

[Bur20] Andres Buritica. Equations of some plane figures. The Mathematical Gazette, 104(560):288–295, July 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/equations-of-some-plane-
figures/0EF29AF269B5892724D4B73B5742CC99.

Burn:2022:FJ

[Bur22] Bob Burn. On feedback July 2022. The Mathematical Gazette, 106(567):552, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-feedback-july-2022/9E6A91DD5BCB3854F66A6ECEF4F1A89D.

Caglayan:2020:PWE

[Cag20] Günhan Caglayan. 104.25 Proof without words: Every central octagonal number is an odd square. The Mathematical Gazette, 104(560):343, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10425-
proof-without-words-every-central-octagonal-number-is-
an-odd-square/094467300306B11F004946F3374AD23A.

Caglayan:2021:DBG

[Cag21] Günhan Caglayan. 105.34 The difference between k-gonal numbers with (2n − 1) sides and (n − 1) sides is (3k − 4)-gonal. The Mathematical Gazette, 105(563):333–334, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10534-the-difference-between-kgonal-
numbers-with-2n-1-sides-and-n-1-sides-is-3k-4gonal/F50D249733BC603FC18C8C4783D1842E.

Caglayan:2022:PNT

[Cag22] Günhan Caglayan. 106.32 Pentagonal numbers and their relationships to other figurate numbers. The Mathematical Gazette, 106(567):497–498, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10632-pentagonal-numbers-and-their-relationships-
to-other-figurate-numbers/459B318E347214EF47D54BCF9AF34757.

Caglayan:2023:DTI

[Cag23] Günhan Caglayan. 107.26 A difference theorem involving kgonal and centred k-gonal numbers. The Mathematical Gazette, 107(569):342–343, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10726-a-difference-theorem-involving-kgonal-
and-centred-kgonal-numbers/05FE50333BADDC4CB68C4EB21517999B.

Calleja:2020:SPI

[Cal20a] James Calleja. A spiral pattern investigation: making mathematical connections. The Mathematical Gazette, 104(560):262–270, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/spiral-pattern-
investigation-making-mathematical-connections/C367658871A92F8D8266EB1FC04CE188.

Calugareanu:2020:MST

[Cal20b] Grigore Calugareanu. Matrices that are similar to their inverses. The Mathematical Gazette, 104(559):116–124, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/matrices-that-
are-similar-to-their-inverses/EA4EE844A828FAAA16ADC035862C8D2A.

Chakraborty:2024:ELR

[CC24] Bikash Chakraborty and Sagar Chakraborty. 108.11 Euler’s limit-revisited. The Mathematical Gazette, 108(571): 144–145, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10811-
eulers-limitrevisited/DCCA87210D5A775CC280D2F075CADA2D.

Cereceda:2023:RGF

[Cer23] José Luis Cereceda. 107.25 A refinement of Griffiths’ formula for the sums of the powers of an arithmetic progression. The Mathematical Gazette, 107(569):340–342, July 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10725-a-refinement-of-griffiths-formula-
for-the-sums-of-the-powers-of-an-arithmetic-progression/
94955F91C5C0B2C096FB1003DA7E999B.

Crilly:2021:UCP

[CF21] Tony Crilly and Colin R. Fletcher. Ubiquitous cousins and parental slides. The Mathematical Gazette, 105(562):27–39, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/ubiquitous-cousins-
and-parental-slides/DFE016C05EB5D2A2E73EBA9ED6DF9C4E.

Chakraborty:2021:VPX

[Cha21] Manishita Chakraborty. 105.23 Visual proofs of x^{3} + y^{3} = (x + y)(x^{2}−xy+y^{2}) and x^{3}−y^{3} = (x−y)(x^{2}+xy+y^{2}). The Mathematical Gazette, 105(563):303–305, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10523-visual-proofs-of-x3-y3-x-yx2-xy-y2-and-
x3-y3-x-yx2-xy-y2/4F14C92D94A0745C8BD030D695CED351.

Chakraborty:2022:VPS

[Cha22] Manishita Chakraborty. 106.39 Visual proofs of sums of powers of positive integers. The Mathematical Gazette, 106(567): 515–516, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10639-
visual-proofs-of-sums-of-powers-of-positive-integers/
8D526BF5D39CF20C54CD2E3D9348FEF1.

Chakraborty:2024:VPW

[Cha24] Bikash Chakraborty. 108.09 A visual proof that b^{e} < e^{b} when b > e. The Mathematical Gazette, 108(571):143, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10809-a-visual-
proof-that-be-eb-when-b-e/C77F8DAD439C56EF26C2B9777D5EACC3.

Chebolu:2023:CMP

[Che23]  Sunil K. Chebolu.  A clock model for planetary con-

junctions. The Mathematical Gazette, 107(570):422–429, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/clock-
model-for-planetary-conjunctions/90E3DE49B5B006246E11C88964B83EC1.

Corless:2020:IFC

[CJS20] Robert M. Corless, David J. Jeffrey, and David R. Stoutemyer. Integrals of functions containing parameters. The Mathematical Gazette, 104(561):412–426, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/integrals-of-functions-containing-parameters/
0D5DEC55DF2CDB7A1AD29FFA029EC782.

Cooper:2021:ELP

[CK21] Curtis Cooper and Robert E. Kennedy. Expected length and probability of winning a tennis game. The Mathematical Gazette, 105(564):490–500, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/expected-length-and-probability-of-winning-a-
tennis-game/0F0397613253368C041889822322EF1D.

Colwell:2024:BRS

[Col24] Sue Colwell. Book review: A student’s guide to Laplace transforms by Daniel Fleisch, pp. 218, £17.99, (paper), ISBN 978-1-009-09629-4, Cambridge University Press (2022). The Mathematical Gazette, 108(571):183–184, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/students-guide-to-laplace-transforms-
by-daniel-fleisch-pp-218-1799-paper-isbn-9781009096294-
cambridge-university-press-2022/7DCE4B95040C0CD2E14CD4DC4496CE99.

Contensou:2023:EPC

[Con23] Matthieu Contensou. 107.16 An expression for the primecomposite characteristic function. The Mathematical Gazette, 107(569):306–307, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10716-an-expression-for-the-primecomposite-characteristic-
function/E292261B33483AD779CDA809E3CE8378.

Cooker:2021:JKH

[Coo21] Mark J. Cooker. Johannes Kepler and his making of the Rudolphine Tables. The Mathematical Gazette, 105(564):425–432, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/johannes-
kepler-and-his-making-of-the-rudolphine-tables/9104D317FB668DB4ACD61497EB9DB9B6.

Crilly:2020:FDS

[Cri20] Tony Crilly. A family of discrete spirals. The Mathematical Gazette, 104(560):215–224, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/family-of-discrete-spirals/4D574D2B84F4AB3621BD52E7C9399D0A.

Crilly:2021:BRS

[Cri21a] Tony Crilly. Book review: From servant to queen: a journey through Victorian mathematics by John Heard, pp. 267, £34.99 (paper), ISBN 978-107-12413-4, Cambridge University Press (2019). The Mathematical Gazette, 105(562):178– 181, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/from-
servant-to-queen-a-journey-through-victorian-mathematics-
by-john-heard-pp-267-3499-paper-isbn-978107124134-cambridge-
university-press-2019/02431A19E0B7026F00D5DC8F176FA715.

Crilly:2021:BRN

[Cri21b] Tony Crilly. Book review: New light on George Boole by Desmond MacHale and Yvonne Cohen, pp. 476, £17.95 (paper), ISBN 978-1-78205-290-6, Atrium (Cork University Press) (2018). The Mathematical Gazette, 105(562):181–184, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/new-light-
on-george-boole-by-desmond-machale-and-yvonne-cohen-
pp-476-1795-paper-isbn-9781782052906-atrium-cork-university-
press-2018/4EDC2DA7B674F60BCCD3020A9B5F2FBE.

Crilly:2021:BRR

[Cri21c] Tony Crilly. Book review: Real quaternionic calculus handbook by João Pedro Morais, Svetlin Georgiev, Wolfgang Sprößig, p. 216, £54.99, ISBN 978-3-0348-0621-3, Birkhäuser (2014). The Mathematical Gazette, 105(563):370–371, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/real-quaternionic-calculus-handbook-

by-joao-pedro-morais-svetlin-georgiev-wolfgang-sprossig-
p-216-5499-isbn-9783034806213-birkhauser-2014/44E31E530705F87CF628847500800D71.

Crilly:2022:BRA

[Cri22a] Tony Crilly. Book review: Arts & minds: how the Royal Society of Arts changed a nation by Anton Howes, pp. 387, £30 (hard), ISBN: 978-0-691-18264-3, Princeton University Press (2020). The Mathematical Gazette, 106(566):362–363, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/arts-minds-how-the-royal-society-of-
arts-changed-a-nation-by-anton-howes-pp-387-30-hard-
isbn-9780691182643-princeton-university-press-2020/E8EBFC94A1BBA2CC5730B7F789C9A1EE.

Crilly:2022:BRFa

[Cri22b] Tony Crilly. Book review: Frank Ramsey: a sheer excess of powers by Cheryl Misak, pp. 537, £25 (hard), ISBN 978-0-19-875535-7, Oxford University Press (2020). The Mathematical Gazette, 106(566):363–366, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/frank-ramsey-a-sheer-excess-of-powers-by-cheryl-
misak-pp-537-25-hard-isbn-9780198755357-oxford-university-
press-2020/9AEF2058EFB68D1ED6E6985528B2BA71.

Crilly:2022:BRM

[Cri22c] Tony Crilly. Book review: Mathematics at the meridian: the history of mathematics at Greenwich edited by Raymond Flood, Tony Mann and Mary Croarken, pp 241, £20.40 (paper), ISBN 978-0-367-36272-0, Chapman & Hall/CRC (2020). The Mathematical Gazette, 106(565):170–172, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mathematics-at-
the-meridian-the-history-of-mathematics-at-greenwich-
edited-by-raymond-flood-tony-mann-and-mary-croarken-
pp-241-2040-paper-isbn-9780367362720-chapman-hallcrc-
2020/84A1D13FCB3921C02FCCA353250AC8F7.

Crilly:2022:BRFb

[Cri22d] Tony Crilly. Book review: The flying mathematicians of World War I by Tony Royle, pp. 269, £22.50 (paper), ISBN 978-0-2280-0373-1, McGill-Queen’s University Press (2020). The Mathematical Gazette, 106(566):367–370, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/flying-mathematicians-of-world-war-i-by-tony-
royle-pp-269-2250-paper-isbn-9780228003731-mcgillqueens-
university-press-2020/6E76CC324C9F845BF7AF34C6E3CAF07C.

Crilly:2024:BRF

[Cri24] Tony Crilly. Book review: Formulations: architecture, mathematics, culture by Andrew Witt, pp. 428, £23.15, (paper), ISBN 978-0-262-54300-2, Massachusetts Institute of Technology Press (2021). The Mathematical Gazette, 108 (571):177–179, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/formulations-architecture-mathematics-culture-
by-andrew-witt-pp-428-2315-paper-isbn-9780262543002-
massachusetts-institute-of-technology-press-2021/F9730564744AFAD00BECCC51D722D9D3.

Cusmariu:2021:GRE

[Cus21] Adolf Cusmariu. 105.07 A golden ratio enigma. The Mathematical Gazette, 105(562):125–126, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10507-a-golden-ratio-enigma/137F3706EA4CD3FCE1D8A974F6C6D7BB.

Dalcin:2021:SAD

[Dal21] Mario Dalcín. 105.32 The side-angle duality in geometry: a direct proof of sufficiency of a cyclic quadrilateral theorem. The Mathematical Gazette, 105(563):329– 331, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10532-
the-sideangle-duality-in-geometry-a-direct-proof-of-
sufficiency-of-a-cyclic-quadrilateral-theorem/A7348E8115E91D39B187DD3CE6DC40AC.

Dalcin:2022:NDC

[Dal22] Mario Dalcín. New dualities in convex quadrilaterals. The Mathematical Gazette, 106(566):269–280, July 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/new-dualities-in-convex-
quadrilaterals/938CBB11FF40078858B15C7AE3F0299B.

Das:2022:APR

[Das22] Himadri Lal Das. 106.42 Another proof of Rolle’s Theorem. The Mathematical Gazette, 106(567):521–522, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10642-another-proof-of-
rolles-theorem/AC2D53EB9DF14C31F778992D526ACD1A.

David:2022:BRG

[Dav22] Hopkins David. Book review: The geometry of musical rhythm: what makes a “good” rhythm good? (2nd edn.) by Godfried T. Toussaint, pp. 352, £37.15 (paperback), ISBN 978-0-8153-7097-0, CRC Press (2020). The Mathematical Gazette, 106(565):187– 188, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/geometry-
of-musical-rhythm-what-makes-a-good-rhythm-good-2nd-
edn-by-godfried-t-toussaint-pp-352-3715-paperback-isbn-
97808153-70970-crc-press-2020/2DCBB138DDE32C52AF5308B96068FE70.

De:2022:IS

[DB22] Prithwijit De and Sutanay Bhattacharya. 106.17 An interesting spin-off. The Mathematical Gazette, 106(566):310–312, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10617-an-
interesting-spinoff/D128B886B256524FAB428D75895B5EBD.

De:2021:TTC

[De21a] Prithwijit De. 105.40 A tale of two cubics. The Mathematical Gazette, 105(564):514–516, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10540-a-tale-of-two-cubics/9ACFBA64A08DC681C9D591664D31F095.

De:2021:LM

[De21b] Prithwijit De. Learning from a mistake. The Mathematical Gazette, 105(563):349, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/learning-from-a-mistake/2F8FF9E0574B9624ACB79632093A06BA.

De:2022:OPI

[De22a] Prithwijit De. 106.06 an optimisation problem involving right circular cones. The Mathematical Gazette, 106(565):127–130, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10606-an-
optimisation-problem-involving-right-circular-cones/
016F0FE7AD39192996DF48E93A0DB22D.

De:2022:IMT

[De22b] Subhranil De. The intriguing mechanics of a tractrix of cards. The Mathematical Gazette, 106(566):281–290, July 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/intriguing-mechanics-of-
a-tractrix-of-cards/603C4478340E3765330DC8C8A107D0E4.

De:2024:SEC

[De24] Subhranil De. A slowly evolving conical pendulum. The Mathematical Gazette, 108(571):105–110, March 2024. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/slowly-evolving-conical-
pendulum/E34D4DC3FA5D2FCF3319BEDFD659A31B.

De:2021:TE

[DL21] Prithwijit De and Gerry Leversha. A triangular exploration. The Mathematical Gazette, 105(564):501–506, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/triangular-exploration/
783E039E63C56863C156C0B1A07FFBAC.

delaRosa:2022:RPT

[dlR22] Félix Martínez de la Rosa. Rewriting polynomials: a tool for teaching secondary mathematics. The Mathematical Gazette, 106(567):544–547, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/rewriting-polynomials-a-tool-for-teaching-secondary-
mathematics/1B10CE4622D2F0564C823B9A0638203C.

Dolan:2020:GUR

[Dol20a] Stan Dolan. The geometric unfolding of recurrence relations. The Mathematical Gazette, 104(561):403–411, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/geometric-unfolding-of-
recurrence-relations/11C2974CAFDA3E476DEBB21EAD8B4092.

Dolan:2020:RHT

[Dol20b] Stan Dolan. Ratios in Heronian triangles. The Mathematical Gazette, 104(560):193–208, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/ratios-in-heronian-triangles/09561344BC7C9068244414E2DB38202D.

Dolan:2020:SPC

[Dol20c] Stan Dolan. Student problem corner. The Mathematical Gazette, 104(560):359–361, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problem-corner/DA40E5437CCC1648ADCFCDE48639FB0F.

Dolan:2021:VSP

[Dol21a] Stan Dolan. 105.38 A very simple proof of the twosquares theorem. The Mathematical Gazette, 105(564):511, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10538-
a-very-simple-proof-of-the-twosquares-theorem/D0CB1CB39CBA0E98905401EA21DCB743.

Dolan:2021:ETG

[Dol21b] Stan Dolan. 105.39 The Eureka theorem of Gauss. The Mathematical Gazette, 105(564):512–514, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10539-the-eureka-theorem-of-gauss/B85583C9C2DECC7D7F479B50E08869C0.

Dolan:2021:SP

[Dol21c] Stan Dolan. Student problems. The Mathematical Gazette, 105(562):175–177, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/309C7F79F12D4CCDCBA45E0C9FFBFB16.

Dolan:2024:ETG

[Dol24] Stan Dolan. The Eureka theorem of Gauss. The Mathematical Gazette, 108(571):94–104, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/eureka-theorem-of-gauss/AF396D6605383A7B9CD7E7CF537F80CB.

Downes:2020:EGJ

[Dow20] Rob Downes. 104.13 An elementary geometric justification for the cofactor expansion of a 3 × 3 determinant. The Mathematical Gazette, 104(560):307–310, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10413-an-elementary-geometric-justification-
for-the-cofactor-expansion-of-a-3-3-determinant/E058B03F4832DDECD03AE1DBE8CEE21E.

Dergiades:2023:IGA

[DT23] Nikolaos Dergiades and Quang Hung Tran. 107.13 An interesting generator of Archimedean circles. The Mathematical Gazette, 107 (568):155–159, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10713-
an-interesting-generator-of-archimedean-circles/2A35D9A6D2581D579CD69B550C028FE7.

Dergiades:2024:EVT

[DT24] Nikolaos Dergiades and Quang Hung Tran. Extensions of Vittas’ Theorem. The Mathematical Gazette, 108(571):53–60, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/extensions-of-
vittas-theorem/6BA822031AAE71DE0B79F7EF5F15396E.

Dubeau:2022:APC

[Dub22] François Dubeau. Archimedes playing with a computer. The Mathematical Gazette, 106(567):427–442, November 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/archimedes-playing-with-
a-computer/A2EC0A2B75DC65A749F39BFCFDDDBCBA.

Dunham:2023:CHM

[Dun23] William Dunham. Cauchy and his modern rivals. The Mathematical Gazette, 107(568):103–113, March 2023. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/cauchy-and-his-modern-rivals/541D4E39677E2A65C1E98119780937DE.

deVilliers:2021:BRI

[dV21a] Michael de Villiers. Book review: Introduction to experimental mathematics by Søren Eilers and Rune Johansen, pp. 303, £34.99 (hard), ISBN 978-1-107-15613-5, (2017). The Mathematical Gazette, 105(563):379–380, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/introduction-to-experimental-mathematics-by-
soren-eilers-and-rune-johansen-pp-303-3499-hard-isbn-
9781107156135-2017/C3FEFEBE3D8F72F79CDA89C559D15B2B.

deVilliers:2021:SMP

[dV21b] Michael de Villiers. Some more properties of the bisectdiagonal quadrilateral. The Mathematical Gazette, 105(564): 474–480, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/some-
more-properties-of-the-bisectdiagonal-quadrilateral/
C772FC0EB68F716883C2649C67A1B837.

Emamyari:2023:AMF

[EH23] Soheila Emamyari and Mehdi Hassani. 107.33 On the antiderivatives of a monotone function and its inverse. The Mathematical Gazette, 107(570):499–501, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10733-on-the-antiderivatives-of-a-monotone-function-
and-its-inverse/6F6ECA16127F2608D55218460C4938B6.

Falbo:2021:AGT

[Fal21] Clement E. Falbo. 105.03 Almost Goldbach theorems. The Mathematical Gazette, 105(562):111–116, March 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10503-almost-goldbach-theorems/5D6B07369AE490E2842F2A681ABEB6BB.

Farhadian:2021:R

[Far21] Reza Farhadian. 105.27 A remark on lim_{n→∞} ^{pn} \sqrt{p_{1}p_{2} ⋯ p_{n}} = e. The Mathematical Gazette, 105(563):311–312, July 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10527-a-remark-on-mathop-
lim-limitsn-to-infty-sqrtpnp1p2pn-e/FFEEDD3FD1AA3C49932C63D26F991283.

Farhadian:2023:RCS

[Far23] Reza Farhadian. 107.30 Remark on Cauchy–Schwarz inequality. The Mathematical Gazette, 107(570):493–495, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10730-remark-on-cauchyschwarz-
inequality/41BCBEE773E77085D59436AD0B629870.

Fellner:2020:DMH

[Fel20] Joseph P. Fellner. 104.15 Divergence of a modified harmonic series. The Mathematical Gazette, 104(560):313– 315, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10415-
divergence-of-a-modified-harmonic-series/54DF18FE1F0E413FE44866BCEE07289D.

Friedman:2022:WSE

[FL22] Eric J. Friedman and Adam S. Landsberg. Winning strategies: the emergence of base 2 in the game of nim. The Mathematical Gazette, 106(566):212–219, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/winning-strategies-the-emergence-of-
base-2-in-the-game-of-nim/569B246E46C6985A15CDFF91006FC904.

Foster:2023:LMI

[Fos23] Colin Foster. Less is more: Improving by removing: (the 2023 Presidential Address). The Mathematical Gazette, 107(570): 385–398, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/less-
is-more-improving-by-removing/14A05E52969D058A7378326D13777A43.

Fox:2020:CTG

[Fox20] Michael Fox. Constructing tetrahedra with given face areas. The Mathematical Gazette, 104(559):63–73, March 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/

mathematical-gazette/article/constructing-tetrahedra-
with-given-face-areas/4332700B63745B4FBA9806D78AB272D7.

Farhadian:2024:IET

[FP24] Reza Farhadian and Vadim Ponomarenko. 108.13 Indeterminate exponentials without tears. The Mathematical Gazette, 108 (571):146–148, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10813-
indeterminate-exponentials-without-tears/377915A9D2AB8FD54F8CCE1F99495EF4.

Frederickson:2021:HDP

[Fre21] Greg N. Frederickson. Hole dissections for planar figures. The Mathematical Gazette, 105(563):237–244, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/hole-dissections-for-planar-figures/
664D32860E24AD2E16434DD1EA5C7E7A.

FreitasGregorio:2022:PWC

[Fre22] Edney Freitas Gregorio. 106.08 Proof without words: the Cauchy–Schwarz inequality using analytic geometry. The Mathematical Gazette, 106(565):132, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10608-proof-without-words-the-cauchyschwarz-
inequality-using-analytic-geometry/1346D5FC0D670C00263038E1CA7924EC.

Fried:2022:LPC

[Fri22] Michael N. Fried. Locus problems concerning centroids of a cyclic quadrilateral and two classic cubic curves. The Mathematical Gazette, 106(566):247–257, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/locus-problems-concerning-centroids-
of-a-cyclic-quadrilateral-and-two-classic-cubic-curves/
A8D93DB53CDD479A964A97E97FA6A597.

Ghanbari:2023:GRE

[GA23] Nima Ghanbari and Saeid Alikhani. A graph related to the Euler ϕ function. The Mathematical Gazette, 107(569):263–272, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/

journals/mathematical-gazette/article/graph-related-
to-the-euler-o-function/1103DB07DA366833BD82215ED2516184.

Gadbois:2020:CCF

[Gad20] Steve Gadbois. 104.12 From calendar coincidence to factorials to Ramanujan. The Mathematical Gazette, 104(560): 304–306, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10412-
from-calendar-coincidence-to-factorials-to-ramanujan/
930D0724FC402DD8F71E95A7A3FC692F.

Gardiner:2023:GHM

[Gar23] Tony Gardiner. Geoffrey Howson (9 May 1931–1 November 2022). The Mathematical Gazette, 107(568):114–119, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/geoffrey-
howson-9-may-1931-1-november-2022/12B8446D897CE4D91FCD208D84E4E30F.

George:2020:LSO

[Geo20a] Glyn George. Limits and second order ordinary differential equations. The Mathematical Gazette, 104(559):44–48, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/limits-and-second-order-
ordinary-differential-equations/6798DD09910649200583C640EB57D120.

George:2020:PP

[Geo20b] Glyn George. Parallel probabilities. The Mathematical Gazette, 104(560):271–280, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/parallel-probabilities/FE5662BBE51AD2FDEEBA379B7A9163FC.

Ghosh:2022:APB

[Gho22a] Sourangshu Ghosh. 106.43 Another proof of e^{x/y} being irrational. The Mathematical Gazette, 106(567):523–525, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10643-another-proof-of-
exy-being-irrational/9F45D24BAD4DE1A544A32AA867439F51.

Ghosh:2022:API

[Gho22b]  Sourangshu Ghosh.  106.44 Another proof of the irra-

tionality of N^{p/q}. The Mathematical Gazette, 106(567): 525–526, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10644-another-proof-of-the-irrationality-of-
npq/E1889BF7FC04442CC69C62109A40539A.

Giblin:2022:BRH

[Gib22a] Peter Giblin. Book review: How to prove it (third edition) by Daniel J. Velleman, pp 458, £29.99 (paper), ISBN 978-1-10843-953-4, Cambridge University Press (2019). (Also available as hardback, and as an e-book ISBN 978-1-10833-745-8). The Mathematical Gazette, 106(567):567–568, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/how-to-prove-it-third-edition-by-daniel-
j-velleman-pp-458-2999-paper-isbn-9781108439534-cambridge-
university-press-2019-also-available-as-hardback-and-
as-an-ebook-isbn-9781108337458/375A40610BA45C39B0BF94503E85D90B.

Giblin:2022:BRL

[Gib22b] Peter Giblin. Book review: Linear algebra for everyone by Gilbert Strang, pp 368, £49.99 (hard), ISBN 978-1-73314-663-0, Cambridge University Press (2020). The Mathematical Gazette, 106(567):574–575, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/linear-algebra-for-everyone-by-gilbert-strang-
pp-368-4999-hard-isbn-9781733146630-cambridge-university-
press-2020/F39492459AB84A7CD85263A699402B69.

Giblin:2022:X

[Gib22c] Peter Giblin. On 105.28. The Mathematical Gazette, 106 (565):156–158, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10528/452D353917341ADCC208D79EFE8DB4D3.

Giblin:2023:FPS

[Gib23a] Peter Giblin. 107.40 Four points, six distances. The Mathematical Gazette, 107(570):517–523, November 2023. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10740-four-points-six-distances/E10D8AB1D023ED769952609F60926B9A.

Giblin:2023:BRY

[Gib23b] Peter Giblin. Book review: 3000 years of analysis by Thomas Sonar, pp. xx + 706, £109.99 (hard), ISBN 978-3-030- 58221-0, also available as an e-book, Birkhäuser (Springer Nature Switzerland AG) 2021. The Mathematical Gazette, 107(570): 567–569, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/3000-
years-of-analysis-by-thomas-sonar-pp-xx706-10999-hard-
isbn-9783030-582210-also-available-as-an-ebook-birkhauser-
springer-nature-switzerland-ag-2021/E27F1CBF139F726F65F267BE72FD93AB.

Giblin:2024:BRM

[Gib24] Peter Giblin. Book review: Mathematics is beautiful by Heinz Klaus Strick, pp. 366, £24.99 (paper), ISBN 978-3-662-62688-7, £19.99 (eBook) ISBN 978-3-662-62689-4, Springer Verlag (2021). The Mathematical Gazette, 108(571):188–189, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mathematics-is-
beautiful-by-heinz-klaus-strick-pp-366-2499-paper-isbn-
9783662626887-1999-ebook-isbn-9783662626894-springer-
verlag-2021/EABE0B30D965EF731E6C061B8AFE81AA.

Girban:2023:DIT

[Gîr23] Alexandru Gîrban. Desargues’ involution theorem: from history to applications. The Mathematical Gazette, 107(568):44– 55, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/desargues-
involution-theorem-from-history-to-applications/25A88AC6960E300A2DD858BB9573878B.

Gordon:2023:ISC

[Gor23a] Russell A. Gordon. Integrating sine and cosine Maclaurin remainders. The Mathematical Gazette, 107(568):96–102, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/integrating-sine-
and-cosine-maclaurin-remainders/3EE35D9155EDB85334C81E2F51F82321.

Goron:2023:MRT

[Gor23b] Paul Goron. 107.09 A metric relation on the triangle and the circle. The Mathematical Gazette, 107(568):146–147, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10709-a-metric-relation-
on-the-triangle-and-the-circle/ED6B146AFFEF24A9C59749316CDC7810.

Griffiths:2020:BRM

[Gri20] Martin Griffiths. Book review: The mathematics of everyday life by Alfred S. Posamentier and Christian Spreitzer, pp. 424, $25 (hard), ISBN 978-1-63388-387-1, Prometheus Books (2018). The Mathematical Gazette, 104(560):367–368, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mathematics-of-everyday-
life-by-alfred-s-posamentier-and-christian-spreitzer-
pp-424-25-hard-isbn-9781633883871-prometheus-books-2018/
C259536B6EED105257751B61B7755D1C.

Griffiths:2021:SCF

[Gri21] Martin Griffiths. 105.01 On some composite Fibonacci expressions. The Mathematical Gazette, 105(562):106–108, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10501-on-some-composite-
fibonacci-expressions/962C6CBAE91D197480CD67D200E9942F.

Griffiths:2024:AT

[Gri24] Jonny Griffiths. ABC-triangles. The Mathematical Gazette, 108(571):78–83, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/abctriangles/707190B527DA80251DD9ADB926B1CE91.

Grimmett:2021:LBP

[GS21] Geoffrey R. Grimmett and David R. Stirzaker. The lost boarding pass and other practical problems. The Mathematical Gazette, 105(563):216–221, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/lost-boarding-pass-and-other-practical-problems/
DD427C86C04E53E808D4BA12FCD44C29.

Girban:2022:PPJ

[GS22] Alexandru Gîrban and Bogdan D. Suceavă. Power of a point: from Jakob Steiner to modern applications. The Mathematical Gazette, 106(565):41–53, March 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/power-of-a-point-from-jakob-steiner-
to-modern-applications/E669C249603C7F87842341A41485FFD1.

Gardiner:2023:WMG

[GT23]  Tony Gardiner and Chris Tew.  What makes a good

maths teacher? The Mathematical Gazette, 107(569):286– 300, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/what-
makes-a-good-maths-teacher/EDE80B73095A6345B5B7112ED362D1B7.

Gupta:2024:DRA

[Gup24] Shyam Sunder Gupta. 108.04 Digital root analysis of Smith numbers. The Mathematical Gazette, 108(571):124–130, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10804-digital-
root-analysis-of-smith-numbers/E5E1B37C2D4F429A0EADEF53EF3EF367.

Hajja:2020:MPG

[Haj20] Mowaffaq Hajja. 104.17 More proofs of the AM–GM inequality. The Mathematical Gazette, 104(560):318–326, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10417-more-proofs-of-the-
amgm-inequality/089FD8AFA839D001F8C8BDC97E438896.

Hajja:2021:EAA

[Haj21] Mowaffaq Hajja. Equality of areas among the ears of the Routh triangle. The Mathematical Gazette, 105(563):245–252, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/equality-
of-areas-among-the-ears-of-the-routh-triangle/545060D24C42DD2C3FBF09A87027EA8C.

Hajja:2022:ASA

[Haj22] Mowaffaq Hajja. The arbitrariness of the semi-angle-bisectors of a triangle. The Mathematical Gazette, 106(565):78–83, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/arbitrariness-
of-the-semianglebisectors-of-a-triangle/3D7C89F878F2ECE1BAFD16D17C8CD343.

Hajja:2023:SLTa

[Haj23a] Mowaffaq Hajja. 107.11 The Steiner–Lehmus Theorem à la Ceva. The Mathematical Gazette, 107(568):149–153, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10711-the-steinerlehmus-
theorem-a-la-ceva/4A187614886683150B08FD98DFCDC405.

Hajja:2023:SLTb

[Haj23b] Mowaffaq Hajja. 107.12 The Steiner-Lehmus theorem à la Euclid. The Mathematical Gazette, 107(568):153–155, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10712-the-steinerlehmus-
theorem-a-la-euclid/A88CFEA770911A883FA037787781DC85.

Hall:2020:BRS

[Hal20a] Peter Hall. Book review: Senior problems by Andrew Jobbings, pp. 285, £16 (paper), ISBN 978-1-90600-133-9, UK Mathematics Trust (2018). The Mathematical Gazette, 104(560):380, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/senior-problems-
by-andrew-jobbings-pp-285-16-paper-isbn-9781906001339-
uk-mathematics-trust-2018/C7306D3F5EFA4238A7568F666734FF8F.

Hall:2020:BRR

[Hal20b] Peter Hall. Book review: The room in the elephant by Chris Pritchard, pp. 208, £18 (paper), ISBN 978-1-91161-600-9, The Mathematical Association (2019). The Mathematical Gazette, 104(560):381, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/room-
in-the-elephant-by-chris-pritchard-pp-208-18-paper-isbn-
9781911616009-the-mathematical-association-2019/401E5F34F5F0CB2AF3954B83A56DEAA8.

Hall:2021:BRM

[Hal21] Peter Hall. Book review: Math makers by Alfred S. Posamentier and Christian Spreizer, pp. 448, $25, ISBN 978-1-63388-520-2, Prometheus Books (2019). The Mathematical Gazette, 105(562):178, March 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/math-makers-by-alfred-s-posamentier-
and-christian-spreizer-pp-448-25-isbn-9781633885202-
prometheus-books-2019/FB224585B3935C31221C6FD8783AAA55.

Hall:2022:BRF

[Hal22a] Peter Hall. Book review: Fibonacci’s rabbits by Adam Hart-Davis, pp. 176, £12.99 (paper), ISBN 978-1-912827-03-9, Modern Books (2019). The Mathematical Gazette, 106(565):191, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonaccis-rabbits-by-adam-hartdavis-pp-176-
1299-paper-isbn-9781-912827039-modern-books-2019/D1C56DFBA72389E0E70A458F2FCA27F5.

Hall:2022:BRP

[Hal22b] Peter Hall. Book review: Pack up a penguin, journeys into the mathematics of area by Chris Pritchard, pp. 240, £19.00 (MA members £13.30), ISBN 978-1-911616-08-5, The Mathematical Association (2020). The Mathematical Gazette, 106(566):361, July 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/pack-up-a-penguin-journeys-into-the-
mathematics-of-area-by-chris-pritchard-pp-240-1900-ma-
members-1330-isbn-9781911616085-the-mathematical-association-
2020/B70EBF05B0A53128931FC7B7B9E2C9FF.

Hall:2022:BRQ

[Hal22c] Peter Hall. Book review: Quantitative reasoning by Eric Zaslow, pp. 227, £26.99 (paper), ISBN 978-1-108-41090-8, Cambridge University Press (2020) (e-version reviewed). The Mathematical Gazette, 106(566):374–375, July 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/quantitative-reasoning-

by-eric-zaslow-pp-227-2699-paper-isbn-9781108410908-
cambridge-university-press-2020-eversion-reviewed/88762EF4BDA123E126632837A159006F.

Hall:2023:BRA

[Hal23] Peter Hall. Book review: All the mathematics you missed (but needed to know for graduate school) (second edition) by Thomas A. Garrity, pp. 388, £22.99 (paper), ISBN 978-1-009-00919-5, Cambridge University Press (2021). The Mathematical Gazette, 107(570):561–562, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/all-the-mathematics-you-missed-but-needed-to-
know-for-graduate-school-second-edition-by-thomas-a-
garrity-pp-388-2299-paper-isbn-9781009009195-cambridge-
university-press-2021/3D92EFA9DE951B86FD31D754D13F0A09.

Hammond:2020:AFD

[Ham20] Christopher N. B. Hammond. 104.16 Another function with discontinuous derivative. The Mathematical Gazette, 104(560): 315–318, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10416-
another-function-with-discontinuous-derivative/3C6F935A8B5E88C014808E96909DD1CC.

Hassani:2021:PW

[Has21a] Mehdi Hassani. 105.22 Proof without words. The Mathematical Gazette, 105(563):303, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10522-proof-without-words/FCEBF22B1B2D9AABC1BC1DDCBF697843.

Hassani:2021:CMT

[Has21b] Mehdi Hassani. 105.24 Conditional 2 × 2 matrices with three prime elements and given determinant. The Mathematical Gazette, 105(563):305–306, July 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10524-conditional-2-2-matrices-with-
three-prime-elements-and-given-determinant/32009ABC926B0DA555A011E577096A90.

Hassani:2022:OPW

[Has22] Mehdi Hassani. 106.09 Observations on a proof without words. The Mathematical Gazette, 106(565):133–134, March 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10609-observations-on-a-
proof-without-words/22985FB4DBB460A49461C1446E609F14.

Hassani:2024:PWL

[Has24] Mehdi Hassani. 108.12 Proof without words: a lower bound for n!. The Mathematical Gazette, 108(571):146, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10812-proof-without-words-
a-lower-bound-for-n/336D4B0054867CB1B076A305308F12D8.

Haugland:2023:ICM

[Hau23] Ole Anton Haugland. 107.46 Illustrating complex mappings with Excel. The Mathematical Gazette, 107(570):537– 542, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10746-
illustrating-complex-mappings-with-excel/4BFEA7B3A101198D1D3C3DC1F6BF5695.

Haworth:2020:BRB

[Haw20] Anne Haworth. Book review: The book of why by Judea Pearl, pp. 419, $32 (hard), ISBN 978-0-46509-760-9, Basic Books (2018). The Mathematical Gazette, 104(560):366–367, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/book-of-why-by-
judea-pearl-pp-419-32-hard-isbn-9780465097609-basic-
books-2018/A301B437EA11971F6AA8C68E382534A8.

Haworth:2021:BRB

[Haw21a] Anne Haworth. Book review: Beyond infinity by Eugenia Cheng, pp. 202, £12.99 (paper), ISBN 978-0-46509-481-3, also available as e-book, Profile Books (2017). The Mathematical Gazette, 105(563):381–382, July 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/beyond-infinity-by-eugenia-cheng-pp-
202-1299-paper-isbn-9780465094813-also-available-as-
ebook-profile-books-2017/71F8A3F954DE90DA754B99F51AFF9ABA.

Haworth:2021:BRM

[Haw21b] Anne Haworth. Book review: More puzzles from Pie edited by Wil Ransome, pp. 96, £10 (paper) (£7 for members), ISBN 978-0-90658-889-5, The Mathematical Association (2016). The Mathematical Gazette, 105(563):381, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/more-puzzles-from-pie-edited-by-wil-ransome-
pp-96-10-paper-7-for-members-isbn-9780906588895-the-
mathematical-association-2016/6864ECC1F18A4F8D5344C9ADBA42DB15.

Haworth:2024:BRC

[Haw24] Anne Haworth. Book review: Can fish count? by Brian Butterworth, pp. 373, £20, (hard), ISBN 978-1-5294-1125-6, Quercus Books (2022). The Mathematical Gazette, 108(571):180, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/can-fish-count-
by-brian-butterworth-pp-373-20-hard-isbn-9781529411256-
quercus-books-2022/EC3306664D275D6DF9A5D0708974A395.

Humenberger:2024:GLA

[HE24] Hans Humenberger and Franz Embacher. 108.17 On a generalisation of the Lemoine axis. The Mathematical Gazette, 108(571): 154–158, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10817-
on-a-generalisation-of-the-lemoine-axis/3BB34C60317015941A375AA96ECA80BF.

Hewitt:2022:BRC

[Hew22] Samuel Hewitt. Book review: Curves for the mathematically curious by Julian Havil, pp. 259, £14.99 (paper), ISBN 978-0-691-20613-4, Princeton University Press (2019). The Mathematical Gazette, 106(565):184–185, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/curves-for-the-
mathematically-curious-by-julian-havil-pp-259-1499-paper-
isbn-9780691206134-princeton-university-press-2019/356F4C2FAB34E7F81A76FF0A70A1D890.

Hewitt:2023:BRT

[Hew23] Samuel Hewitt. Book review: Trigonometric delights by Eli Maor, pp. 236, £17.95 (paper), ISBN 978-0-691-20219-8, Princeton University Press (2020). The Mathematical Gazette, 107(568):185–186, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/trigonometric-delights-by-eli-maor-pp-236-1795-
paper-isbn-9780691202198-princeton-university-press-
2020/76A92D706A13F331F8EF5A95F240C490.

Hiller:2021:THM

[Hil21] Josh Hiller. 105.41 Topology haiku matrix. The Mathematical Gazette, 105(564):516, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10541-topology-haiku-matrix/189F75B7E0D50BA8E820F7D9643F5636.

Hajja:2023:MCP

[HK23a] Mowaffaq Hajja and Panagiotis T. Krasopoulos. More characterisations of parallelograms. The Mathematical Gazette, 107 (568):76–83, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/more-
characterisations-of-parallelograms/6323236D93C189C11954F2C895A845F1.

Hajja:2023:YMC

[HK23b] Mowaffaq Hajja and Panagiotis T. Krasopoulos. Yet more characterisations of parallelograms. The Mathematical Gazette, 107(569):225–233, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/yet-more-characterisations-of-parallelograms/
94EA844B4918F65A46D399579D9AE941.

Hassani:2023:PWI

[HL23] Mehdi Hassani and Gerry Leversha. 107.21 Proof without words: An inverse tangent inequality. The Mathematical Gazette, 107(569):323–324, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10721-proof-without-words-an-inverse-tangent-
inequality/0127F00C97F7C25D5AF04421A5B8890C.

Heffernan:2024:EPP

[HLM24] Robert Heffernan, Nick Lord, and Des MacHale. Euler’s primeproducing polynomial revisited. The Mathematical Gazette, 108(571):69–77, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/eulers-primeproducing-polynomial-revisited/1F76183BF30CECE688E31E8F10F37AC8.

Hoare:2020:BRM

[Hoa20] Graham T. Q. Hoare. Book review: Millions billions zillions by Brian W. Kernighan, pp. 160, £17.99 (hard), ISBN 978-0-691-18277-3, Princeton University Press (2018). The Mathematical Gazette, 104(560):362, July 2020. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/millions-billions-zillions-by-brian-
w-kernighan-pp-160-1799-hard-isbn-9780691182773-princeton-
university-press-2018/05F7430558CF469ECE3B524DA30A5678.

Hopkins:2022:BRS

[Hop22a] David Hopkins. Book review: Stochastic modelling of reaction– diffusion processes by Radek Erban and S. Jonathan Chapman, pp. 308, £36.99 (paperback), ISBN 978-1-108-70300-0, Cambridge University Press (2020). The Mathematical Gazette, 106(565):186–187, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/stochastic-modelling-of-reactiondiffusion-processes-
by-radek-erban-and-s-jonathan-chapman-pp-308-3699-paperback-
isbn-9781108703000-cambridge-university-press-2020/BDC1A60984AC1905A6A3B6AE7DB16214.

Hopkins:2022:DP

[Hop22b] David Hopkins. Dropping plates. The Mathematical Gazette, 106(566):193–205, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/dropping-plates/3EDCDD2CB832F1AB76D385C54BF5414E.

Haque:2023:PID

[HP23] Nazrul Haque and Ángel Plaza. 107.06 Proving inequalities via definite integration: a visual approach. The Mathematical Gazette, 107(568):136–140, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10706-proving-inequalities-via-definite-integration-
a-visual-approach/FF7D29BA8BA244E5EBFE59F059248389.

Hamzic:2020:TPG

[HS20] Dina Kamber Hamzić and Zenan Sabanac. Two plane geometry problems approached through analytic geometry. The Mathematical Gazette, 104(560):255–261, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/two-plane-geometry-problems-approached-through-
analytic-geometry/2323F2093AEFC6C3469630F3B4120D56.

Hu:2022:GIC

[Hu22] Hailiang Hu. 106.04 a geometric interpretation of Cramer’s rule. The Mathematical Gazette, 106(565):124–125, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10604-a-geometric-interpretation-
of-cramers-rule/145C6FF01298FF7FF56F7940B845C866.

Humenberger:2020:FTG

[Hum20] Hans Humenberger. 104.08 Finding triangles with given circummedial triangle. The Mathematical Gazette, 104(559):164–168, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10408-finding-
triangles-with-given-circummedial-triangle/FC3417E92EF5765EC736C7879DC6DA7A.

Hunacek:2020:BRF

[Hun20a] Mark Hunacek. Book review: A first course in differential geometry by L. M. Woodward and J. Bolton, pp. 263, £29.99 (paper), ISBN 978-1-108-44102-5, Cambridge University Press (2018). The Mathematical Gazette, 104(560): 374–376, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/first-
course-in-differential-geometry-by-l-m-woodward-and-
j-bolton-pp-263-2999-paper-isbn-9781108441025-cambridge-
university-press-2018/698118AFF0D74DE3D00B8E6067534088.

Hunacek:2020:BRS

[Hun20b] Mark Hunacek. Book review: A short course in differential topology by Bjorn Ian Dundas, pp. 251, £39.99 (hard), ISBN 978-1-10842-579-7, Cambridge University Press (2018). The Mathematical Gazette, 104(560):371–373, July 2020. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/short-course-in-differential-topology-by-bjorn-
ian-dundas-pp-251-3999-hard-isbn-9781108425797-cambridge-
university-press-2018/3638FC2C39FC0CD6EC6688A5BBA3C250.

Hunt:2020:BRB

[Hun20c] Francis Hunt. Book review: Is that a big number? by Andrew C. A. Elliott, pp. 338, £18.99 (hard), ISBN 978-0-19882-122-9, Oxford University Press (2018). The Mathematical Gazette, 104(560):362–364, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/is-that-a-big-number-by-andrew-c-a-elliott-pp-
338-1899-hard-isbn-9780198821229-oxford-university-press-
2018/F39CDE62611B0921E2E10DE43EC61682.

Hunacek:2021:BRGb

[Hun21a] Mark Hunacek. Book review: A gentle course in local class field theory by Pierre Guillot, pp. 293, £28.99 (paper), ISBN 978-1-108-43224-5, Cambridge University Press (2018). The Mathematical Gazette, 105(563):371–373, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/gentle-course-in-local-class-field-theory-by-
pierre-guillot-pp-293-2899-paper-isbn-9781108432245-
cambridge-university-press-2018/F414F5140A3F2B728ADD04462F0D47FD.

Hunacek:2021:BRD

[Hun21b] Mark Hunacek. Book review: Discrete harmonic analysis by Tullio Ceccherini-Silberstein, Fabio Scarabotti and Filippo Tolli, pp. 572, £74.99 (hard), ISBN 978-1-10718-233-2, also available as ebook, Cambridge University Press (2018). The Mathematical Gazette, 105(563):373–375, July 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/discrete-harmonic-analysis-by-tullio-
ceccherinisilberstein-fabio-scarabotti-and-filippo-tolli-
pp-572-7499-hard-isbn-9781107182332-also-available-as-
ebook-cambridge-university-press-2018/924592A60D6B05C040F7E77C086C9F84.

Hunacek:2021:BRGa

[Hun21c] Mark Hunacek. Book review: Geometric and topological inference by Jean-Daniel Boissonnat, Frédéric Chazal and Mariette Yvinec, pp. 233, £28.99 (paper), ISBN 978-1-108-41089-2, Cambridge University Press (2018). The Mathematical Gazette, 105(562):184–185, March 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/geometric-and-topological-inference-
by-jeandaniel-boissonnat-frederic-chazal-and-mariette-
yvinec-pp-233-2899-paper-isbn-9781108410892-cambridge-
university-press-2018/0B23CA25507EF647BB17A5ADB04E2855.

Hunt:2021:BRMb

[Hun21d] Francis Hunt. Book review: Making up your own mind by Edward B. Burger, pp. 136, £16.99 (hard), ISBN 978-0-69-118278-0, Princeton University Press (2018). The Mathematical Gazette, 105(563):369–370, July 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/making-up-your-own-mind-by-edward-b-
burger-pp-136-1699-hard-isbn-9780691182780-princeton-
university-press-2018/4223490C6965B6ABB2D56FD71247B345.

Hunt:2021:BRMa

[Hun21e] Francis Hunt. Book review: Maths on the back of an envelope: Clever ways to (roughly) calculate anything by Rob Eastaway, pp. 208, £9.99 (hard), ISBN 978-0-00-832458-2, Harper Collins (2019). The Mathematical Gazette, 105(562):190, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/maths-on-the-back-of-an-
envelope-clever-ways-to-roughly-calculate-anything-by-
rob-eastaway-pp-208-999-hard-isbn-9780008324582-harper-
collins-2019/DA4E650C2327DBB7EA25EA0542BA9E85.

Hunacek:2022:BRIb

[Hun22a] Mark Hunacek. Book review: An introduction to functional analysis by James C. Robinson, pp 248, £29.99 (paper), ISBN 978-0-521-72839-3, Cambridge University Press (2020). The Mathematical Gazette, 106(566):375–376, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-introduction-to-functional-analysis-by-james-

c-robinson-pp-248-2999-paper-isbn-9780521728393-cambridge-
university-press-2020/2298842DCF45C2A64FE3DDD2D3EB6043.

Hunacek:2022:BRF

[Hun22b] Mark Hunacek. Book review: Fundamentals of graph theory by Allan Bickle, pp. 336, $85, ISBN 978-1-4704-5342-8, American Mathematical Society (2020). The Mathematical Gazette, 106(566):379–380, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fundamentals-of-graph-theory-by-allan-bickle-
pp-336-85-isbn-9781470453428-american-mathematical-society-
2020/459590BB34F0B1F30DBBF856862B6E67.

Hunacek:2022:BRIa

[Hun22c] Mark Hunacek. Book review: Introduction to approximate groups by Matthew C. H. Tointon, pp. 205, £26.99 (paper), ISBN 978-1-108-45644-9, Cambridge University Press (2019). The Mathematical Gazette, 106(565):175–176, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/introduction-to-approximate-groups-by-matthew-
c-h-tointon-pp-205-2699-paper-isbn-9781108456449-cambridge-
university-press-2019/6CCBD5448A72EAF275CA1F04B71FF4B8.

Hunacek:2022:BRLa

[Hun22d] Mark Hunacek. Book review: Linear algebra for the young mathematician by Steven Weintraub, pp, 389, $89.00 (hard), ISBN 978-1-4704-5084-7, American Mathematical Society (AMS) (2019). The Mathematical Gazette, 106(565):176–178, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/linear-algebra-
for-the-young-mathematician-by-steven-weintraub-pp-389-
8900-hard-isbn-9781470450847-american-mathematical-society-
ams-2019/6E6B06016019B6078E0F9F5F39B73B03.

Hunacek:2022:BRLb

[Hun22e] Mark Hunacek. Book review: Linear algebra I by Frederick P. Greenleaf and Sophie Marques, pp. 261, $51 (paper), ISBN 978-1-4704-4871-4, American Mathematical Society (2019). The Mathematical Gazette, 106(565):178–179, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/linear-algebra-
i-by-frederick-p-greenleaf-and-sophie-marques-pp-261-
51-paper-isbn-9781470448714-american-mathematical-society-
2019/286BBE9D02ADF85D8B1EB22D8776476B.

Hunacek:2022:BRLc

[Hun22f] Mark Hunacek. Book review: Linear algebra II by Frederick P. Greenleaf and Sophie Marques, pp. 312, £58.95 (paper), ISBN 978-1-4704-5425-8, American Mathematical Society (2020). The Mathematical Gazette, 106(565):180–181, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/linear-algebra-
ii-by-frederick-p-greenleaf-and-sophie-marques-pp-312-
5895-paper-isbn-9781470454258-american-mathematical-
society-2020/653446EB4A1BA296FFF1687662783E68.

Hunacek:2022:BRRb

[Hun22g] Mark Hunacek. Book review: Representations of finite groups of Lie type (2nd edn.) by François Digne and Jean Michel, pp 172, £37.99 (paper), ISBN 978-1-108-72262-9, Cambridge University Press (2020). The Mathematical Gazette, 106(566):372–373, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/representations-of-finite-groups-of-lie-type-
2nd-edn-by-francois-digne-and-jean-michel-pp-172-3799-
paper-isbn-9781108722629-cambridge-university-press-
2020/9053169F5758B30BE68B79EE943C6108.

Hunacek:2022:BRRa

[Hun22h] Mark Hunacek. Book review: Republic of numbers by David Lindsay Roberts, pp. 244, £22, ISBN 978-1-42143-308-0, Johns Hopkins University Press (2019). The Mathematical Gazette, 106(565):172–173, March 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/republic-of-numbers-by-david-lindsay-
roberts-pp-244-22-isbn-978142143-3080-johns-hopkins-
university-press-2019/54FCF243840F71AE8F8DF8B934A8B942.

Hung:2022:NPD

[Hun22i] Tran Quang Hung. 106.12 A new proof of the n-dimensional Pythagorean theorem. The Mathematical Gazette, 106(565): 136–137, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10612-
a-new-proof-of-the-ndimensional-pythagorean-theorem/
AD84063E7E09BA6F288CBEF64F92EEC8.

Hunacek:2023:BRG

[Hun23a] Mark Hunacek. Book review: Geometry Transformed by James R. King, pp 282, £91.50 (paper), ISBN 978-1-4704-6307-6, American Mathematical Society (2021). The Mathematical Gazette, 107(569):376–379, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/geometry-transformed-by-james-r-king-pp-282-
9150-paper-isbn-9781470463076-american-mathematical-
society-2021/09A0F0D1AA749CFD39259B2385431148.

Hunacek:2023:BRN

[Hun23b] Mark Hunacek. Book review: The new era in American mathematics: 1920–1950 by Karen Hunger Parshall, pp. 640, £40 (paper), ISBN 978-0-691-23524-0, Princeton University Press (2022). The Mathematical Gazette, 107 (570):560–561, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/new-era-in-american-mathematics-19201950-by-
karen-hunger-parshall-pp-640-40-paper-isbn-9780691235240-
princeton-university-press-2022/31240D64D635A8D5481E87313BCF407E.

Hunt:2023:BRW

[Hun23c] Francis Hunt. Book review: When least is best: how mathematicians discovered many clever ways to make things as small (or as large) as possible by Paul J. Nahin, pp. 392, £20 (paper), 978-0-69121-876-2, Princeton University Press (2021). The Mathematical Gazette, 107(568):183–185, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/when-least-is-
best-how-mathematicians-discovered-many-clever-ways-
to-make-things-as-small-or-as-large-as-possible-by-paul-

j-nahin-pp-392-20-paper-9780691218762-princeton-university-
press-2021/D4C24613DA32C007C99E390088F39830.

Hunacek:2024:BRL

[Hun24] Mark Hunacek. Book review: Lost in the Math Museum by Colin Adams, pp 209, $35 (paperback), ISBN 978-1-4704-6858-3, American Mathematical Society (2022). The Mathematical Gazette, 108(571):189–190, March 2024. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/lost-in-the-math-museum-
by-colin-adams-pp-209-35-paperback-isbn-9781470468583-
american-mathematical-society-2022/BD792CD2B25409F060DA2A5E4AC1DAD0.

Jacob:2020:BRU

[Jac20] Niels Jacob. Book review: Understanding topology, a practical introduction by Shaun V. Ault, pp. 416, £74.00 (hard), ISBN 978-1-42142-407-1, Johns Hopkins University Press (2018). The Mathematical Gazette, 104(559):189–190, March 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/understanding-topology-
a-practical-introduction-by-shaun-v-ault-pp-416-7400-
hard-isbn-9781421424071-johns-hopkins-university-press-
2018/1E744E48732D2723E0A6DD02F0D394CC.

Jacoby:2021:BRC

[Jac21] Michael Jacoby. Book review: Calculating the cosmos by Ian Stewart, pp 352, £20.00 (hard), ISBN 978-1-78125-433-2, Basic Books (2016). The Mathematical Gazette, 105(563):376–377, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/calculating-the-cosmos-by-ian-stewart-pp-352-
2000-hard-isbn-9781781254332-basic-books-2016/F151B1A7614509E27CE0B0D60987FDC5.

Jahangiri:2022:GCO

[Jah22] Jay Jahangiri. 106.45 A generalisation of a classical opentop box problem. The Mathematical Gazette, 106(567):526– 531, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10645-
a-generalisation-of-a-classical-opentop-box-problem/
2FC966248811D86999DF08D473A34FB2.

Jameson:2020:HCA

[Jam20] G. J. O. Jameson. How close is the approximation by Bernstein polynomials? The Mathematical Gazette, 104 (561):482–494, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/how-close-is-the-approximation-by-bernstein-
polynomials/8977AED8E7B8FE95850F47F084567E86.

Jameson:2021:MRF

[Jam21a] G. J. O. Jameson. 105.09 Monotonic ratios of functions. The Mathematical Gazette, 105(562):129–134, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10509-monotonic-ratios-of-functions/
89C32AA76F70B81ABE953CB018781746.

Jameson:2021:REO

[Jam21b] G. J. O. Jameson. 105.20 Revisiting even and odd squarefree numbers. The Mathematical Gazette, 105(563):299– 300, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10520-
revisiting-even-and-odd-squarefree-numbers/BCEC5D4187475B19959139F3589DD3EA.

Jameson:2021:MMT

[Jam21c] G. J. O. Jameson. Monotonicity of the midpoint and trapezium estimates for integrals. The Mathematical Gazette, 105 (564):433–441, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/monotonicity-of-the-midpoint-and-trapezium-estimates-
for-integrals/AE092CDA744E3CBFB6044684BADB8950.

Jameson:2021:TWG

[Jam21d] G. J. O. Jameson. Two ways to generate monotonic sequences: convexity and ratios. The Mathematical Gazette, 105(562):16– 26, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/two-ways-
to-generate-monotonic-sequences-convexity-and-ratios/
2E4671537901ECAB78C7A5C952E4550D.

Jameson:2022:ESS

[Jam22a] G. J. O. Jameson. Equal sums, sums of squares and sums of cubes. The Mathematical Gazette, 106(565):54– 60, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/equal-
sums-sums-of-squares-and-sums-of-cubes/E32D2F1CD2E965ED9CB8ADC97E6C07C9.

Jameson:2022:MPC

[Jam22b] G. J. O. Jameson. The majorisation principle for convex functions. The Mathematical Gazette, 106(565):95–102, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/majorisation-principle-
for-convex-functions/8ABE6BA3AD1B1D69E8102F58F7198BE0.

Jameson:2022:PSR

[Jam22c] Graham Jameson. On ‘A pretty series revisited’. The Mathematical Gazette, 106(566):350, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-a-pretty-series-revisited/E73324C990567DD2DBEA6113A38742AF.

Jameson:2023:ALC

[Jam23a] G. J. O. Jameson. Approximating Lipschitz and continuous functions by polynomials; Jackson’s theorem. The Mathematical Gazette, 107(569):273–285, July 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/approximating-lipschitz-
and-continuous-functions-by-polynomials-jacksons-theorem/
5AC47AD5026199C84A46EC8F21BFDEAB.

Jameson:2023:FCT

[Jam23b] G. J. O. Jameson. Fourier coefficients of 1/sinθ and θ/sinθ. The Mathematical Gazette, 107(570):412–421, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/fourier-coefficients-of-
1-over-sin-theta-and-thetaover-sin-theta/C3DE806D8C73E45E9BFA9482835860A9.

Jameson:2023:HIA

[Jam23c] G. J. O. Jameson. Hardy’s inequality for averages. The Mathematical Gazette, 107(568):25–34, March 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/hardys-inequality-for-averages/D92B41B6401F05D485419405EF7465B0.

Jameson:2023:VPP

[Jam23d] G. J. O. Jameson. Various perspectives on a pair of simple inequalities for log x. The Mathematical Gazette, 107 (570):430–437, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/various-perspectives-on-a-pair-of-simple-inequalities-
for-log-x/FEED7E17AD406EBEC698CA8C8561BE97.

Jameson:2023:X

[Jam23e] Graham Jameson. On 107.27. The Mathematical Gazette, 107(570):549, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10727/62350751882E080D733611BDD3866D98.

Jewess:2024:OIE

[Jew24] Michael Jewess. xy = cos(x + y) and other implicit equations that are surprisingly easy to plot. The Mathematical Gazette, 108(571):1–11, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/xy-cos-x-y-and-other-implicit-equations-that-
are-surprisingly-easy-to-plot/53662F33414BB389B93DD76A75B66D85.

Jiang:2022:IGB

[Jia22] Wei-Dong Jiang. 106.29 An improvement on the garfunkelbankoff inequality. The Mathematical Gazette, 106(566):342– 344, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10629-an-
improvement-on-the-garfunkelbankoff-inequality/5ED044878DA96B136F9F075166726078.

Johnson:2021:PMS

[Joh21] Clive Johnson. 105.28 Periodic Möbius sequences. The Mathematical Gazette, 105(563):312–318, July 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10528-periodic-mobius-sequences/A66FF07C94C343566F0103EF2B247E6C.

Josefsson:2020:CMQ

[Jos20] Martin Josefsson. 104.20 A characterisation of midsquare quadrilaterals. The Mathematical Gazette, 104(560):331– 335, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10420-
a-characterisation-of-midsquare-quadrilaterals/E2F464B9D979E0B0F96E9627334F9D2C.

Josefsson:2022:NCB

[Jos22] Martin Josefsson. New characterisations of bicentric quadrilaterals. The Mathematical Gazette, 106(567):414–426, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/new-characterisations-
of-bicentric-quadrilaterals/FC39A554F8619199FB58D8D24BA4C712.

Kaczkowski:2020:CSB

[Kac20] Stephen Kaczkowski. Concurrent sequences of Bernoulli trials. The Mathematical Gazette, 104(561):435–448, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/concurrent-sequences-of-
bernoulli-trials/D750A8BB7B5F8436CB1E20D50EEA3B31.

Katsuura:2023:HCC

[Kat23] Hidefumi Katsuura. 107.24 How to cut cubes into dodecahedra and icosahedra. The Mathematical Gazette, 107 (569):332–340, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10724-how-to-cut-cubes-into-dodecahedra-and-
icosahedra/BA65C4FC1578722CA2BBE92AE4A5A72D.

Kilner:2021:PTA

[KF21a] Steven J. Kilner and David L. Farnsworth. Pairing theorems about parabolas through duality. The Mathematical Gazette, 105(564):385–396, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/

article/pairing-theorems-about-parabolas-through-duality/
2AADE00BD2476863827269D71FBE81E3.

Kilner:2021:PC

[KF21b] Steven J. Kilner and David L. Farnsworth. Parabolic coordinates. The Mathematical Gazette, 105(563):226–236, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/parabolic-
coordinates/E444B0EFA705570E7486A82A24A6B6CD.

Kilner:2023:OMG

[KF23] Steven J. Kilner and David L. Farnsworth. 107.31 Obtaining a more general result from a functional equation by not differentiating. The Mathematical Gazette, 107 (570):495–497, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10731-obtaining-a-more-general-result-from-a-
functional-equation-by-not-differentiating/C699F30D1FE96C48172976EFE0FB57C1.

Kilner:2024:RCP

[KF24] Steven J. Kilner and David L. Farnsworth. Relating constructions and properties through duality. The Mathematical Gazette, 108(571):27–35, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/relating-constructions-and-properties-through-
duality/FCB93A70D27DA1CA59ED7D8F189C5D37.

Kiradjiev:2020:PCT

[Kir20] Kristian B. Kiradjiev. 104.31 Polygons and complex trigonometric identities without complex numbers. The Mathematical Gazette, 104(561):522–527, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10431-polygons-and-complex-trigonometric-identities-
without-complex-numbers/F10340B4846EE84F1BCD54413C4EAEA8.

Kitagawa:2021:PSE

[Kit21] Tomoko L. Kitagawa. Passionate souls: Elisabeth of Bohemia and René Descartes. The Mathematical Gazette, 105(563):193–200, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/passionate-souls-elisabeth-of-bohemia-and-rene-
descartes/0EB78BACD0A3043CEC467A7A17B907BF.

Kataria:2021:GBT

[KM21] Kuldeep Kumar Kataria and Raj Kumar Mistri. 105.42 Generalised binomial theorem via Laplace transform technique. The Mathematical Gazette, 105(564):516–520, November 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10542-generalised-binomial-
theorem-via-laplace-transform-technique/304CABA85C1E28EFDFB3FA8893A9543E.

Kubo:2023:LBP

[KNS23] Shohei Kubo, Toshio Nakata, and Naoki Shiraishi. The lost boarding pass problem: converse results. The Mathematical Gazette, 107(569):234–240, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/lost-boarding-pass-problem-converse-results/
5CD00B2B637EE4CE08C05708118EE6B8.

Kobal:2020:MZP

[Kob20] Damjan Kobal. Matrix zeros of polynomials. The Mathematical Gazette, 104(559):27–35, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/matrix-zeros-of-polynomials/26893ED9800719FF827030074B9FF96B.

Kobal:2023:MPA

[Kob23] Damjan Kobal. A mathematical promenade along parallel paths. The Mathematical Gazette, 107(570):445–453, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mathematical-promenade-
along-parallel-paths/E76B589E7383DDC20C67360CBE8983B1.

Kolosov:2022:UIO

[Kol22] Petro Kolosov. 106.37 An unusual identity for odd-powers. The Mathematical Gazette, 106(567):509–513, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/

mathematical-gazette/article/10637-an-unusual-identity-
for-oddpowers/D6BCDCF54480B913CCDB5E6CA4FF5458.

Koner:2020:UWF

[Kon20] Sourav Koner. 104.02 Unifying Wilson’s and Fermat’s congruence theorems. The Mathematical Gazette, 104(559):146–150, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10402-unifying-
wilsons-and-fermats-congruence-theorems/04A0030239F50ED5DEC70E1175E70FC2.

Kronberg:2024:UPT

[Kro24] Alexander Kronberg. 108.01 A use of Pythagorean triples in a problem in elementary geometry. The Mathematical Gazette, 108 (571):118–119, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10801-
a-use-of-pythagorean-triples-in-a-problem-in-elementary-
geometry/9B05A1EE9F48AEEA3FFA20474F32C15C.

Kuhapatanakul:2021:LPC

[KS21] Kantaphon Kuhapatanakul and Lalitphat Sukruan. Linearly periodic continued fractions. The Mathematical Gazette, 105 (564):442–449, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/linearly-periodic-continued-fractions/C70C17F8DEDAD21F88E5C26A8B151EA3.

Khan:2023:QPT

[KS23] Rasul Khan and Allan Silberger. 107.22 Quick proofs of two inequalities related to the digamma function. The Mathematical Gazette, 107(569):324–327, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10722-quick-proofs-of-two-inequalities-related-
to-the-digamma-function/23FDD7A56180BE050751CA82ECE99B9E.

Kulkarni:2021:SDB

[Kul21] Raghavendra G. Kulkarni. 105.05 Simpler derivation of Bring radical. The Mathematical Gazette, 105(562):120–121, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/

journals/mathematical-gazette/article/10505-simpler-
derivation-of-bring-radical/C1AF458601A7E6C769FC282FB698E4D4.

L:2021:PCa

[L.21a] N. J. L. Problem corner. The Mathematical Gazette, 105 (562):169–174, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/1D0A196DB83027551D47F1A274614D0D.

L:2021:PCb

[L.21b] N. J. L. Problem corner. The Mathematical Gazette, 105(563):358–364, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/BBE1A601BA4119DB9E3F7D036EAE296A.

L:2022:PCa

[L.22a] N. J. L. Problem corner. The Mathematical Gazette, 106 (565):159–166, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/AAB67A850DECE7911AD96FF92BFA7AE9.

L:2022:PCb

[L.22b] N. J. L. Problem corner. The Mathematical Gazette, 106(566):352–357, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/48C0D84FD5C5E0675A67F9727FA18444.

L:2022:PC

[L.22c] N. J. L. Problem corner. The Mathematical Gazette, 106 (567):553–559, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/6A3FC71432B3104C913B23AFD4312084.

L:2023:PC

[L.23] N. J. L. Problem corner. The Mathematical Gazette, 107 (570):550–555, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/22CE02B10231850D616B4E558EF8E244.

L:2024:PC

[L.24] N. J. L. Problem corner. The Mathematical Gazette, 108 (571):167–171, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/951433CC34EFD91BDE1ECF4EF030D509.

Lamphere:2021:CNC

[Lam21] Robert L. Lamphere. 105.15 Could Newton have calculated the deflection angle of starlight? The Mathematical Gazette, 105 (562):154–158, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10515-
could-newton-have-calculated-the-deflection-angle-of-
starlight/32D1E2D1183594AD676C3B47403C93EB.

Laudano:2022:LTF

[Lau22] Francesco Laudano. 106.40 The law of tangents and the formulae of Mollweide and Newton. The Mathematical Gazette, 106(567): 516–517, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10640-
the-law-of-tangents-and-the-formulae-of-mollweide-and-
newton/4E97B3AA3BE8C936598ABD4D25676646.

Laudano:2023:NPE

[Lau23a] Francesco Laudano. 107.38 c^{2} = a^{2} + bd, a new proof of an extension of the Pythagorean theorem. The Mathematical Gazette, 107(570):512–514, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10738-c2-a2-bd-a-new-proof-of-an-extension-of-
the-pythagorean-theorem/FD9B18497733D1BB1E48697EC655AAD5.

Laudano:2023:GAP

[Lau23b] Francesco Laudano. 107.43 Generalised averages of polygons. The Mathematical Gazette, 107(570):528–530, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/

mathematical-gazette/article/10743-generalised-averages-
of-polygons/63BA72A128A4F732C456EAE9BC0142CC.

Lescot:2020:AQR

[Les20] Paul Lescot. An arithmetical question related to perfect numbers. The Mathematical Gazette, 104(559):20–26, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/an-arithmetical-question-
related-to-perfect-numbers/55E6BB7DCA03A3BCDDDA807ED3F93BD7.

Leversha:2020:BRBa

[Lev20a] Gerry Leversha. Book review: The best writing on mathematics 2017 by Mircea Pitici (ed.), pp. 224, £20.00, ISBN 978-0-691-17863-9, Princeton University Press (2018). The Mathematical Gazette, 104(560):368–369, July 2020. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/best-writing-on-mathematics-2017-by-
mircea-pitici-ed-pp-224-2000-isbn-9780691178639-princeton-
university-press-2018/F806460C6FB7E7CB3414DD58C58B77EB.

Leversha:2020:BRBb

[Lev20b] Gerry Leversha. Book review: The best writing on mathematics 2018 by Mircea Pitici (ed.), pp. 250, £20.00, ISBN 978-0-691-18276-6, Princeton University Press (2019). The Mathematical Gazette, 104(560):370–371, July 2020. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/best-writing-on-mathematics-2018-by-
mircea-pitici-ed-pp-250-2000-isbn-9780691182766-princeton-
university-press-2019/D43DE464B48D48613C677B6EEE87CE74.

Leversha:2021:CTC

[Lev21a] Gerry Leversha. 105.30 A Cinderella theorem in circle geometry. The Mathematical Gazette, 105(563):323–327, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10530-a-cinderella-theorem-
in-circle-geometry/605C5A9D6868FAADC39513EE577EA771.

Leversha:2021:WMG

[Lev21b] Gerry Leversha. What makes a good Proof without Words? The Mathematical Gazette, 105(563):271–281, July 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/what-makes-a-good-proof-
without-words/BF589EB1A692A2D7369138A4A34A6444.

Leversha:2022:A

[Lev22a] Gerry Leversha. Acknowledgements. The Mathematical Gazette, 106(567):575, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/acknowledgements/C2798406D08333A84E8581606FE241BF.

Leversha:2022:BRM

[Lev22b] Gerry Leversha. Book review: Mathematics is the poetry of science by Cédric Villani, translated by Malcolm DeBevoise, pp 69, £9.99, ISBN 978-0-19-884643-7, Oxford University Press (2020). The Mathematical Gazette, 106(566):371–372, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mathematics-is-
the-poetry-of-science-by-cedric-villani-translated-by-
malcolm-debevoise-pp-69-999-isbn-9780198846437-oxford-
university-press-2020/3DB191054DC07FA63DF9BBDC32D77C44.

Leversha:2022:BROb

[Lev22c] Gerry Leversha. Book review: OCR A level Further Mathematics for Core Year 2 (A) by Ben Sparks and Claire Baldwin, pp. 342, £24, ISBN 978-1-4718-8648-5, Hodder Education (2017). The Mathematical Gazette, 106(567):566, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/ocr-a-level-further-mathematics-for-
core-year-2-a-by-ben-sparks-and-claire-baldwin-pp-342-
24-isbn-9781471886485-hodder-education-2017/78B4A79F3D5BD795E4E853F10A36A4AD.

Leversha:2022:BROa

[Lev22d] Gerry Leversha. Book review: OCR A level Mathematics (A) for Year 2 by Sophie Goldie, Val Hanrahan, Cath Moore, Jean-Paul Muscat and Susan Whitehouse, pp. 600, £29.99, ISBN 978-1-4718-5307-4, Hodder Education (2018). The Mathematical Gazette, 106(567):564–565, November 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/ocr-a-level-mathematics-a-for-year-2-
by-sophie-goldie-val-hanrahan-cath-moore-jeanpaul-muscat-
and-susan-whitehouse-pp-600-2999-isbn-9781471853074-
hodder-education-2018/110DEAD0F0C8449F1D0A112D5B699FB9.

Leversha:2022:BRB

[Lev22e] Gerry Leversha. Book review: The best writing on mathematics 2019 by Mircea Pitici (ed.), pp. 272, £20.00, ISBN 978-0-691-19835-4, Princeton University Press (2019). The Mathematical Gazette, 106(566):381–382, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/best-writing-on-mathematics-2019-by-
mircea-pitici-ed-pp-272-2000-isbn-9780691198354-princeton-
university-press-2019/FC6B89FFEBBCFE76A7AFF28818918501.

Leversha:2022:BRW

[Lev22f] Gerry Leversha. Book review: The wonder book of geometry by David Acheson, pp. 280, £12.99, ISBN 978-0-19-884638-3, Oxford University Press (2020). The Mathematical Gazette, 106(566):377–378, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/wonder-book-of-geometry-by-david-acheson-pp-
280-1299-isbn-9780198846383-oxford-university-press-
2020/ACB874749D41C5099E2592033FF7C45E.

Leversha:2023:BROc

[Lev23a] Gerry Leversha. Book review: OCR A level Further Mathematics Discrete (A) by Nick Geere, pp. 162, £21, ISBN 978-1-5104-3337-3, Hodder Education (2018). The Mathematical Gazette, 107 (568):179–181, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/ocr-a-
level-further-mathematics-discrete-a-by-nick-geere-pp-
162-21-isbn-9781510433373-hodder-education-2018/581640237F62E5920F79EF99DA7678FE.

Leversha:2023:BROa

[Lev23b] Gerry Leversha. Book review: OCR A level Further Mathematics Mechanics (A) by Jean-Paul Muscat, pp. 362, £24, ISBN 978-1-5104-1451-8, Hodder Education (2018). The Mathematical Gazette, 107(568):177, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/ocr-a-level-further-mathematics-mechanics-a-
by-jeanpaul-muscat-pp-362-24-isbn-9781510414518-hodder-
education-2018/42ADB3B37C6D9FB629C1A0DF49759003.

Leversha:2023:BROb

[Lev23c] Gerry Leversha. Book review: OCR A level Further Mathematics Statistics (A) by Jean-Paul Muscat, pp. 362, £24, ISBN 978-1-5104-1451-8, Hodder Education (2018). The Mathematical Gazette, 107(568):178–179, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/ocr-a-level-further-mathematics-statistics-a-
by-jeanpaul-muscat-pp-362-24-isbn-9781510414518-hodder-
education-2018/AE05BFAA331888C550FC081CF3C4A53D.

Leversha:2023:BRB

[Lev23d] Gerry Leversha. Book review: The best writing on mathematics 2021 by Mircea Pitici (ed.), pp. 288, £20.00, ISBN 978-0-691-22570-8, Princeton University Press (2022). The Mathematical Gazette, 107(570):570–571, November 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/best-writing-on-mathematics-2021-by-
mircea-piticied-pp-288-2000-isbn-978-0691225708-princeton-
university-press-2022/9B49E1F4B7BD76F9E4968007ADC437A6.

Leversha:2024:BRP

[Lev24] Gerry Leversha. Book review: The polyhedrists: art and geometry in the long sixteenth century by Noam Andrews, pp. 300, $44.95, ISBN 978-0-262-04664-0, MIT Press (2022). The Mathematical Gazette, 108(571):175, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/polyhedrists-art-and-geometry-in-the-
long-sixteenth-century-by-noam-andrews-pp-300-4495-isbn-
9780262046640-mit-press-2022/A0C1DEFDCF637318F53213F70653A0CF.

Lund-Hansen:2022:CAD

[LH22] Lars Lund-Hansen. On ‘Correct answer — dodgy method’. The Mathematical Gazette, 106(566):350–351, July 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/on-correct-answer-dodgy-
method/53B0C1C96BF5EABDCEA5FF72BE56702B.

Ligo:2023:GII

[Lig23] Richard G. Ligo. 107.07 A geometric illustration for infinite sequences and series. The Mathematical Gazette, 107(568): 140–144, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10707-
a-geometric-illustration-for-infinite-sequences-and-
series/9451B22CACC32BFF193858C89E8CE8EA.

Lord:2024:IMC

[LM24] Nick Lord and Des MacHale. Infinitely many composites. The Mathematical Gazette, 108(571):20–26, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/infinitely-many-composites/A522BC5767166463A1B00FECE6B68BD4.

Levrie:2022:CII

[LN22a] Paul Levrie and Amrik Singh Nimbran. 106.21 A class of interesting integrals. The Mathematical Gazette, 106(566):323–325, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10621-a-class-
of-interesting-integrals/A594E61557E9565D664CF8E6F3EDB89B.

Lucas:2022:MSF

[LN22b] Stephen K. Lucas and Amrik Singh Nimbran. Monotonic series for fractions near π and their convergents. The Mathematical Gazette, 106(566):300–309, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/monotonic-series-for-fractions-near-
and-their-convergents/73CAD6F4E882A4643E7C626066EA5891.

Lord:2020:NNG

[Lor20a] Ems Lord. Nurturing the next generation of mathematicians: the case for mathematical fluency. The Mathematical Gazette, 104(561):385–394, November 2020. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/nurturing-the-next-generation-of-mathematicians-
the-case-for-mathematical-fluency/40BEE6C0E8B7A9B4297DBE7C402A0175.

Lord:2020:BRC

[Lor20b] Nick Lord. Book review: Complex analysis by Ian Stewart and David Tall (2nd edn.), pp. 389, £29.99 (paper), ISBN 978-1-108-43679-3, Cambridge University Press (2018). The Mathematical Gazette, 104(559):187–189, March 2020. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/complex-analysis-by-ian-stewart-and-
david-tall-2nd-edn-pp-389-2999-paper-isbn-9781108436793-
cambridge-university-press-2018/3E5D7FBF161D1B8086D5BE2723854561.

Lord:2020:PCa

[Lor20c] Nick Lord. Problem corner. The Mathematical Gazette, 104(559):176–183, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/D96C4A53746F05B976172B71A5BA89B2.

Lord:2020:PCb

[Lor20d] Nick Lord. Problem corner. The Mathematical Gazette, 104(560):353–358, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/96940A86EFBFAB06649EE453B5E02815.

Lord:2021:RRM

[Lor21a] Nick Lord. 105.35 Reconciling remainders in Maclaurin expansions. The Mathematical Gazette, 105(563):334–338, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10535-reconciling-remainders-
in-maclaurin-expansions/054912B55DA7D34FCB75E8C643790016.

Lord:2021:NEA

[Lor21b] Nick Lord. 105.36 New error analyses for some old mensuration formulae. The Mathematical Gazette, 105(563): 339–343, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10536-
new-error-analyses-for-some-old-mensuration-formulae/
74563C94C5272EF73FEF5B723C86F987.

Lord:2021:BRV

[Lor21c] Nick Lord. Book review: 99 variations on a proof by Philip Ording, pp. 260, £20 (hard), ISBN 978-0-691-15883-9, Princeton University Press (2019). The Mathematical Gazette, 105(562):188–189, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/99-variations-on-a-proof-by-philip-ording-pp-
260-20-hard-isbn-9780691158839-princeton-university-
press-2019/BFA699B940677402AA1D6C53441C6142.

Lord:2021:SNS

[Lor21d] Nick Lord. A small note on small angle approximations. The Mathematical Gazette, 105(562):159–161, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/small-note-on-small-angle-approximations/
A0B87CEE7B1731AEF1D51CFBDE384877.

Lord:2022:SOI

[Lor22a] Nick Lord. 106.19 Some observations on inequalities related to Huygens’ inequality. The Mathematical Gazette, 106(566): 316–318, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10619-
some-observations-on-inequalities-related-to-huygens-
inequality/9EF131A4FE51C5E16D546530BDCC667B.

Lord:2022:NPP

[Lor22b] Nick Lord. 106.26 The nested polygons problem revisited. The Mathematical Gazette, 106(566):335–338, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10626-the-nested-polygons-problem-revisited/
A401A0D1D8B66FA7F966607DB4849C90.

Lord:2022:BRC

[Lor22c] Nick Lord. Book review: Calculus for cranks by Nets Hawk Katz, pp. 251, £20 (paper), ISBN 978-0-30024- 279-9, Yale University Press (2021). The Mathematical Gazette, 106 (567):569–570, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/calculus-for-cranks-by-nets-hawk-katz-pp-251-
20-paper-isbn-978030024-2799-yale-university-press-2021/
C7133A3A7D373D619B442DEBFBAC65D9.

Lord:2022:BRI

[Lor22d] Nick Lord. Book review: Introduction to complex variables and applications by Mark J. Ablowitz and Athanassios S. Fokas, pp. 420, £39.99 (paper), ISBN 978-1-10895-972-8, Cambridge University Press (2021). The Mathematical Gazette, 106(567):570–572, November 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/introduction-to-complex-
variables-and-applications-by-mark-j-ablowitz-and-athanassios-
s-fokas-pp-420-3999-paper-isbn-9781108959728-cambridge-
university-press-2021/2B57F3DB706AEABEDF974FB42A8E51DE.

Lord:2022:BRW

[Lor22e] Nick Lord. Book review: Where do numbers come from? by T. W. Körner, pp. 260, £24.99 (paper), ISBN 978-1-108-73838-5, Cambridge University Press (2020). The Mathematical Gazette, 106(565):188–189, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/where-do-numbers-come-from-by-tw-korner-pp-260-
2499-paper-isbn-9781108738385-cambridge-university-press-
2020/AD8727E7600C1E1979C9F98214A67E17.

Lord:2022:FSI

[Lor22f] Nick Lord. The full story of invariant lines. The Mathematical Gazette, 106(567):547–548, November 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/full-story-of-invariant-
lines/2A46A6F8AAA57E00DC4DA6DBA75686BE.

Lord:2022:HIC

[Lor22g] Nick Lord. How to impress a chemist (again!). The Mathematical Gazette, 106(565):149–150, March 2022. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/how-to-impress-a-chemist-again/97B1CC0C126D7B5A6BD3CEC50F7B7ADC.

Lord:2022:QSE

[Lor22h] Nick Lord. A quick simultaneous evaluation of ∑ k^{3} and ∑ k^{2}. The Mathematical Gazette, 106(565):149, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/quick-simultaneous-evaluation-of-k3-
and-k2/11CBA304F15CB14E6FECABB8515260DC.

Lord:2022:TUO

[Lor22i] Nick Lord. Two unusual optimisation problems. The Mathematical Gazette, 106(565):151–154, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/two-unusual-optimisation-problems/8C7C4F081339CADA0306E7090D8784E9.

Lord:2023:IES

[Lor23a] Nick Lord. 107.08 An interesting equivalent of squaring the circle. The Mathematical Gazette, 107(568):144–145, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10708-an-
interesting-equivalent-of-squaring-the-circle/44DBDE50FA7CF6D9F32604091A52AADF.

Lord:2023:TCR

[Lor23b] Nick Lord. 107.17 Two curios related to lattice polygons. The Mathematical Gazette, 107(569):307–312, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10717-two-curios-related-to-lattice-
polygons/D443B0551BD31BBB42792DEED5A61ED1.

Lord:2023:TVA

[Lor23c] Nick Lord. 107.18 A two-variable approach to some standard optimisation problems. The Mathematical Gazette, 107(569): 312–316, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10718-
a-twovariable-approach-to-some-standard-optimisation-
problems/01CE045F7E47B2D27284355D609CF5DB.

Lord:2023:BRC

[Lor23d] Nick Lord. Book review: A course in complex analysis by Saeed Zakeri, pp. 428, £50 (hard), ISBN 978-0-691- 20758-2, Princeton University Press (2021). The Mathematical Gazette, 107(570):566–567, November 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/course-in-complex-analysis-by-saeed-
zakeri-pp-428-50-hard-isbn-9780691-207582-princeton-
university-press-2021/1CB457707667ED5D905C55A2C7EC497C.

Lord:2023:BRA

[Lor23e] Nick Lord. Book review: Algebra and geometry, an introduction to university mathematics (second edition) by Mark V. Lawson, pp. 424, £49.99 (paper), ISBN 978-0-367-56303-5, CRC/Taylor and Francis (2021). The Mathematical Gazette, 107(569):368–369, July 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/algebra-and-geometry-an-introduction-
to-university-mathematics-second-edition-by-mark-v-lawson-
pp-424-4999-paper-isbn-9780367563035-crctaylor-and-francis-
2021/622D65C9FA1DCFAF5EADB1499B3A4B67.

Lord:2023:OFA

[Lor23f] Nick Lord. An odd fact about Cayley tables. The Mathematical Gazette, 107(570):543–544, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-odd-fact-about-cayley-tables/DA1A910D2F7168179C94A8612085BA29.

Lord:2023:Xa

[Lor23g] Nick Lord. On 107.03. The Mathematical Gazette, 107(570):548, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/on-10703/
7AF2C02C48AA417C69754229EDD3667B.

Lord:2023:Xb

[Lor23h] Nick Lord. On 107.09. The Mathematical Gazette, 107 (570):548–549, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10709/90A58D96D03F48B04C9B5F2FF6A46C07.

Lord:2023:FJ

[Lor23i] Nick Lord. On feedback for July 2022. The Mathematical Gazette, 107(569):356, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-feedback-for-july-2022/9A4A7899049054A37CD7995181724841.

Lord:2023:PCa

[Lor23j] Nick Lord. Problem corner. The Mathematical Gazette, 107(568):166–173, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/631B62DE05E7190903681A064478F646.

Lord:2023:PCb

[Lor23k] Nick Lord. Problem corner. The Mathematical Gazette, 107(569):359–364, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/8549C2691BE7D6572B502052C804B838.

Lord:2023:WDG

[Lor23l] Nick Lord. When does (1/4)(a + b + c + d) give the centre of mass of a quadrilateral? The Mathematical Gazette, 107(568):162–163, March 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/when-does-give-the-centre-of-mass-of-
a-quadrilateral/DD48E695DFC524FAC5DDF3A76B88E53E.

Lord:2024:OLP

[Lor24a] Nick Lord. 108.18 A one-line proof of the Finsler–Hadwiger inequality. The Mathematical Gazette, 108(571):158–160, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10818-a-oneline-
proof-of-the-finslerhadwiger-inequality/A306441207B1B51F1A9C5DEFDD225F02.

Lord:2024:AAG

[Lor24b] Nick Lord. Another appearance of the golden ratio. The Mathematical Gazette, 108(571):163–165, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/another-appearance-of-the-golden-ratio/
C766E0E135D0D589B1FE9B88A6F29036.

Lord:2024:BRC

[Lor24c] Nick Lord. Book review: Comic sections plus by Des MacHale, pp. 264, £15.67 (paper), ISBN 978-1-4717-6147-8 Logic Press (2022). The Mathematical Gazette, 108(571): 190–191, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/comic-
sections-plus-by-des-machale-pp-264-1567-paper-isbn-
9781471761478-logic-press-2022/B8640CEC4C61405F93FF397C90E50BB7.

Lord:2024:CTA

[Lor24d] Nick Lord. A cautionary tale about the pole of polar coordinates. The Mathematical Gazette, 108(571):162–163, March 2024. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/cautionary-tale-about-the-
pole-of-polar-coordinates/E52980E889872B9109FDCB0AE2F2E2CD.

Lord:2024:DAB

[Lor24e] Nick Lord. The Devil’s Advocate and the binomial expansion. The Mathematical Gazette, 108(571):161–162, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/devils-advocate-and-the-
binomial-expansion/58A96A0A528396B88BC15C756219FC86.

Lord:2021:GHA

[LR21] Nick Lord and Jenny Ramsden. Graham Hoare: an appreciation. The Mathematical Gazette, 105(563):356–357, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/graham-hoare-an-appreciation/
9269F0D938142FDF45089E7F21CDA6BB.

Lukarevski:2021:BD

[LS21] Martin Lukarevski and J. A. Scott. 105.31 On the Brocard disc. The Mathematical Gazette, 105(563):327–328, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10531-on-
the-brocard-disc/186B296371623743B04B3C944B5B2B59.

Lukarevski:2022:TDI

[LS22] Martin Lukarevski and J. A. Scott. 106.25 Three discs for the incentre. The Mathematical Gazette, 106(566):332– 335, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10625-
three-discs-for-the-incentre/74F4D84CC05E2930C0117D422144619D.

Liao:2023:RGC

[LSS23] Hsin-Chieh Liao, Mark Saul, and Peter J.-S Shiue. Revisiting the general cubic: a simplification of Cardano’s solution. The Mathematical Gazette, 107(570):438–444, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/revisiting-the-
general-cubic-a-simplification-of-cardanos-solution/
7D6BD635E7D6FCE7B11FB9BE09EC03F7.

Lucas:2024:MTD

[Luc24] T. N. Lucas. 108.07 De Moivre’s theorem via difference equations. The Mathematical Gazette, 108(571):136–140, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10807-de-
moivres-theorem-via-difference-equations/F3546BA590C4B840B51286E81B8F885F.

Lukarevski:2020:IAE

[Luk20a] Martin Lukarevski. 104.07 An inequality for the altitudes of the excentral triangle. The Mathematical Gazette, 104 (559):161–164, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10407-an-inequality-for-the-altitudes-of-the-
excentral-triangle/FCD13CE98CB0FB59C91AD04AA632DF1E.

Lukarevski:2020:CTF

[Luk20b] Martin Lukarevski. 104.21 The circummidarc triangle and the Finsler–Hadwiger inequality. The Mathematical Gazette, 104 (560):335–338, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10421-
the-circummidarc-triangle-and-the-finslerhadwiger-inequality/
7B25169B1C62D038A9A87A3205B78111.

Lukarevski:2021:PII

[Luk21a] Martin Lukarevski. 105.13 Proximity of the incentre to the inarc centres. The Mathematical Gazette, 105(562):142– 147, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10513-
proximity-of-the-incentre-to-the-inarc-centres/FBF2F3BF7A64B85E4E4AAD9236C77654.

Lukarevski:2021:X

[Luk21b] Martin Lukarevski. On 105.13. The Mathematical Gazette, 105 (563):355, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/on-10513/
7702211132127BC12ADB0B40AE5AF358.

Lukarevski:2022:ERF

[Luk22a] Martin Lukarevski. 106.14 Exarc radii and the Finsler– Hadwiger inequality. The Mathematical Gazette, 106(565): 138–143, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10614-
exarc-radii-and-the-finslerhadwiger-inequality/4F8579EE9661D8CBA09BFE1B0335E0ED.

Lukarevski:2022:WMG

[Luk22b] Martin Lukarevski. On ‘What makes a good proof without words’. The Mathematical Gazette, 106(566):349, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/on-what-makes-
a-good-proof-without-words/2F5A39C7198B0BF8D89C64F3FA214F8F.

Lukarevski:2023:LIC

[Luk23a] Martin Lukarevski. 107.23 Location of the inarc circle and its point of contact with the circumcircle. The Mathematical Gazette, 107(569):327–331, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10723-location-of-the-inarc-circle-and-its-point-
of-contact-with-the-circumcircle/822F4AC72AB9753C912D9892AA8F137D.

Lukarevski:2023:BRS

[Luk23b] Martin Lukarevski. Book review: Series and products in the development of mathematics (second edition) by Ranjan Roy, Vol. 1 pp. 776, £69.99 (paperback), ISBN 978-1-108-70945-3; Vol. 2 pp. 476, £45.99 (paperback), ISBN 978-1-108-70937-8, Cambridge University Press (2021). The Mathematical Gazette, 107(569):373–376, July 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/series-and-products-in-
the-development-of-mathematics-second-edition-by-ranjan-
roy-vol-1-pp-776-6999-paperback-isbn-9781108709453-vol-
2-pp-476-4599-paperback-isbn-9781108709378-cambridge-
university-press-2021/F6DEACDFF3EA62D5103990D89309A03D.

Lukarevski:2023:X

[Luk23c] Martin Lukarevski. On 106.12. The Mathematical Gazette, 107(568):164–165, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10612/6638417F7F43AA72501E3001A4849745.

Lukarevski:2023:WIR

[Luk23d] Martin Lukarevski. Wolstenholme’s inequality and its relation to the Barrow and Garfunkel–Bankoff inequalities. The Mathematical Gazette, 107(568):70–75, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/wolstenholmes-inequality-and-its-relation-to-
the-barrow-and-garfunkelbankoff-inequalities/809CCB99431AF6E0973EF53D306A035F.

Lukarevski:2024:PWT

[Luk24a] Martin Lukarevski. 108.10 Proof without words: tan(π/12) = 2 − \sqrt{3}, tan(5π/12) = 2 + \sqrt{3}. The Mathematical Gazette, 108 (571):143–144, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10810-

proof-without-words-rmtan-piover-rm12rm2-sqrt-3-rm-tan-
rm5pi-over-rm12rm2-sqrt-rm3/6D49104D14BA2482BF4B44FB8E8AA649.

Lukarevski:2024:BRT

[Luk24b] Martin Lukarevski. Book review: Theory of infinite sequences and series by Ludmila Bourchtein and Andrei Bourchtein, pp. 377, £54.99, (paper), ISBN 978-3-030-79430-9, Springer Verlag (2022). The Mathematical Gazette, 108(571):182–183, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/theory-of-
infinite-sequences-and-series-by-ludmila-bourchteinand-
andrei-bourchtein-pp-377-5499-paper-isbn-9783030794309-
springer-verlag-2022/468EA80413165396B0AEB899A6CAA1A9.

MacGregor:2020:BRA

[Mac20] P. MacGregor. Book review: Applied linear algebra and matrix analysis (2nd edn.) by Thomas S. Shores, pp. 479, £59.99 (hard), ISBN 978-3-319-74747-7, also available as e-book, Springer Verlag (2018). The Mathematical Gazette, 104(560):376, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/applied-linear-algebra-
and-matrix-analysis-2nd-edn-by-thomas-s-shores-pp-479-
5999-hard-isbn-9783319747477-also-available-as-ebook-
springer-verlag-2018/9C7CEE60E38A78299ADDB5250D8A7822.

MacHale:2021:SER

[Mac21a] Des MacHale. 105.16 Some elementary results in number theory. The Mathematical Gazette, 105(563):282–285, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10516-some-elementary-results-
in-number-theory/558C392C89C924A8DE9C95110E4B622A.

MacHale:2021:GM

[Mac21b] Des MacHale. 105.29 A geometric memory. The Mathematical Gazette, 105(563):318–323, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10529-a-geometric-memory/BC8105BE11725C6033473C56466A7C2D.

MacHale:2021:CAD

[Mac21c] Des MacHale. Correct answer — dodgy method. The Mathematical Gazette, 105(564):507–510, November 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/correct-answer-dodgy-method/D3681D2A91E5B227CAFB433F50408608.

MacHale:2021:VNI

[Mac21d] Des MacHale. Verifying non-isomorphism of groups. The Mathematical Gazette, 105(564):467–473, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/verifying-nonisomorphism-of-groups/E7B4791E90B036C88482467226A6E050.

MacHale:2022:DAS

[Mac22] Des MacHale. Dissecting attached squares. The Mathematical Gazette, 106(566):258–268, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/dissecting-attached-squares/5B1C3CE7F6B35E8A50B081B055C25E43.

MacGregor:2023:BRD

[Mac23a] P. G. MacGregor. Book review: The discrete mathematical charms of Paul Erdős by Vasek Chvátal, pp 248, £22.99 (paper), ISBN 978-1-108-92740-6, Cambridge University Press (2021). The Mathematical Gazette, 107(569):382– 383, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/discrete-
mathematical-charms-of-paul-erdos-by-vasek-chvatal-pp-
248-2299-paper-isbn-9781108927406-cambridge-university-
press-2021/F872B26F82E04945243A9527DB1F51F3.

MacGregor:2023:BRP

[Mac23b] Peter MacGregor. Book reviews: A prelude to quantum field theory by John Donoghue and Lorenzo Sorbo, pp. 160, £25 (paper), ISBN 978-0-691-22348-3, Princeton University Press (2022) — What is a quantum field theory? - a first introduction for mathematicians by Michel Talagrand, pp. 741, £69.99 (hard), ISBN 978-1-316-51027-8, Cambridge University Press (2022). The Mathematical Gazette, 107 (570):573–575, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/prelude-to-quantum-field-theory-by-john-donoghue-
and-lorenzo-sorbo-pp-160-25-paper-isbn-9780691223483-
princeton-university-press-2022-what-is-a-quantum-field-
theory-a-first-introduction-for-mathematicians-by-michel-
talagrand-pp-741-6999-hard-isbn-9781316510278-cambridge-
university-press-2022/E9BE6A264038962C3575D219276B23BD.

MacHale:2023:X

[Mac23c] Des MacHale. 2 + 2 = 5. The Mathematical Gazette, 107 (568):164, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/2-2-5/
AD65C675ED23269B11C4906989421603.

MacGregor:2024:BRC

[Mac24] P. MacGregor. Book review: Change and variations, a history of differential equations to 1900 by Jeremy Gray, pp. 261, £29.99, (hard), ISBN 978-3-030-70574-9, Springer Verlag (2021). The Mathematical Gazette, 108(571):176–177, March 2024. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/change-and-variations-a-
history-of-differential-equations-to-1900-by-jeremy-
gray-pp-261-2999-hard-isbn-9783030705749-springer-verlag-
2021/2FBA99A7F8712E3DA395E1D1F8E2DE74.

Mahony:2020:EVS

[Mah20] John D. Mahony. 104.29 On the evaluation of a very specific integral. The Mathematical Gazette, 104(561):518–519, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10429-on-
the-evaluation-of-a-very-specific-integral/9CA2B20716CC5C7AF7BADBE8DDDBDEDA.

Mahony:2021:SMA

[Mah21a] John Mahony. Some mathematical appreciations in the physical sciences. The Mathematical Gazette, 105(562):4– 15, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/some-
mathematical-appreciations-in-the-physical-sciences/
694A3A17F4D65B085F6BCADFA374E280.

Mahony:2021:PB

[Mah21b] John D. Mahony. 105.37 A potter’s brief. The Mathematical Gazette, 105(563):343–348, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10537-a-potters-brief/00DC253E1B5163DC826024B46D949D9A.

Mahony:2021:DQT

[Mah21c] John D. Mahony. Developing quadrature themes. The Mathematical Gazette, 105(564):458–466, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/developing-quadrature-themes/160983DFE86E681CACAD19E64CEE8F1F.

Mahony:2022:C

[Mah22a] John D. Mahony. Correspondence. The Mathematical Gazette, 106(565):155, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/correspondence/9CDD49FE0FB1CD608F00EF6137142203.

Mahony:2022:MAP

[Mah22b] John D. Mahony. A mathematical approximation in the physical sciences. The Mathematical Gazette, 106(566):220–232, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/mathematical-approximation-
in-the-physical-sciences/C3B6CA5FD44DC31A58F209EB980F6FAA.

Mala:2022:BRA

[Mal22] Firdous Ahmad Mala. Book review: Africa and mathematics: from colonial findings back to the Ishango Rods by Dirk Huylebrouck, pp. 229, £27.99 (hardback), ISBN 978-3-030-04036-9 Springer Verlag (2019). The Mathematical Gazette, 106(566):370–371, July 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/africa-and-mathematics-from-colonial-
findings-back-to-the-ishango-rods-by-dirk-huylebrouck-
pp-229-2799-hardback-isbn-9783030040369-springer-verlag-
2019/8311A5C5E0F9CDC0C5EA15BCB6977C41.

Markov:2022:TSP

[Mar22] Lubomir Markov. Two short proofs of the formula ∑_{n=0}^{∞} \frac{1}{(2n+1)^{2}} =

\frac{π^{2}}{8} . The Mathematical Gazette, 106(565):28–31, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/two-short-proofs-of-the-
formula-sumlimitsn-0infty-1-over-2n-12-pi-2-over-8/19A3C95629C55C5BAE764DA40FDE0E3A.

McBride:2021:PMN

[MC21] Adam McBride and Barbara Cullingworth. Peter Michael Neumann OBE: (28 December 1940–18 December 2020). The Mathematical Gazette, 105(562):1–3, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/peter-michael-neumann-obe/4CEE751AE81731197FCEB1EB96FC30C7.

McBride:2021:TTS

[McB21] Adam McBride. 105.17 A tale of two sixes. The Mathematical Gazette, 105(563):285–290, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10517-a-tale-of-two-sixes/CC99078EA11C7E6BFB26055B624C817E.

McLean:2020:WS

[McL20] K. Robin McLean. What is the significance of b^{2} = \frac{1}{2} (3 + \sqrt{5}) ac. The Mathematical Gazette, 104(559):107– 115, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/what-
is-the-significance-of-b2-frac12left-3-sqrt-5-rightac/
290A93ED64C3AACC29AAABBE758B792D.

Melman:2021:PNR

[Mel21] Aaron Melman. 105.04 Polynomials with no real zeros. The Mathematical Gazette, 105(562):117–120, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10504-polynomials-with-no-real-zeros/
EF820828887AD53825571642B9BE2AE0.

Mercer:2023:ECS

[Mer23] Peter R. Mercer. 107.20 Euler’s constant and the speed of convergence. The Mathematical Gazette, 107(569):320– 323, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10720-
eulers-constant-and-the-speed-of-convergence/9B5F9810114F8C9AF0A40F7BFCC679E0.

Mestrovic:2023:NIP

[Mes23] Romeo Mestrović. 107.32 A new inductive proof of the AM– GM inequality. The Mathematical Gazette, 107(570):497– 499, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10732-
a-new-inductive-proof-of-the-am-gm-inequality/05ABEFD2218BF948BCC57F1BBDB40BF6.

Meyer:2020:EWB

[Mey20] Joerg Meyer. An easy way to Brahmagupta’s formula for the area of a cyclic quadrilateral. The Mathematical Gazette, 104(559):174–175, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-easy-way-to-brahmaguptas-formula-for-the-
area-of-a-cyclic-quadrilateral/8623C6553B394ED271CE3D63BD96F169.

Moreno:2020:ID

[MGC20] Samuel G. Moreno and Esther M. García-Caballero. 104.01 Irrationality of \sqrt{2} with determinants. The Mathematical Gazette, 104(559):143–146, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10401-irrationality-of-sqrt-2-with-determinants/
397CE6DE02738D48FE46C7D9D08428F7.

Majumdar:2023:FDE

[MS23] Dipramit Majumdar and B. Sury. 107.15 Fruit diophantine equation. The Mathematical Gazette, 107(569):302–306, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10715-fruit-diophantine-
equation/0E52021D65DC7DC91F666AECF8F718A9.

Mukherjee:2023:GTR

[Muk23] Rajib Mukherjee. 107.04 A geometric telescope revisited. The Mathematical Gazette, 107(568):128–130, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10704-a-geometric-telescope-revisited/
85316751AD1D6B9DC7DF404529D7959C.

Munkhammar:2020:RZF

[Mun20] Joakim Munkhammar. 104.32 The Riemann zeta function as a sum of geometric series. The Mathematical Gazette, 104 (561):527–530, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10432-the-riemann-zeta-function-as-a-sum-of-
geometric-series/D7C4EFAA90E409FB542712C041453639.

Missa:2022:AFC

[MY22a] Abdel Missa and Chrif Youssfi. An alternative formula for the cubic equation. The Mathematical Gazette, 106 (567):474–479, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-alternative-formula-for-the-cubic-equation/
97633DB5BE051C9FF60177C1A454F15A.

Missa:2022:NMS

[MY22b] Abdel Missa and Chrif Youssfi. A novel method to solve the quartic equation. The Mathematical Gazette, 106(567):480– 486, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/novel-
method-to-solve-the-quartic-equation/9DB7A0415E71EAA955BB6E4F19DD7C5A.

Nakata:2022:CEP

[Nak22] Toshio Nakata. Characterisation of equalisation problems via random walks. The Mathematical Gazette, 106(565):61–67, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/characterisation-
of-equalisation-problems-via-random-walks/792EF12A09071278DF8C2D6504F0B9E7.

Narlikar:2020:BT

[Nar20] Jayant V. Narlikar. 104.18 Beetham’s triangle. The Mathematical Gazette, 104(560):327–330, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10418-beethams-triangle/E0C77403B0CE4BE670B72DB131C324D2.

Nathanson:2021:HSC

[Nat21] Melvyn B. Nathanson. 105.06 The Hermite–Sylvester criterion for real-rooted polynomials. The Mathematical Gazette, 105 (562):122–125, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10506-
the-hermitesylvester-criterion-for-realrooted-polynomials/
7F5C5395F235F18A21A6816C0A78799C.

Nathanson:2022:RRP

[Nat22] Melvyn B. Nathanson. 106.03 real-rooted polynomials and a generalised Hermite–Sylvester theorem. The Mathematical Gazette, 106(565):120–124, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10603-realrooted-polynomials-and-a-generalised-
hermitesylvester-theorem/E149E1809654C5E0EA8FCD4AD9D0E2CF.

Nichols:2020:CSV

[Nic20a] Owen Nichols. 104.11 The curious story of a very old puzzle. The Mathematical Gazette, 104(559):171–173, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10411-the-curious-story-
of-a-very-old-puzzle/A62A10382697F2FE2768E2DFF8CBC6F4.

Nickalls:2020:CC

[Nic20b] R. W. D. Nickalls. 104.04 The complementary cubic. The Mathematical Gazette, 104(559):155–158, March 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10404-the-complementary-
cubic/0EA86B611918DC72B81BE6F2C08D96CA.

Nickalls:2021:SPR

[Nic21] R. W. D. Nickalls. On the structure of polynomial roots. The Mathematical Gazette, 105(563):253–262, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/on-the-structure-of-polynomial-roots/
34793671C17FA38994797585EAAE1960.

Niel:2021:RAT

[Nie21] Blanca Isabel Niel. 105.10 Right-angled triangles link circular and hyperbolic functions. The Mathematical Gazette, 105(562): 134–135, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10510-
rightangled-triangles-link-circular-and-hyperbolic-functions/
9A26F039234E28A84F74F65CC149D0FD.

Nimbran:2022:EEM

[Nim22] Amrik Singh Nimbran. Evolution of the Euler–Maclaurin sum formula. The Mathematical Gazette, 106(567):443–457, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/evolution-
of-the-eulermaclaurin-sum-formula/656A224808B2D98C05D818FE6B8BEBBE.

Nadarajah:2022:TIF

[NO22] Saralees Nadarajah and Idika E. Okorie. On the tail integral formulae for real-valued random variables. The Mathematical Gazette, 106(567):487–493, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-the-tail-integral-formulae-for-realvalued-
random-variables/D4449FA5D3CFB01380413B8B27ECE946.

Northshield:2020:TC

[Nor20] S. Northshield. Tropical cycles. The Mathematical Gazette, 104(560):225–234, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/tropical-cycles/03198208E593E69B1EF99C7A916DA199.

Nowicki:2020:SFE

[Now20] Andrzej Nowicki. 104.05 Some finite extensions of the rational numbers. The Mathematical Gazette, 104(559):159– 160, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10405-
some-finite-extensions-of-the-rational-numbers/C927F3485AC588CA50D8C11C5D09A280.

Nurkanovic:2023:CII

[NT23] Mehmed Nurkanović and Mirsad Trumić. Computing indefinite integrals by difference equations. The Mathematical Gazette, 107(570):474–487, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/computing-indefinite-integrals-by-difference-
equations/B94F66E58B0887169D737DD435141EDE.

Ohyama:2020:ECS

[Ohy20] Hiroshi Ohyama. 104.28 Extending Cardano’s solution of the cubic. The Mathematical Gazette, 104(561):511–517, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10428-
extending-cardanos-solution-of-the-cubic/66E61A3E301939FAF57B488B31D65F93.

Ohyama:2022:ISQ

[Ohy22] Hiroshi Ohyama. 106.18 Impossibility of solving the quintic using Cardano’s solution. The Mathematical Gazette, 106(566): 312–315, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10618-
impossibility-of-solving-the-quintic-using-cardanos-
solution/3F766AF12394298A0E58CEF5D1E69E49.

Ohyama:2023:FSQ

[OI23] Hiroshi Ohyama and Koichiro Ike. 107.05 The final solution of a quasi-palindromic. The Mathematical Gazette, 107(568):130– 136, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10705-the-
final-solution-of-a-quasipalindromic/CAEB89E3625ACD2170599347345C5C5E.

Okumura:2023:EPW

[Oku23] Hiroshi Okumura. 107.39 Enomoto’s problem in Wasan geometry. The Mathematical Gazette, 107(570):514–516, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10739-enomotos-
problem-in-wasan-geometry/9A8AEA04FBEFE4457140FEC94B3C6BB3.

Ortega:2020:API

[Ort20] Ryan N. Ortega. An alternative proof of the integral of a logarithm. The Mathematical Gazette, 104(560):344–345, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/an-alternative-proof-of-
the-integral-of-a-logarithm/B2485AC521E240D5A35B71F4B9B6C0CB.

Oxman:2022:PTI

[OS22] Victor Oxman and Moshe Stupel. 106.10 PWW: Trigonometric inequality. The Mathematical Gazette, 106(565):134, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10610-pww-
trigonometric-inequality/8FDA126A602AB85DE83C1CF7E5033C1D.

Oxman:2023:DTE

[OS23] Victor Oxman and Moshe Stupel. 107.14 Does a trapezium exist whose side lengths form a geometric progression? The Mathematical Gazette, 107(569):301–302, July 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10714-does-a-trapezium-
exist-whose-side-lengths-form-a-geometric-progression/
74AD75F27B1A112A2269F7BFEA7971B0.

Ottewill:2024:RCD

[Ott24] Chris Ottewill. The role of convexity in defining regular polyhedra. The Mathematical Gazette, 108(571):61– 68, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/role-
of-convexity-in-defining-regular-polyhedra/F3145ADEB5184226F0F8536158A95FA3.

Paseau:2021:WTN

[Pas21] A. C. Paseau. Why are there no infinite left-sided decimal expansions? The Mathematical Gazette, 105(562):78–86, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/why-are-there-
no-infinite-leftsided-decimal-expansions/A54242E8EE69515FAEE3717CDA6872ED.

Patil:2022:MCH

[Pat22] Sammedkumar M. Patil. 106.36 A method to calculate the harmonic number and other related sums. The Mathematical Gazette, 106(567):506–508, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10636-a-method-to-calculate-the-harmonic-number-
and-other-related-sums/007C98EC90661D02F859B332A9AA2CA2.

Paul:2021:VSM

[Pau21] Prabir Kumar Paul. Visually suggestive but mathematically incorrect. The Mathematical Gazette, 105(562):98–105, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/visually-
suggestive-but-mathematically-incorrect/A712240F83CCBC8AD6EFDACA4C9412DC.

Pellegrinetti:2022:ESP

[PdV22] Dario Pellegrinetti and Michael de Villiers. An extension of the six-point circle theorem for a generalised Van Aubel configuration. The Mathematical Gazette, 106(567): 400–407, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-extension-of-the-sixpoint-circle-theorem-
for-a-generalised-van-aubel-configuration/7A873D773AF6E60E2AA65C804FE486AB.

Pellegrinetti:2022:SPB

[Pel22] Dario Pellegrinetti. 106.11 On a synthetic proof of Bottema’s theorem. The Mathematical Gazette, 106(565):135–136, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10611-on-
a-synthetic-proof-of-bottemas-theorem/49CEBD5EB08BE68558301DE42DAF08E2.

Plaza:2020:MZZ

[Pla20a] Ángel Plaza. 104.14 More on zero-over-zero limits of special type. The Mathematical Gazette, 104(560):310–313, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10414-more-on-
zerooverzero-limits-of-special-type/751B5E152686B0D627E473F733F1FE65.

Plaza:2020:PWM

[Pla20b] Ángel Plaza. 104.22 Proof without words: Minimum perimeter of an inscribed quadrangle to a square. The Mathematical Gazette, 104(560):338–339, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10422-proof-without-words-minimum-perimeter-
of-an-inscribed-quadrangle-to-a-square/25FDB8C144816D437442396FD219D569.

Plaza:2022:FBP

[Pla22a] Ángel Plaza. 106.07 A function-based proof of the harmonic mean — geometric mean — arithmetic mean inequalities. The Mathematical Gazette, 106(565):130–131, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10607-a-functionbased-proof-of-the-harmonic-
mean-geometric-mean-arithmetic-mean-inequalities/61AA72E9B6B3A99B4A12F6C768665CFB.

Plaza:2022:PWR

[Pla22b] Ángel Plaza. 106.24 Proof without words: a Riemann sum. The Mathematical Gazette, 106(566):331, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10624-proof-without-words-a-riemann-
sum/8774FD80B362B7E785C6DD29002E40FE.

Porteous:2022:C

[Por22] Hugh Porteous. Correspondence. The Mathematical Gazette, 106(565):155–156, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/correspondence/5205156558C7B4A2A15514960EA452DB.

Phoopha:2023:DM

[PPP23] Niphawan Phoopha, Prapanpong Pongsriiam, and Phakhinkon Napp Phunphayap. Digit maps. The Mathematical Gazette, 107 (568):35–43, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/digit-
maps/38CE608D2E198D834ABD59E6D1B94F3A.

Pratoussevitch:2021:BRJ

[Pra21] Anna Pratoussevitch. Book review: The Joy of SET: the many mathematical dimensions of a seemingly simple card game by Liz McMahon, Gary Gordon, Hannah Gordon and Rebecca Gordon, pp. 306, £16.99 (paper), ISBN 978-0-691-19232-1, Princeton University Press (2017). The Mathematical Gazette, 105(563):375–376, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/joy-of-set-the-many-mathematical-dimensions-
of-a-seemingly-simple-card-game-by-liz-mcmahon-gary-
gordon-hannah-gordon-and-rebecca-gordon-pp-306-1699-
paper-isbn-9780691192321-princeton-university-press-
2017/74C8C149ABC34789B946B0866ECA9545.

Pritchard:2022:FVP

[Pri22] Chris Pritchard. Focus on the visual (the 2022 Presidential address). The Mathematical Gazette, 106(567):386–399, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/focus-on-
the-visual-the-2022-presidential-address/8DADC55CCA549D9E8296B378839DC6D1.

Padmanabhan:2022:WDW

[PS22a] R. Padmanabhan and Alok Shukla. 106.20 When do we have 1 + 1 = 11 and 2 + 2 = 5 ? The Mathematical Gazette, 106(566):319–323, July 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10620-when-do-we-have-1-1-11-and-2-2-
5/B467F321C86FFB6F6C5CB4FB23E4F50F.

Prochno:2022:PWD

[PS22b] Joscha Prochno and Michael Schmitz. A probabilistic way to discover the rainbow. The Mathematical Gazette, 106(565):103–115, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/probabilistic-
way-to-discover-the-rainbow/D6F43C8D54EAF95A7BD01E959D3ACA91.

Read:2022:CIT

[Rea22] Emrys Read. On the class of an integer triangle. The Mathematical Gazette, 106(566):291–299, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/on-the-class-of-an-integer-triangle/
F71A8F7D3DC50636296EC06D157E6073.

Read:2023:ITI

[Rea23] Emrys Read. Integer triangles with integer circumradii. The Mathematical Gazette, 107(569):241–248, July 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/integer-triangles-with-
integer-circumradii/E4F578E83CC6D37207997C0644F193A4.

Retkes:2023:X

[Ret23] Zoltan Retkes. On 107.03. The Mathematical Gazette, 107(569): 358, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/on-10703/
25B4925002210D3CDD28610145650E2E.

Richardson:2021:CIa

[Ric21a] Bill Richardson. 1933 cumulative index. The Mathematical Gazette, 105(563):270, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/1933-cumulative-index/FA78FAA4A03CF0E60F76244A6BABA7C9.

Richardson:2021:CIb

[Ric21b] Bill Richardson. 1933 cumulative index. The Mathematical Gazette, 105(563):348, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/1933-cumulative-index/9DF2E3C6C686EC9BD7D1436361AAA987.

Richardson:2024:C

[Ric24] Bill Richardson. Correspondence. The Mathematical Gazette, 108(571):166, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/correspondence/210C8DEAB9FAA93CE421562D5B3877D4.

Ridley:2021:CCG

[Rid21a] J. N. Ridley. 105.14 Cubes, cones and the Gauss–Bonnet theorem. The Mathematical Gazette, 105(562):148–153, March 2021.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10514-cubes-cones-and-the-
gaussbonnet-theorem/7878A70E595164170A3ABB21B39E3B26.

Ridley:2021:RS

[Rid21b] J. N. Ridley. Rectangles and spirals. The Mathematical Gazette, 105(564):416–424, November 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/rectangles-and-spirals/B8AC7D6DEB9D6EA77897DE2C26424367.

Roberts:2020:SP

[Rob20] Lewis Roberts. Student problems. The Mathematical Gazette, 104(559):184–186, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/1E278F6E448F1C22FF5D4890F188884F.

Roidos:2023:SF

[Roi23] Nikolaos Roidos. 107.34 On a staircase function. The Mathematical Gazette, 107(570):501–503, November 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10734-on-a-staircase-function/B7AF4306CA541D651D74D3106B7A92E6.

Rout:2020:BRA

[Rou20a] Stephen Rout. Book review: Algebraic inequalities by Hayk Sedrakyan and Nairi Sedrakyan, pp. 243, £32.99 (hard), ISBN 978-3-319-77835-8, Springer Verlag (2018). The Mathematical Gazette, 104(560):373–374, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/algebraic-inequalities-by-hayk-sedrakyan-and-
nairi-sedrakyan-pp-243-3299-hard-isbn-9783319778358-
springer-verlag-2018/D26B49FFB8EDF087308A00CCC8D6B0EC.

Rout:2020:BRS

[Rou20b] Stephen Rout. Book review: The stair-step approach in mathematics, by Hayk Sedrakyan and Nairi Sedrakyan, pp. 530, £64.99 (hard), ISBN 978-3-319-70631-3, Springer Verlag (2018). The Mathematical Gazette, 104(560):376–378, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/stairstep-approach-in-mathematics-by-
hayk-sedrakyan-and-nairi-sedrakyan-pp-530-6499-hard-
isbn-9783319706313-springer-verlag-2018/CACBFE6738E7E7CBACD33EE65BC100E5.

Rowland:2022:MNN

[Row22] Tim Rowland. Mathematics in ‘the news’: number theory and number sense. The Mathematical Gazette, 106 (567):467–473, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/mathematics-in-the-news-number-theory-and-number-
sense/64ACD0312D5E69B4FCDB7E23E822BEB3.

Ruane:2023:BRE

[Rua23a] P. N. Ruane. Book review: Emmy Noether — mathematician extraordinaire by David E. Rowe, pp 339, £74.99 (paper), ISBN 978-3-030-63812-2, Springer Verlag (2021). The Mathematical Gazette, 107(570):559–560, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/emmy-noether-mathematician-extraordinaire-by-
david-e-rowe-pp-339-7499-paper-isbn-9783030638122-springer-
verlag-2021/41F996DAF5B78DD0B6892BAB6A7EAED3.

Ruane:2023:BRD

[Rua23b] P. N. Ruane. Book review: The doctrine of triangles by Glen Van Brummelen, pp. 376, £25 (hard), ISBN 978-0-69117-941-4, Princeton University Press (2021). The Mathematical Gazette, 107(568):186–188, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/doctrine-of-triangles-by-glen-van-brummelen-
pp-376-25-hard-isbn-9780691179414-princeton-university-
press-2021/69458FB978F8113DAB09B7C5198C4A80.

Rossetto:2024:CRG

[RV24] Silvano Rossetto and Giovanni Vincenzi. A characterisation of regular n-gons via (in)commensurability. The Mathematical Gazette, 108(571):43–52, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/

article/characterisation-of-regular-ngons-via-incommensurability/
CE078BB48F9D21D7B767F41DB00213FA.

Ryba:2020:JHC

[Ryb20] Alex Ryba. John Horton Conway FRS: 26 December 1937– 11 April 2020 : Honorary Member of The Mathematical Association 2017. The Mathematical Gazette, 104(561):395– 402, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/john-
horton-conway-frs/EACD432A66B76F7169A050CB7E1E166E.

Rose:2021:PBC

[RZ21] David Rose and Li Zhou. 105.33 A property of bifocal conics, illuminated by Ptolemy’s theorem. The Mathematical Gazette, 105 (563):331–332, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10533-
a-property-of-bifocal-conics-illuminated-by-ptolemys-
theorem/6E5721E1DC33C4E945A0BCD4CA66380D.

Stubbs:2020:PEL

[SA20] John Stubbs and Jacob Adetunji. Paying for end of life care in the UK. The Mathematical Gazette, 104(561):495–506, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/paying-for-end-
of-life-care-in-the-uk/52FCB05577108A7A5E2DBE3EA90C3FA3.

Sa:2023:SPa

[Sa23a] Tuya Sa. Student problems. The Mathematical Gazette, 107(568):174–176, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/C269D1D70C8F6E1033327915A6DE71E6.

Sa:2023:SPb

[Sa23b] Tuya Sa. Student problems. The Mathematical Gazette, 107(569):365–367, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/453233BF48E2BC460651249D382DCA5A.

Sa:2023:SP

[Sa23c] Tuya Sa. Student problems. The Mathematical Gazette, 107 (570):556–558, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/CED2F6E033CA85CDFB529F68CC1AC55D.

Stubbs:2023:RFD

[SA23d] John Stubbs and Jacob Adetunji. The repayment of financial debt: some mathematical considerations. The Mathematical Gazette, 107(569):204–217, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/repayment-of-financial-debt-some-mathematical-
considerations/78FC37F52F78F323C1F8DD5448127A9D.

Sa:2024:SP

[Sa24] Tuya Sa. Student problems. The Mathematical Gazette, 108(571):172–174, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/DF07ADAA5B0EFD6737061EEC7AED069A.

Samtani:2023:CCC

[Sam23] Sur Samtani. 107.02 Collatz conjecture: coalescing orbits and conditions on a minimum counterexample. The Mathematical Gazette, 107(568):123–126, March 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10702-collatz-conjecture-coalescing-
orbits-and-conditions-on-a-minimum-counterexample/152EAADB9F3425F567F357CC1DC9B966.

Sangwin:2023:SFO

[San23] Chris Sangwin. Sums of the first n odd integers. The Mathematical Gazette, 107(568):10–24, March 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/sums-of-the-first-n-odd-
integers/F5EF6F7261C11FF7793A543CB0CFC7E9. See comment

[Ste23b].

Sangwin:2024:C

[San24] Chris Sangwin. Correspondence. The Mathematical Gazette, 108(571):166, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/correspondence/A44084CD206A9977D2EFAA3696728A8C.

Saouter:2020:NPS

[Sao20] Yannick Saouter. New pancake series for π. The Mathematical Gazette, 104(560):296–303, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/new-pancake-series-for/4DE51EA2A1D97B4F9925FCA9F4AF1FF0.

Sarp:2023:VCB

[Sar23] Umit Sarp. Visualising connections between types of polygonal number. The Mathematical Gazette, 107(568):56–64, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/visualising-connections-
between-types-of-polygonal-number/07657392E2E8E5F366726C6C45FCE007.

Schilling:2020:BRC

[Sch20] René L. Schilling. Book review: Calculus of variations by Filip Rindler, pp. 444, £39.99 (paper), ISBN 978-3-319-77636-1, also available as e-book, Springer Verlag (2018). The Mathematical Gazette, 104(559):191, March 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/calculus-of-variations-
by-filip-rindler-pp-444-3999-paper-isbn-9783319776361-
also-available-as-ebook-springer-verlag-2018/396EDE9306F55B43573A4BE132802555.

Schilling:2021:BRM

[Sch21a] René L. Schilling. Book review: Mathematical constants II by Steven R. Finch, pp. 769, £125 (hard), ISBN 978-1-10847-059-9, Cambridge University Press (2018). The Mathematical Gazette, 105(563):380–381, July 2021. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/mathematical-constants-ii-by-steven-
r-finch-pp-769-125-hard-isbn-9781108470599-cambridge-
university-press-2018/DAD50ED6F86CAEEBE24A2DFE0C9FD3D5.

Schmitz:2021:SEP

[Sch21b] Michael Schmitz. 105.02 Some essential properties of Farey series in one Pic(k)ture. The Mathematical Gazette, 105(562): 108–111, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10502-
some-essential-properties-of-farey-series-in-one-pickture/
3BB59FB138F322D3A4580C17F86C98BD.

Scimone:2022:GRT

[Sci22] Aldo Scimone. Golden right triangles and the golden quadrilateral. The Mathematical Gazette, 106(565):9–20, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/golden-right-triangles-
and-the-golden-quadrilateral/E5C2686BC2899DAB6826F2030EAEBA8C.

Scimone:2024:GTF

[Sci24] Aldo Scimone. 108.16 Golden triangles founded on Kepler’s triangle. The Mathematical Gazette, 108(571):152–154, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10816-golden-triangles-
founded-on-keplers-triangle/655744B93C8B7BF51AD2AC4BDAE91908.

Scott:2020:HRT

[Sco20] J. A. Scott. 104.09 A harmonic range for the triangle. The Mathematical Gazette, 104(559):168–169, March 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10409-a-harmonic-range-
for-the-triangle/3B808EF47007556DEE7A018AD07F8696.

Scott:2022:GIT

[Sco22a] J. A. Scott. 106.46 On the Gerretsen inequalities in trigonometrical form. The Mathematical Gazette, 106(567):532–533, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10646-on-
the-gerretsen-inequalities-in-trigonometrical-form/422C80607EB72B36B0BA0D736BCF6E4C.

Scott:2022:FFP

[Sco22b] J. A. Scott. 106.48 On the first Fermat point for the triangle. The Mathematical Gazette, 106(567):539–541, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10648-on-
the-first-fermat-point-for-the-triangle/D4DFAB02359A64387960AD90F43DEFD9.

Scott:2023:LLT

[Sco23] J. A. Scott. 107.42 On the Lemoine line for the triangle. The Mathematical Gazette, 107(570):526–528, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10742-on-
the-lemoine-line-for-the-triangle/632C0582AE7FDB459CC1186A5EAD34D6.

Sadek:2020:PFN

[SE20a] Jawad Sadek and Russell Euler. 104.03 On periods of Fibonacci numbers using modular arithmetic on the Binet formula. The Mathematical Gazette, 104(559):150–154, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10403-on-
periods-of-fibonacci-numbers-using-modular-arithmetic-
on-the-binet-formula/30FE1AC16B389F17D16C5CBB420D9448.

Sadek:2020:UFL

[SE20b] Jawad Sadek and Russell Euler. 104.26 Unusual Fibonacci, Lucas and Pell congruence relations. The Mathematical Gazette, 104(561):507–509, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10426-unusual-fibonacci-lucas-and-pell-congruence-
relations/E22427749D42B30385C1C8CCD8D8E902.

Sebastian:2022:TDC

[Seb22] Sabu Sebastian. 106.34 A theorem on divisibility by congruence. The Mathematical Gazette, 106(567):501–504, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10634-a-theorem-on-divisibility-
by-congruence/2287DCC35ADCAF8A69D6C49E17A8B255.

Shahali:2024:RPP

[Sha24] H. A. Shahali. 108.03 Remarks on perfect powers. The Mathematical Gazette, 108(571):122–124, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10803-remarks-on-perfect-powers/1741C2E132A8B366EDE7036C2A5D0F25.

Shenoy:2023:DRT

[She23] Rohan Manojkumar Shenoy. 107.27 The discrete renewal theorem with bounded interevent times. The Mathematical Gazette, 107(569):343–348, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10727-the-discrete-renewal-theorem-with-bounded-
interevent-times/888F0FD065BCD42A908580FFD58CA59B.

Shiu:2020:TST

[Shi20] Peter Shiu. The three-square theorem of Gauss and Legendre. The Mathematical Gazette, 104(560):209–214, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/threesquare-theorem-of-
gauss-and-legendre/F5E4C8F7CD8BEC11CCB4F89488414648.

Shiu:2022:BRN

[Shi22a] Peter Shiu. Book review: A new year’s present from a mathematician by Snezana Lawrence, pp. 177, £29.99 (paper), ISBN 978-0-36721-936-9, CRC Press (2019). The Mathematical Gazette, 106(567):563, November 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/new-years-present-from-
a-mathematician-by-snezana-lawrence-pp-177-2999-paper-
isbn-9780367219369-crc-press-2019/6EDD9FBFAC4DB3C3DCA0FC02EABBA7C4.

Shiu:2022:BRM

[Shi22b] Peter Shiu. Book review: The maths of life and death by Kit Yates, pp. 333, £20 (paper), ISBN 978-1-78747-542-7, Quercus Books (2019). The Mathematical Gazette, 106(565): 189–190, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/maths-

of-life-and-death-by-kit-yates-pp-333-20-paper-isbn-
978178747-5427-quercus-books-2019/170FC1ADF251D7B007EDD81B0180EBD5.

Shiu:2023:BRB

[Shi23a] Peter Shiu. Book review: Bounded gaps between primes by Kevin Broughan, pp. 590, £39.99 (paper), ISBN 978-1-108-79920-1, Cambridge University Press (2021). The Mathematical Gazette, 107(568):181–183, March 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/bounded-gaps-between-primes-by-kevin-
broughan-pp-590-3999-paper-isbn-9781108799201-cambridge-
university-press-2021/FF9D6F9D36450D15C522983221D3195B.

Shiu:2023:BRC

[Shi23b] Peter Shiu. Book review: Complex analysis by Andrei Bourchtein and Ludmila Bourchtein, pp. 346, £54.99 (paper), ISBN 978-981-15-9218-8, Springer Verlag (2021). The Mathematical Gazette, 107(570):569–570, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/complex-analysis-by-andrei-bourchteinand-ludmila-
bourchtein-pp-346-5499-paper-isbn-9789911592218-springer-
verlag-2021/5F60B5C95154579F312F7D4F2CE2A11E.

Shiu:2023:BRE

[Shi23c] Peter Shiu. Book review: Essential mathematics for undergraduates by Simon G. Chiossi, pp 490, £54.99 (hard), ISBN 978-3-030-87173-4, Springer Verlag (2021). The Mathematical Gazette, 107(570):563–565, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/essential-mathematics-for-undergraduates-by-
simon-g-chiossi-pp-490-5499-hard-isbn-9783030871734-
springer-verlag-2021/9F9C2B2147F17224BADA63BFDDD7B922.

Shiu:2024:BRI

[Shi24] Peter Shiu. Book review: Irrationality and transcendence in number theory by David Angell, pp. 242, £59.99, (hard), ISBN 978-0-367-62837-6, Chapman and Hall/CRC (2022). The Mathematical Gazette, 108(571):180–182, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/irrationality-and-transcendence-in-number-
theory-by-david-angell-pp-242-5999-hard-isbn-9780367628376-
chapman-and-hallcrc-2022/63CFF2C64BAE74C9063EBA8E914F800D.

Sibley:2023:HEE

[Sib23] Thomas Q. Sibley. How effective is the efficiency gap? The Mathematical Gazette, 107(569):218–224, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/how-effective-is-the-efficiency-gap/
C52C7C83E1DA83058634C8AEDFE8E11B.

Siems:2020:MCM

[Sie20] Tobias Siems. Markov Chain Monte Carlo on finite state spaces. The Mathematical Gazette, 104(560):281–287, July 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/markov-chain-monte-carlo-
on-finite-state-spaces/F410EAF29400580A62DEE5B7C73134CB.

Silvester:2020:DPK

[Sil20] John R. Silvester. Desargues, Pascal and Kirkman. The Mathematical Gazette, 104(559):125–135, March 2020. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/desargues-pascal-and-kirkman/1F4285CFB2C8842536D3096B5E3BCFE3.

Silvester:2021:FFR

[Sil21a] John R. Silvester. 105.21 Factorial factors revisited. The Mathematical Gazette, 105(563):301–303, July 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10521-factorial-factors-
revisited/9F2E47D8E32A42BF51D18E5B051C5184.

Silvester:2021:CMT

[Sil21b] John R. Silvester. On cardioids and Morley’s theorem. The Mathematical Gazette, 105(562):40–51, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/on-cardioids-and-morleys-theorem/4DF2DF083F16857327410C931C97B00D.

Silvester:2022:TLP

[Sil22] John R. Silvester. The trisectrix and Langley’s problem. The Mathematical Gazette, 106(565):21–27, March 2022. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/trisectrix-and-langleys-
problem/79CAD1D7AB6148D5C5CAA328CF865C9B.

Silvester:2023:CCT

[Sil23a] John R. Silvester. 107.44 On converses of circle theorems. The Mathematical Gazette, 107(570):530–532, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10744-on-
converses-of-circle-theorems/0AF3379299A47B20B2FB1E0C0F45C4D5.

Silvester:2023:STT

[Sil23b] John R. Silvester. Some triangle theorems by complex numbers. The Mathematical Gazette, 107(570):454–473, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/some-triangle-theorems-
by-complex-numbers/AA6213AFC04573130FEE6FEB21376FBB.

Silberger:2024:RPB

[Sil24] Allan J. Silberger. 108.05 Ramanujan’s proof of Bertrand’s postulate. The Mathematical Gazette, 108(571):130–134, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10805-ramanujans-
proof-of-bertrands-postulate/72FED9594A96508B22FEFEAA80549817.

Singh:2020:IR

[Sin20] Angad Singh. 104.30 An integral relating π and ζ(3). The Mathematical Gazette, 104(561):520–522, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10430-an-integral-relating-and-3/B2D24B72A8CB7788D18E93AB1490E9E3.

Singh:2021:ICF

[Sin21] Angad Singh. 105.19 An inverted cone and Fermat’s last theorem. The Mathematical Gazette, 105(563):298, July 2021.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10519-an-inverted-cone-
and-fermats-last-theorem/56511E11C3B5696A30C91883EBFD2090.

Singh:2022:NDL

[Sin22] Angad Singh. 106.01 the number of divisors of the LCM of the first n natural numbers. The Mathematical Gazette, 106(565): 116–117, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10601-
the-number-of-divisors-of-the-lcm-of-the-first-n-natural-
numbers/CCEA08EC1E3F75A0B2DEC4B9DB8DE913.

Singh:2024:CI

[Sin24] Angad Singh. 108.08 Cone and integral. The Mathematical Gazette, 108(571):140–142, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10808-cone-and-integral/6D258C2A6026AE6429DAE45C04411E6E.

Skorczewski:2020:MLY

[Sko20] Tyler Skorczewski. Modelling learning in youth archery. The Mathematical Gazette, 104(561):427–434, November 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/modelling-learning-in-youth-
archery/39FE0CCAB78CD4B608EB7C1413011FD8.

Slomson:2024:BRO

[Slo24] Alan Slomson. Book review: The one true logic: a monist manifesto, by Owen Griffiths and A. C. Paseau, pp 232, ISBN 978-0-19-882971-3, Oxford University Press (2022). The Mathematical Gazette, 108(571):186–188, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/one-true-logic-
a-monist-manifesto-by-owen-griffiths-and-a-c-paseau-
pp-232-isbn-9780198829713-oxford-university-press-2022/
96B1D4DCC0C80C6E7B6CE29E8B3A9C60.

Stupel:2023:TWS

[SO23] Moshe Stupel and Victor Oxman. 107.10 Trapezia whose sidelengths form an arithmetic progression. The Mathematical Gazette, 107(568):147–149, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10710-trapezia-whose-sidelengths-form-an-arithmetic-
progression/28BB79428D783C14212B5FDEC67DCA6E.

Sporn:2021:GPT

[Spo21a] Howard Sporn. A group of Pythagorean triples using the inradius. The Mathematical Gazette, 105(563):209–215, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/group-of-
pythagorean-triples-using-the-inradius/E88C97B72AEBE740093A07E1ABAC41E7.

Sporn:2021:GSG

[Spo21b] Howard Sporn. A group structure on the golden triples. The Mathematical Gazette, 105(562):87–97, March 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/group-structure-on-the-
golden-triples/2C00D0296460B3EF961CBE421182B394.

Sporn:2022:FLH

[Spo22a] Howard Sporn. Fibonacci–Lucas hyperbolas. The Mathematical Gazette, 106(566):242–246, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonaccilucas-hyperbolas/0B51B37A75BC9B51020FDF631AA6169C.

Sporn:2022:FFC

[Spo22b] Howard Sporn. Fibonacci fraction circles. The Mathematical Gazette, 106(565):1–8, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonacci-fraction-circles/40C40BE93CB0477893321A2F62FCA713.

Sporn:2023:MAA

[Spo23] Howard Sporn. Multiplication is to addition as addition is to what? The Mathematical Gazette, 107(568):84–95, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/multiplication-
is-to-addition-as-addition-is-to-what/69855BFEFA2463E3593BCBA626C51121.

Sporn:2024:PAS

[Spo24] Howard Sporn. Patterns among square roots of the 2 × 2 identity matrix. The Mathematical Gazette, 108(571):84–93, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/patterns-among-
square-roots-of-the-2-2-identity-matrix/456EDD8B24409B9CB8BF4F7BE5DA179B.

Skurnick:2020:MCM

[SR20] Ronald Skurnick and Christopher Roethel. A more conclusive and more inclusive second derivative test. The Mathematical Gazette, 104(560):247–254, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/more-conclusive-and-more-inclusive-second-derivative-
test/7D48E66EF263873083E8FA573AEF51AB.

Switkes:2020:WAE

[SS20] Jennifer Switkes and Randall Swift. Worst average encountered highway velocity. The Mathematical Gazette, 104(559):12–19, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/worst-average-
encountered-highway-velocity/B257A43CA64BD559086FE5A8709D7314.

Shahbari:2022:TI

[SS22] Juhaina A. Shahbari and Moshe Stupel. 106.13 A triangle inequality. The Mathematical Gazette, 106(565):138, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10613-a-triangle-inequality/
A88AD35D6F5D1D166B2D04396CDFE821.

Shahali:2024:WG

[SS24] Manija Shahali and H. A. Shahali. Walk on a grid. The Mathematical Gazette, 108(571):111–117, March 2024. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/walk-on-a-grid/6BDE88C625E149F29116C304A29282AB.

Subramaniam:2022:PWS

[ST22] K. B. Subramaniam and Aji Thomas. 106.23 Proof without words: sin 3x = 3 sin x − 4 sin^{3} x. The Mathematical Gazette, 106(566):330, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10623-
proof-without-words-sin-3x-3-sin-x-4-sin3x/7DD6F4D3BEC3F5A17B4A3E702CDAAB61.

Sangwin:2023:DNP

[ST23] Chris Sangwin and Fenner Stanley Tanswell. Developing new picture proofs that the sums of the first n odd integers are squares. The Mathematical Gazette, 107(569):249–262, July 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/developing-new-picture-
proofs-that-the-sums-of-the-first-n-odd-integers-are-
squares/6B0F55C5036E9C7A4621294F9A4244A8.

Stanley:2020:TCS

[Sta20] P. Stanley. 104.23 Tessellation of a conical surface. The Mathematical Gazette, 104(560):339–341, July 2020. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10423-tessellation-of-a-
conical-surface/1592102B77AE547C66294EB33C4C31CA.

Stanley:2021:FFT

[Sta21a] P. Stanley. On the frequency of Friday the thirteenth. The Mathematical Gazette, 105(563):222–225, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/on-the-frequency-of-friday-the-thirteenth/
579BAAFD36E2C57C0CEB629DB2F1166C.

Starr:2021:NLC

[Sta21b] Chris Starr. 105.18 Notes on listener crossword 4595 by Elap. The Mathematical Gazette, 105(563):291–298, July 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10518-notes-on-listener-
crossword-4595-by-elap/7A469A2E4FE3BCC7AACABF16A8D7963C.

Stephenson:2020:CSP

[Ste20a] Paul Stephenson. 104.06 A consecutive squares property. The Mathematical Gazette, 104(559):160–161, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10406-a-consecutive-squares-property/
9E86EDE6AC2373586F2678E9EB2B42A5.

Stephenson:2020:CSV

[Ste20b] Paul Stephenson. 104.19 Centroid of a sector from Viète. The Mathematical Gazette, 104(560):330–331, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10419-centroid-of-a-sector-from-viete/
77714328CAC7CA05717C6B2190CFFEFB.

Stewart:2020:CCI

[Ste20c] Seán M. Stewart. A Catalan constant inspired integral odyssey. The Mathematical Gazette, 104(561):449–459, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/catalan-constant-inspired-
integral-odyssey/48229249A53E7D5F6C5BCDEB42233726.

Stephenson:2021:FDS

[Ste21a] Paul Stephenson. On ‘A family of discrete spirals’. The Mathematical Gazette, 105(563):354–355, July 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/on-a-family-of-discrete-
spirals/F8AED4E800E32880519E3AAAC9A7CBA3.

Stewart:2021:PSR

[Ste21b] Seán M. Stewart. A pretty series revisited. The Mathematical Gazette, 105(564):450–457, November 2021. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/pretty-series-revisited/
E29CCC38B5192FA27DF01351D8334CC4.

Stephenson:2022:HIC

[Ste22a] Paul Stephenson. On ‘How to impress a chemist — again!’. The Mathematical Gazette, 106(567):551–552, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/on-how-to-impress-a-chemist-
again/4E897421628D3137ABD3BA575D1B465F.

Stewart:2022:BRA

[Ste22b] Sean M. Stewart. Book review: Advanced calculus explored by Hamza Alsamraee, pp. 448, $26.99 (paper), ISBN 978-0-578-61682-7, Curious Math Publications (2019). The Mathematical Gazette, 106(565):181–183, March 2022. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/advanced-calculus-explored-by-hamza-
alsamraee-pp-448-2699-paper-isbn-9780578616827-curious-
math-publications-2019/0AEE44787147D97A0F71299F805655F1.

Stewart:2022:LBL

[Ste22c] Seán M. Stewart. A look back at a long-forgotten trigonometric function: the versine function and its inverse. The Mathematical Gazette, 106(565):84–94, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/look-back-at-a-longforgotten-trigonometric-function-
the-versine-function-and-its-inverse/72C0952C4E93DDD8428CEDC6376EED9C.

Stephenson:2023:BPT

[Ste23a] Paul Stephenson. Bicentric polygons, their incircles, circumcircles and the circles between. The Mathematical Gazette, 107(568):160–161, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/bicentric-polygons-their-incircles-circumcircles-
and-the-circles-between/1F4ADCFAFE495F9A31338BF5C3D76218.

Stephenson:2023:SFI

[Ste23b] Paul Stephenson. On ‘Sums of the first n integers’. The Mathematical Gazette, 107(570):545–547, November 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/on-sums-of-the-first-n-
integers/82BBC718C7441D8FA1A65FE93C230202. See [San23].

Stephenson:2023:TIM

[Ste23c] Paul Stephenson. Trigonometric identities from the mystic rose. The Mathematical Gazette, 107(569):349–355, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/

journals/mathematical-gazette/article/trigonometric-
identities-from-the-mystic-rose/E2E3C4A45D606131149EFD7985EC7C33.

Stephenson:2023:WMG

[Ste23d] Paul Stephenson. What makes a good Proof without Words. The Mathematical Gazette, 107(568):165, March 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/what-makes-a-good-proof-
without-words/ABB2F0E1D4B2A114F9B535FB4B9CCDA1.

Stewart:2023:SIR

[Ste23e] Seán M. Stewart. 107.01 A simple integral representation of the Fibonacci numbers. The Mathematical Gazette, 107(568): 120–123, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10701-
a-simple-integral-representation-of-the-fibonacci-numbers/
ADCAC1613B42D6660E37932E748B1A28.

Stephenson:2024:TNI

[Ste24] Paul Stephenson. 108.14 A triangle number identity. The Mathematical Gazette, 108(571):148, March 2024. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10814-a-triangle-number-
identity/BE80E0E4659348E0CC9A20EF76B119B2.

Sullivan:2022:CL

[Sul22] Jerry Sullivan. A computer look at N!. The Mathematical Gazette, 106(566):233–241, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/computer-look-at-n/F71176C8058EA4101299CAA1EF6F7545.

Sury:2020:AAU

[Sur20] B. Sury. 104.27 An additive analogue of an unsolved multiplicative problem. The Mathematical Gazette, 104(561):510–511, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10427-an-
additive-analogue-of-an-unsolved-multiplicative-problem/
778F47B0A352F8793C6013B49117F177.

Stupel:2021:GTA

[SWT21] Moshe Stupel, Shula Weissman, and Idan Tal. Geometrical theorems about conservation of area for use in classroom activities. The Mathematical Gazette, 105(562):61–69, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/geometrical-theorems-about-
conservation-of-area-for-use-in-classroom-activities/
4234C22CA527198A6947987DD42F342B.

Tho:2022:RDP

[Tho22a] Nguyen Xuan Tho. 106.02 the rational distance problem revisited. The Mathematical Gazette, 106(565):117–120, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10602-the-
rational-distance-problem-revisited/9F5B2E87122B06BC2E358D864267FED6.

Tho:2022:ISS

[Tho22b] Nguyen Xuan Tho. 106.05 on the irrationality of sums of square roots. The Mathematical Gazette, 106(565):125–127, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10605-on-
the-irrationality-of-sums-of-square-roots/58AE53E90D31C775B647114449ED0C7D.

Tho:2022:PLC

[Tho22c] Nguyen Xuan Tho. 106.15 A proof of Lukarevski’s conjecture. The Mathematical Gazette, 106(565):143–147, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10615-a-proof-
of-lukarevskis-conjecture/05AA082625DDE1CD29770C50C6208A99.

Tho:2022:IAP

[Tho22d] Nguyen Xuan Tho. 106.27 An interesting application of Ptolemy’s inequality. The Mathematical Gazette, 106(566):338– 340, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10627-an-
interesting-application-of-ptolemys-inequality/F6445A0C40B9106416202F3A4E2C643B.

Tho:2022:III

[Tho22e] Nguyen Xuan Tho. 106.28 Inequalities involving the inradius and altitudes of a triangle. The Mathematical Gazette, 106 (566):341–342, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10628-
inequalities-involving-the-inradius-and-altitudes-of-
a-triangle/1960B1B4B7D1C1A90B4517ADAAE72014.

Toller:2021:BRD

[Tol21a] Owen Toller. Book review: Do dice play God? by Ian Stewart, pp. 320, £15.99 (paper), ISBN 978-1-78816-228-9, Profile Books (2019). The Mathematical Gazette, 105 (562):186–187, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/do-dice-play-god-by-ian-stewart-pp-320-1599-
paper-isbn-9781788162289-profile-books-2019/4769EDDEF5E215284EEAE0C32093BDDF.

Toller:2021:BRG

[Tol21b] Owen Toller. Book review: Game changers by Rudolf Taschner, pp. 237, $ 18 (paper), ISBN 978-1-63388-373-4, Prometheus Books (2017). The Mathematical Gazette, 105(563):378–379, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/game-changers-
by-rudolf-taschner-pp-237-18-paper-isbn-9781633883734-
prometheus-books-2017/07736CEF8A8A3EB5F1830FCDA44874E8.

Toller:2021:BRI

[Tol21c] Owen Toller. Book review: Introductory mathematics and statistics through sport by Tricia Muldoon Brown and Eric B. Kahn, pp. 116, £25 (hard), ISBN 978-0-19-883567-7, Oxford University Press (2019). The Mathematical Gazette, 105(562):187–188, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/introductory-mathematics-and-statistics-through-
sport-by-tricia-muldoon-brown-and-eric-b-kahn-pp-116-
25-hard-isbn-9780198835677-oxford-university-press-2019/
9B89568A2EA774B87EDB44623E972697.

Toller:2021:BRJ

[Tol21d] Owen Toller. Book review: The joy of statistics by Steve Selvin, pp. 211, £19.99 (hard), ISBN 978-0-19-883344-4, Oxford University Press (2019). The Mathematical Gazette, 105(562):186–187, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/joy-of-statistics-by-steve-selvin-pp-211-1999-
hard-isbn-9780198833444-oxford-university-press-2019/
A2C0AC6A5809789A1116154ABB44A5DA.

Toller:2021:BRT

[Tol21e] Owen Toller. Book review: Truth or truthiness, distinguishing fact from fiction by learning to think like a data scientist by Howard Wainer, pp. 210, £19.99 (hard), ISBN 978-1-107-13057-9, Cambridge University Press (2016). The Mathematical Gazette, 105(563):368, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/truth-or-truthiness-distinguishing-fact-from-
fiction-by-learning-to-think-like-a-data-scientist-by-
howard-wainer-pp-210-1999-hard-isbn-9781107130579-cambridge-
university-press-2016/651DA9E15AAAB53DA99DBFB3B0AB95E6.

Toller:2021:IQ

[Tol21f] Owen Toller. An instructive question. The Mathematical Gazette, 105(563):350–353, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-instructive-question/B62FAECEFAAEA3279C63BC289E1D7A36.

Toller:2022:BRMa

[Tol22a] Owen Toller. Book review: A modern introduction to differential equations (3rd edn.) by Henry J. Ricardo, pp. 539, £115 (hard), ISBN 978-0-12-823417-4, Academic Press/Elsevir (2020). The Mathematical Gazette, 106(566):376–377, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/modern-introduction-to-differential-
equations-3rd-edn-by-henry-j-ricardo-pp-539-115-hard-
isbn-9780128234174-academic-presselsevir-2020/B9DE504D22F088183796385903612B99.

Toller:2022:BRI

[Tol22b] Owen Toller. Book review: Introduction to the theory of optimization in Euclidean space by Samia Challal, pp. 318, £73.59 (hard), ISBN 978-0-367-19557-1, also available as ebook, CRC Press (2019). The Mathematical Gazette, 106 (565):185–186, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/introduction-to-the-theory-of-optimization-in-
euclidean-space-by-samia-challal-pp-318-7359-hard-isbn-
9780367195571-also-available-as-ebook-crc-press-2019/
63DA064B9A552473728B39E88DC89E22.

Toller:2022:BRN

[Tol22c] Owen Toller. Book review: Number and letter puzzles by Des MacHale, pp. 98, £9.21 (paper), ISBN 978-0-24423-100-2, Logic Press (2019). The Mathematical Gazette, 106(565):191, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/number-and-letter-puzzles-by-des-machale-pp-
98-921-paper-isbn-9780-244231002-logic-press-2019/A249A35832C65B78888E503C5D216E75.

Toller:2022:BRO

[Tol22d] Owen Toller. Book review: One thousand exercises in probability (3rd edition) by Geoffrey R. Grimmett and David R. Stirzaker, pp. 580, £29.99 (paper), ISBN 978-0198847618, Oxford University Press (2020). The Mathematical Gazette, 106 (567):568–569, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/one-thousand-exercises-in-probability-3rd-edition-
by-geoffrey-r-grimmett-and-david-r-stirzaker-pp-580-
2999-paper-isbn-9780198847618-oxford-university-press-
2020/D2D88478A396F3BD44395D9D96A15D24.

Toller:2022:BRP

[Tol22e] Owen Toller. Book review: Probability and random processes (4th edition) by Geoffrey R. Grimmett and David R. Stirzaker, pp. 669, £40 (paper), ISBN 978-019884759, Oxford University Press (2020). The Mathematical Gazette, 106 (567):568–569, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/probability-and-random-processes-4th-edition-
by-geoffrey-r-grimmett-and-david-r-stirzaker-pp-669-
40-paper-isbn-978019884759-oxford-university-press-2020/
0222046279595EB113671FD07CC80F57.

Toller:2022:BRMb

[Tol22f] Owen Toller. Book review: The mathematics lover’s companion by Edward Scheinerman, pp. 274, £12.99 (paper), ISBN 978-0-300-25539-3, Yale University Press (2021). The Mathematical Gazette, 106(566):383, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/mathematics-lovers-companion-by-edward-scheinerman-
pp-274-1299-paper-isbn-9780300255393-yale-university-
press-2021/34DF5F765F9BCD3CE5C0FD5F9771CE9B.

Toller:2022:BRS

[Tol22g] Owen Toller. Book review: The secret formula by Fabio Toscano, pp. 161, £22 (hard), ISBN 978-0-691-18367-1, Princeton University Press (2020). The Mathematical Gazette, 106(566):366–367, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/secret-formula-
by-fabio-toscano-pp-161-22-hard-isbn-9780691183671-princeton-
university-press-2020/E40C254FD737A0245D8F5223EF447667.

Toller:2022:BRTb

[Tol22h] Owen Toller. Book review: Thinking clearly with data by Ethan Bueno de Mesquita and Anthony Fowler, pp. 432, £25 (paper), ISBN 978-0-69121-435-1, Princeton University Press (2021). The Mathematical Gazette, 106(567): 572–574, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/thinking-clearly-with-data-by-ethan-bueno-de-
mesquita-and-anthony-fowler-pp-432-25-paper-isbn-9780691214351-
princeton-university-press-2021/980F4D18F2D2C822C0717148DAF37B09.

Toller:2022:BRTa

[Tol22i] Owen Toller. Book review: Thinking probabilistically by Ariel Amir, pp. 242, £39.99 (paper), ISBN 978-1-108-78998-1, Cambridge University Press (2021) (e-copy reviewed). The Mathematical Gazette, 106(566):378–379, July 2022. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/thinking-probabilistically-by-ariel-
amir-pp-242-3999-paper-isbn-9781108789981-cambridge-
university-press-2021-ecopy-reviewed/1BA43F1E6B4AB11FE7E4D25DA0F2BDC5.

Toller:2022:X

[Tol22j] Owen Toller. On 105.49. The Mathematical Gazette, 106(565): 158, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/on-10549/
84E04C77881F69A958936D51256D2319.

Toller:2023:BRF

[Tol23a] Owen Toller. Book review: A first course in group theory by Bijan Davvaz, pp. 291, £44.99 (hard), also available as e-book, ISBN 978-981-16-6364-2, Springer Verlag (2021). The Mathematical Gazette, 107(570):565–566, November 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/first-course-in-group-theory-by-bijan-
davvaz-pp-291-4499-hard-also-available-as-ebook-isbn-
9789811663642-springer-verlag-2021/825728B09A114FCF9D6AA38DBDCAA976.

Toller:2023:BRAb

[Tol23b] Owen Toller. Book review: Abstract algebra, a comprehensive introduction by John W. Lawrence and Frank A. Zorzitto, pp. 619, £64.99 (hard), ISBN 978-1-108-83665-4, Cambridge University Press (2021). The Mathematical Gazette, 107(569):372– 373, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/abstract-
algebra-a-comprehensive-introduction-by-john-w-lawrence-
and-frank-a-zorzitto-pp-619-6499-hard-isbn-9781108836654-
cambridge-university-press-2021/E60652F070FAEA359BAEC7367FE53AC7.

Toller:2023:BRA

[Tol23c] Owen Toller. Book review: Algebra, notes from the underground by Paolo Aluffi, pp. 488, £29.99 (paper), ISBN 978-1-108-95823-3, Cambridge University Press (2021). The Mathematical Gazette, 107(570):562–563, November 2023. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/algebra-notes-from-the-underground-by-paolo-
aluffi-pp-488-2999-paper-isbn-9781108958233-cambridge-
university-press-2021/75255C33843681F8CE8F1EF0B9748791.

Toller:2023:BRAa

[Tol23d] Owen Toller. Book review: Anachronisms in the history of mathematics edited by Niccolò Guicciardini, pp. 366, £110 (hard), ISBN 978-1-108-83496-4, Cambridge University Press (2021). The Mathematical Gazette, 107(569):369–372, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/anachronisms-in-the-history-
of-mathematics-edited-by-niccolo-guicciardini-pp-366-
110-hard-isbn-9781108834964-cambridge-university-press-
2021/3F9164647F70C508FA5355CC63E6946C.

Toller:2023:BRB

[Tol23e] Owen Toller. Book review: Basic statistics with R by Stephen C. Loftus, pp 283, £57.95 (paper), ISBN 978-0-12-820788-8, Academic Press/Elsevier (2021). The Mathematical Gazette, 107(569):379–380, July 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/basic-statistics-with-r-by-stephen-c-
loftus-pp-283-5795-paper-isbn-9780128207888-academic-
presselsevier-2021/0BF5A0142A9544DF2FCCFB018F7415A3.

Toller:2023:BRD

[Tol23f] Owen Toller. Book review: Do not erase by Jessica Wynne, pp 252, £30.00 (hard), ISBN 978-0-69119-922-1, Princeton University Press (2021). The Mathematical Gazette, 107(568):191, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/do-not-erase-by-jessica-wynne-pp-252-3000-hard-
isbn-9780691199221-princeton-university-press-2021/40917EF1832D697C8A4BC1A673788D51.

Toller:2023:BRI

[Tol23g] Owen Toller. Book review: Introduction to linear and matrix algebra by Nathaniel Johnston, pp. 482, £49.99 (hard), ISBN 978-3-03052-810-2, Springer Verlag (2021). The Mathematical Gazette, 107(568):188–189, March 2023. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/introduction-to-linear-
and-matrix-algebra-by-nathaniel-johnston-pp-482-4999-
hard-isbn-9783030528102-springer-verlag-2021/4ABF2DE4473BF903BF7D7A9BC1896EA9.

Toller:2023:BRW

[Tol23h] Owen Toller. Book review: What are the chances of that? by Andrew Elliott, pp. 356, £25 (hard), ISBN 978-0-19-886902-3, Oxford University Press (2021). The Mathematical Gazette, 107 (569):381–382, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/what-
are-the-chances-of-that-by-andrew-elliott-pp-356-25-
hard-isbn-9780198869023-oxford-university-press-2021/
47B99232D00C4207E4F39E544F4C7B10.

Toller:2023:X

[Tol23i] Owen Toller. On 106.34. The Mathematical Gazette, 107(569): 356–358, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/on-10634/
BC929E5C9B64EFE3EE545C48D0D14E5F.

Toller:2024:BRS

[Tol24] Owen Toller. Book review: Science by simulation, volume 1 by Andrew French, pp 288, £40 (paper), ISBN 978-1-80061-121-4, World Scientific (2022). The Mathematical Gazette, 108(571):185–186, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/science-by-simulation-volume-1-by-andrew-french-
pp-288-40-paper-isbn-9781800611214-world-scientific-
2022/92356A6887C1B3DBE64455FB873D5D9C.

Tossavainen:2023:CBA

[Tos23] Timo Tossavainen. Cryptography based on algebraic perpendicularities. The Mathematical Gazette, 107(568):65–69, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/cryptography-based-on-algebraic-
perpendicularities/0DA495AF75CA913027E3934A536DB7BA.

Tran:2023:SGW

[Tra23] Quang Hung Tran. Some generalisations of Weitzenböck’s inequality. The Mathematical Gazette, 107(570):405–411, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/some-
generalisations-of-weitzenbocks-inequality/EFB8243BFC597F72A8AFC1DC87DA6E75.

Tran:2024:SGE

[Tra24] Quang Hung Tran. Some generalisations and extensions of a remarkable geometry puzzle. The Mathematical Gazette, 108 (571):36–42, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/some-
generalisations-and-extensions-of-a-remarkable-geometry-
puzzle/281A761BCB3227B94A647B9819452F6C.

Tho:2024:FLE

[TT24] Nguyen Xuan Tho and Nguyen Quynh Tram. 108.02 Fermat-like equations for fractional parts. The Mathematical Gazette, 108 (571):120–122, March 2024. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10802-
fermatlike-equations-for-fractional-parts/83E93DB70CD11E9B1E54CC96A611D877.

Uhlmann:2023:CER

[Uhl23] Jeffrey Uhlmann. 107.29 A cautionary example relating to the interpretation of numerical computations. The Mathematical Gazette, 107(570):490–493, November 2023. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10729-a-cautionary-example-relating-
to-the-interpretation-of-numerical-computations/D6D8710999C0B8096D50AD38D38E80B9.

Viglione:2020:TCK

[Vig20] Raymond Viglione. The Thébault configuration keeps on giving. The Mathematical Gazette, 104(559):74–81, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/thebault-configuration-
keeps-on-giving/92093E80A2CA1AFD6D8D40E91C5E6EE0.

Villarino:2023:QHA

[Vil23] Mark B. Villarino. 107.19 A quadratic harmonic approximation. The Mathematical Gazette, 107(569):316–320, July 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10719-a-quadratic-harmonic-
approximation/52042ACA94287700B2DFD5869B2747FC.

Volenec:2022:APU

[Vol22] Vladimir Volenec. 106.49 An area problem using barycentric coordinates. The Mathematical Gazette, 106(567):541–543, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10649-an-area-
problem-using-barycentric-coordinates/1282A9BF1F6354C2ACF8B3D01F117D86.

Vu:2021:MPM

[Vu21] Thanh Tung Vu. 105.11 Median-parallelogic and medianorthologic triangles. The Mathematical Gazette, 105(562): 136–139, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10511-
medianparallelogic-and-medianorthologic-triangles/70306B95CABEFC6159D695DE1AC2FA9E.

Vu:2022:RRN

[Vu22] Thanh Tung Vu. 106.35 On the representation of rational numbers in the form \frac{(ϕ(m^{r}))^{a}}{(ϕ(n^{s}))^{b}}. The Mathematical Gazette, 106(567): 504–505, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10635-
on-the-representation-of-rational-numbers-in-the-form/
EC9CD640E8FBB24765C8BF90CE1B5978.

Viher:2021:ASC

[VVK21] Radimir Viher, Dragutin Viher, and Helena Koncul. Applications of the sine and cosine rules for a quadrilateral. The Mathematical Gazette, 105(562):70–77, March 2021. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/applications-of-the-sine-and-cosine-
rules-for-a-quadrilateral/2E4EDE1AC578FB954DA5C52EB2486C05.

Wapner:2020:GNA

[Wap20] Leonard M. Wapner. GPS navigation apps and the price of anarchy. The Mathematical Gazette, 104(560):235–240, July 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/gps-navigation-apps-and-
the-price-of-anarchy/E9059E0CAFB7D8EE708E659FDA813C5C.

Wapner:2021:UCT

[Wap21] Leonard M. Wapner. An unexpected characteristic of tournament predictive power. The Mathematical Gazette, 105 (563):201–208, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/an-unexpected-characteristic-of-tournament-predictive-
power/236438ED72A1EDDAA4183F20D033398D.

Wapner:2022:PQS

[Wap22] Leonard M. Wapner. Probability: a questionable science of the uncertain. The Mathematical Gazette, 106(567):458–466, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/probability-a-
questionable-science-of-the-uncertain/5B105B401912FEB9A824E911252CB603.

Wapner:2023:SPT

[Wap23]  Leonard M. Wapner.  107.45 A surprising property of

a tennis-like game. The Mathematical Gazette, 107(570): 533–537, November 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10745-a-surprising-property-of-a-tennislike-
game/1DA6D229B68C94075ABA9A82523A4931.

Whitty:2022:HTG

[Whi22] Robin Whitty. 106.47 Halving a triangle in a given direction. The Mathematical Gazette, 106(567):534–538, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10647-halving-
a-triangle-in-a-given-direction/C8EE4718F19B7B08F007C96561488CAA.

Williams:2023:AU

[Wil23] Hollis Williams. 107.03 Archimedes and the ungula. The Mathematical Gazette, 107(568):127–128, March 2023. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10703-archimedes-and-the-ungula/D98065C561D8E5BA5F95ADD18AAC4436.

Woollacott:2021:SP

[Woo21] Beth Woollacott. Student problems. The Mathematical Gazette, 105(563):365–367, July 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/7D2FBB7E102B61A1DFE9CB4CEA0B9724.

Woollacott:2022:SPa

[Woo22a] Beth Woollacott. Student problems. The Mathematical Gazette, 106(565):167–169, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/102E8C5D587B9F1E790C0EDDB9BD83B5.

Woollacott:2022:SPb

[Woo22b] Beth Woollacott. Student problems. The Mathematical Gazette, 106(566):358–360, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/06EB67D328827851336B41B29FB04F6C.

Woollacott:2022:SP

[WS22] Beth Woollacott and Tuya Sa. Student problems. The Mathematical Gazette, 106(567):560–562, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/student-problems/7975388C3F5FDD5E119F9B863BDC2CB7.

Young:2022:X

[YC22] Robert M. Young and Jack Calcut. On 106.06. The Mathematical Gazette, 106(567):549–550, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/on-10606/9731A367850151AA5F2393007558A9FE.

Young:2023:AFN

[YC23] Robert M. Young and Jack Calcut. Amendment to feedback: On 106.06 in November 2022. The Mathematical Gazette, 107(568): 165, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/amendment-
to-feedback-on-10606-in-november-2022/ECC4450EE32376C5EC8BBC521825D56D.

Yeo:2020:BRF

[Yeo20] Dominic Yeo. Book review: A first course in analysis by John B. Conway, £39.99, ISBN 978-1-10717-314-9, Cambridge University Press (2017). The Mathematical Gazette, 104(559): 187, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/first-
course-in-analysis-by-john-b-conway-3999-isbn-9781107173149-
cambridge-university-press-2017/E1DA89FC94606B7A61C940A5595F23D8.

Yeo:2023:BRC

[Yeo23] Dominic Yeo. Book review: Counterexamples in measure and integration by René L. Schilling and Franziska Kühn, pp. 399, £34.99 (paper), ISBN 978-1-00900-162-5, Cambridge University Press (2021). The Mathematical Gazette, 107(568):189–190, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/counterexamples-in-measure-and-integration-by-
rene-l-schilling-and-franziska-kuhn-pp-399-3499-paper-
isbn-9781009001625-cambridge-university-press-2021/D051730B7D9ACA1D9174B432DACD6DBC.

Yoshida:2022:GRR

[Yos22] Norio Yoshida. 106.22 The golden ratio represented by a tangent. The Mathematical Gazette, 106(566):325–329, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10622-the-golden-ratio-
represented-by-a-tangent/0ED6EFC3A896960EBC2CD69A98A293CE.

Zamfir:2020:BRP

[Zam20] Rica Zamfir. 104.33 Bounds for roots of polynomials with increasing coefficients. The Mathematical Gazette, 104(561):530– 532, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10433-
bounds-for-roots-of-polynomials-with-increasing-coefficients/
14097D5F0D13B0E72A867713C458F40C.